### v1

# Segre forms of singular metrics on vector bundles and Lelong numbers

Abstract.
Let $E\to X$ be a holomorphic vector bundle with a singular Griffiths negative Hermitian metric $h$ with analytic singularities. One can define, given a smooth reference metric $h_{0}$, a current $s(E,h,h_{0})$ that is called the associated Segre form, which defines the expected Bott-Chern class and coincides with the usual Segre form of $h$ where it is smooth. We prove that $s(E,h,h_{0})$ is the limit of the Segre forms of a sequence of smooth metrics if the metric is smooth outside the unbounded locus, and in general as a limit of Segre forms of metrics with empty unbounded locus. One can also define an associated Chern form $c(E,h,h_{0})$ and we prove that the Lelong numbers of $s(E,h,h_{0})$ and $c(E,h,h_{0})$ are integers if the singularities are integral, and non-negative for $s(E,h,h_{0})$. We also extend these results to a slightly larger class of singular metrics with analytic singularities.
1. Introduction
Inspired by the importance during several decades of singular metrics on line bundles in algebraic geometry, with origins in [12], several efforts have been done to study singular metrics on higher rank holomorphic vector bundles. In particular, such metrics appeared in the work of Berndtsson-Păun, [9], in connection with pseudoeffectivity of twisted relative canonical bundles. Contrary to the line bundle case, it is not clear what the curvature and connection matrix for a singular metric should be, cf. [20, Theorem 1.5]. This would appear to be an obstacle for defining associated Chern forms (or more general characteristic forms like, e.g., Segre forms), since the standard definition is as certain polynomials in the entries of the curvature matrix. However, for appropriate metrics, it has still been possible to introduce such associated forms, or more precisely currents of order 00, see, e.g., [15],[16],[17],[18] and [21]. In this article, we will further the study of such associated forms in more general situations.
Assume that $E\to X$ is a holomorphic vector bundle over a complex manifold $X$ with a Griffiths negative singular metric $h$ with analytic singularities; this means that $\log|\cdot|_{h}$ is plurisubharmonic (psh) with analytic singularities on $E\setminus 0_{E}$, see Section 3. In [18] was introduced an associated Segre form [^1] We keep the notion ’Segre form’ although in the singular case $s(E,h,h_{0})$ is in general a current (of order 00). that represents the expected Bott-Chern class, and coincides with the usual Segre form where the metric is non-singular. The form $s(E,h,h_{0})$ slightly depends on a choice of smooth reference metric $h_{0}$ on $E$. For a singular metric $h$, we let the unbounded locus $Z=Z_{h}\subset X$ denote the complement of the set of $x\in X$ where $\log|\cdot|^{2}_{h}$ is locally bounded on $E\setminus 0_{E}$ above a neighborhood of $x$. Assume for the moment that in addition $h$ is smooth outside of $Z$. Let $h_{\epsilon}=h+{\epsilon}h_{0}$, i.e.,
(1.1) $|\cdot|^{2}_{h_{\epsilon}}=|\cdot|^{2}_{h}+{\epsilon}|\cdot|^{2}_{h_{0}}.$
Then $h_{\epsilon}$ is a smooth metric on $E$. Thus the associated Segre forms $s(E,h_{\epsilon})$ are well-defined and our first main result states that
(1.2) $s(E,h,h_{0})=\lim_{{\epsilon}\to 0}s(E,h_{\epsilon}).$
This further suggests that the definition of $s(E,h,h_{0})$ is natural.
In this paper we will consider bundles with metrics such that $\psi=\log|\cdot|^{2}$ is quasi-plurisubharmonic (qpsh) with analytic singularities on $E$. The latter condition means that locally $\psi=c\log|f|^{2}+b$, where $f$ is a tuple of holomorphic functions, $c>0$ and $b$ is bounded. We say that the singularities are integral (rational/real) if $c$ is a positive integer (rational/real number). Such a metric, as well as the associated bundle will be called qas (qpsh with analytic singularities). For such metrics, the unbounded locus will be an analytic subset of $X$. We extend the definition of Segre form $s(E,h,h_{0})$ to qas bundles $E\to X$. For simplicity, we will assume that qas metrics have integral singularities, which is practical in order to keep track on arithmetic information. However, almost all the results extend easily to the case of rational or real singularities, see Remark 6.9.
Example 1.1.
Assume that $g\colon E\to F$ is a holomorphic vector bundle morphism and $F$ has a smooth metric, or a qas metric with empty unbounded locus. Then the metric $|\alpha|^{2}=|g(x)\alpha|^{2}$ on $E$ inherited from $F$ is qas, see Corollary 3.4. ∎
Example 1.2.
An example of a singular metric $e^{\psi}$ on a line bundle $L$ is given by $\psi=\log(|\sigma_{1}|^{2}+\cdots+|\sigma_{m}|^{2})$, where $\sigma_{j}$ are sections of $L^{-1}$. It is covered by Example 1.1 via the mapping $L\to X\times\oplus_{1}^{m}{\mathbb{C}}$, defined by $(\sigma_{1},\ldots,\sigma_{m})$. Such metrics are important in modern complex algebraic geometry. ∎
By a smooth metric $h$ on $E$, we mean (as usual) that with respect to a local trivialization, the matrix representing $h$ has entries which are smooth functions, and $h$ is strictly positive definite. We begin by considering the basic case of qas metrics with empty unbounded locus, i.e., which are still strictly positive definite, bot not necessarily smooth. Then we have the following result.
Proposition 1.3.
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$, and assume that the unbounded locus $Z$ of $h$ is empty. Then the Segre form $s(E,h,h_{0})$ associated to $h$ is independent of the choice of smooth reference metric $h_{0}$, and may hence be denoted $s(E,h)$. Furthermore, there exists locally a sequence of smooth metrics $h_{\epsilon}$ decreasing to $h$ such that
(1.3) $\lim_{{\epsilon}\to 0}s(E,h_{\epsilon})=s(E,h).$
If $h$ is additionally Griffiths negative, then this is a special case of the results in [17].
We consider now a general qas metric $h$ on $E$. Then, the metric $h_{\epsilon}$ defined by (1.1) is a qas metric with empty unbounded locus, cf. Lemma 3.5 (iv). Hence, $s(E,h_{\epsilon})$ is well-defined by the above proposition, independent of any choice of smooth reference metric.
The components $s_{k}(E,h,h_{0})$ of the Segre form are closed currents of order 00, but not necessarily positive unless both $h$ and $h_{0}$ are Griffiths negative. There are anyway well-defined multiplicities $\operatorname{mult}_{x}s_{k}(E,h,h_{0})$ at each point $x$, defined in the same way as Lelong numbers, see Section 4, and they turn out to be non-negative integers. Here is our first main theorem:
Theorem 1.4.
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$, and let $Z$ be the unbounded locus of $h$. Let $h_{0}$ be a smooth reference metric, and let $s(E,h,h_{0})$ denote the associated Segre form.
(i) If $h_{\epsilon}$ is the metric defined by (1.1), then
$s(E,h_{\epsilon})\to s(E,h,h_{0}).$
(ii) There is a current $w$ on $X$ such that
(1.4) $dd^{c}w=s(E,h,h_{0})-s(E,h_{0}).$
(iii) We have the decomposition
(1.5) $s(E,h,h_{0})=s^{\prime}(E,h)+M^{E,h,h_{0}},$
where $M^{E,h,h_{0}}$ has support on $Z$ and $s^{\prime}(E,h)={\bf 1}_{X\setminus Z}s(E,h,h_{0})$ is independent of $h_{0}$ and if $h$ is smooth on $X\setminus Z$, $s^{\prime}(E,h)$ coincides with the usual Segre form of $h$ there.
(iv) If $\pi\colon X^{\prime}\to X$ is a modification, then $\pi^{*}h$ is a qas metric on $\pi^{*}E$, and
(1.6) $\pi_{*}s(\pi^{*}E,\pi^{*}h,\pi^{*}h_{0})=s(E,h,h_{0}).$
(v) If $E^{\prime}\to X$ is another holomorphic vector bundle, equipped with a smooth metric $h^{\prime}$, then
(1.7) $s(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})=s(E^{\prime},h^{\prime})\wedge s(E,h,h_{0}).$
(vi) For each integer $k\geq 0$, there is a decomposition
(1.8) $M_{k}^{E,h,h_{0}}=S^{E,h}_{k}+N^{E,h,h_{0}}_{k},$
where $S_{k}^{E,h}$ is the Lelong current of an analytic cycle of codimension $k$, and the non-vanishing multiplicities of $s_{k}^{\prime}$ and $N_{k}$ are contained in an analytic set with codimension $\geq k+1$. The multiplicities of $s_{k}^{\prime}$, $S_{k}$ and $N_{k}$ are non-negative integers, independent of the choice of $h_{0}$.
(vii) If $h^{\prime}$ is a metric on $E\to X$ comparable to $h$, and $h_{0}^{\prime}$ is a smooth metric on $E$, then $\operatorname{mult}_{x}M_{k}^{E,h,h_{0}}=\operatorname{mult}_{x}M_{k}^{E,h^{\prime},h_{0}^{\prime}}$ for all $x\in X$ and $k\geq 0$, and similarly if $M_{k}$ is replaced by $s_{k}^{\prime}$. In particular $S_{k}^{E,h}=S_{k}^{E,h^{\prime}}$, $k=0,1,2,\ldots$.
Remark 1.5.
(i) Equation (1.4) means that $s(E,h,h_{0})$ represents the Bott-Chern class of $E$.
(ii) The properties (ii) to (iv) are (at least implicitly) already proved in [18].
(iii) Equality (1.7) implies that
$s^{\prime}(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})$ $=s(E^{\prime},h^{\prime})\wedge s^{\prime}(E,h,h_{0})\text{ and }$
$M^{E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0}}$ $=s(E^{\prime},h^{\prime})\wedge M^{E,h,h_{0}}.$
(iv) Notice that if $M_{k}^{E,h,h_{0}}$ is a positive current, (1.8) is precisely the Siu decomposition. ∎
In [9], there is a general definition of singular metrics of vector bundles, and a definition of Griffiths negative metrics as mentioned above, and a definition of Griffiths positive metrics, which by definition are dual metrics of Griffiths negative metrics (which coincides with the standard definition when the metric is smooth). In e.g. the references mentioned above concerning definitions of Chern and Segre forms of singular metrics, the results have mainly been formulated for Griffiths positive metrics, which is natural as in the smooth case, Griffiths positive metrics seem to be more frequently of interest than Griffiths negative metrics. In this article, we consider metrics which are generalizations of Griffiths negative metrics instead of Griffiths positive metrics. However, by considering duals of such metrics, one obtains corresponding statements concerning Griffiths positive metrics, cf. Remark 6.2. We opted here for the formulation in the negatively curved case, as it leads to simpler formulations, in particular with less signs and dual metrics.
Recall that in algebraic geometry, characteristic classes of vector bundles such as the Chern class, are defined as endomorphisms on Chow groups, i.e., the ${\mathbb{Z}}$-module of analytic cycles ${\mathcal{Z}}(X)$ modulo rational equivalence. In some cases, as for instance in case of a Hermitian holomorphic vector bundle, there is a canonical representative of its Segre and Chern classes.
Inspired by [5] we introduce a ${\mathbb{Z}}$-module ${\mathcal{QZ}}(X)$ of currents of order 00 that contains the module ${\mathcal{Z}}(X)$ of analytic cycles as a submodule, if cycles are identified with their associated Lelong currents. In particular the $(0,0)$-current $1$ is an element in ${\mathcal{QZ}}(X)$. Any element ${\mathcal{QZ}}(X)$ of a fixed dimension has a well-defined multiplicity (Lelong number) at each point. We also introduce certain quotient classes ${\mathcal{QB}}(X)$ that, contrary to the Chow classes, keep local information such as multiplicities; this fact is crucial in this paper.
A second main result is that the Segre form $s(E,h,h_{0})$ extends to an endomorphism ${\bf s}(E,h,h_{0})$ on ${\mathcal{QZ}}(X)$, such that ${\bf s}_{0}(E,h,h_{0})$ is the identity and ${\bf s}(E,h,h_{0})1=s(E,h,h_{0})$. An important fact is that ${\bf s}(E,h,h_{0})$ defines a well-defined endomorphism ${\bf s}(E,h)$ of ${\mathcal{QB}}(X)$, independent of $h_{0}$. There is an analogue of Theorem 1.4 for ${\bf s}(E,h,h_{0})$, see Theorem 6.5.
Since ${\bf s}(E,h,h_{0})-I$ has positive degree,
(1.9) ${\bf c}(E,h,h_{0})=\sum_{\ell=0}^{\infty}(-1)^{\ell}({\bf s}(E,h,h_{0})-{I})^{\ell}$
is a well-defined endomorphism, such that
(1.10) ${\bf c}(E,h,h_{0})\wedge{\bf s}(E,h,h_{0})={\bf s}(E,h,h_{0})\wedge{\bf c}(E,h,h_{0})={I}.$
It is natural to say that $c(E,h,h_{0}):={\bf c}(E,h,h_{0})1$ is the Chern form associated with $(E,h,h_{0})$. The same definition of $c(E,h,h_{0})$ appeared in [18] but formulated somewhat differently.
Theorem 1.6.
Let $(E,h,h_{0})$ be as in Theorem 1.4. Properties (ii) to (vii) of Theorem 1.4 hold if $s(E,h,h_{0})$ is replaced by $c(E,h,h_{0})$, except for that the multiplicities are not necessarily non-negative.
As already mentioned, the starting point for this paper is [18] and the ideas there. The regularization (1.1) relies on results in [6, 10]. The arithmetic analysis of the Lelong numbers relies on ideas in [1], and the modules ${\mathcal{QZ}}(X)$ and ${\mathcal{QB}}(X)$ that are modelled on the somewhat smaller module ${\mathcal{GZ}}(X)$ and ${\mathcal{B}}(X)$ in [5].
Acknowledgements
We are grateful to Håkan Samuelsson Kalm, Antonio Trusiani and David Witt Nyström for helpful discussions.
2. Preliminaries
Throughout the article, $X$ is assumed to be a complex manifold. Given a holomorphic vector bundle $E\to X$ we let $p\colon{\mathbb{P}}(E)\to X$ denote the projectivization, so that the fiber above $x$ is the set of all lines [^2] There is another common convention in the literature, where the fiber instead is all hyperplanes through the origin. through the origin in $E_{x}$. Then the pullback $p^{*}E\to{\mathbb{P}}(E)$ has the tautological sub(line)bundle $L_{E}\to{\mathbb{P}}(E)$ whose fiber over $(x,[\alpha])\in{\mathbb{P}}(E)$ is the line in $E_{x}$ determined by $\alpha$.
If $E$ has a smooth Hermitian metric $h$, then $p^{*}E$ and hence $L_{E}$ inherit smooth Hermitian metrics. Notice that a local section $s$ of $L_{E}$ has the form $s=\tilde{s}(x,\alpha)\alpha$, where $\tilde{s}$ is a holomorphic function on $E\setminus 0_{E}$ which is $-1$-homogeneous in $\alpha$. Thus any local section can be identified with such an $\tilde{s}$. The induced norm is
(2.1) $|s|^{2}_{e^{\psi}}:=|\tilde{s}|^{2}|\alpha|^{2}_{h}.$
In a local frame for $L_{E}$, say $1/\alpha_{j}$, the section $s$ is represented by the function $\tilde{s}\alpha_{j}$, and from (2.1) we have that
$|s|^{2}_{e^{\psi}}=|\tilde{s}\alpha_{j}|^{2}e^{\log(|\alpha|_{h}^{2}/|\alpha_{j}|^{2})}.$
Thus the metric $e^{\psi}$ on $L_{E}$ with respect to the local frame $1/\alpha_{j}$ is given by
(2.2) $\psi=\log(|\alpha|^{2}_{h}/|\alpha_{j}|^{2}).$
Note that even though $\log|\alpha|_{h}^{2}$ is not well-defined on ${\mathbb{P}}(E)$, $dd^{c}\log|\alpha|_{h}^{2}$ is a well-defined form on ${\mathbb{P}}(E)$, and if $s$ is a non-vanishing section of $L_{E}$, then $dd^{c}\log|s|_{h}^{2}=dd^{c}\log|\alpha|_{h}^{2}$. Thus, if $c_{1}(L_{E},e^{\psi})$ and $s_{1}(L_{E},e^{\psi})$ are the first Chern form and Segre form, respectively, then
$-c_{1}(L_{E},e^{\psi})=s_{1}(L_{E},e^{\psi})=dd^{c}\log|\alpha|^{2}_{h},$
so that
$s(L_{E},e^{\psi})=\frac{1}{c(L_{E},e^{\psi})}=\frac{1}{1+c_{1}(L_{E},e^{\psi})}=\frac{1}{1-s_{1}(L_{E},e^{\psi})}=\sum_{\ell=0}^{\infty}(dd^{c}\log|\alpha|_{h}^{2})^{\ell}.$
The Segre form of $(E,h)$ is defined as
(2.3) $s(E,h)=p_{*}s(L_{E},e^{\psi})=p_{*}\sum_{\ell=0}^{\infty}(dd^{c}\log|\alpha|_{h}^{2})^{\ell},$
and it is a smooth form since $p$ is a submersion. We write $s(E,h)=\sum s_{k}(E,h)$, where $s_{k}(E,h)$ has bidegree $(k,k)$. It turns out that $s_{0}(E,h)=1$ and thus $s(E,h)-1$ has positive bidegree. We have the Chern form
(2.4) $c(E,h)=\frac{1}{s(E,h)}=\sum_{\ell=0}^{\infty}(-1)^{\ell}(s(E,h)-1)^{\ell}.$
By [19] (see also i.e., [14]), these definitions of $c(E,h)$ and $s(E,h)$ coincide with the more standard definitions of Chern form defined as the sum of the elementary symmetric polynomials acting on ($i/2\pi$ times) the curvature tensor $\Theta(h)=D^{2}$ of $(E,h)$, where $D$ denotes the Chern connection with respect to $h$, and the Segre form defined as $s(E,h)=1/c(E,h)$.
2.1. Qpsh functions with analytic singularities
Let $Y$ be a complex manifold. Recall that a real valued function $u$ is qpsh (quasi-plurisubharmonic) if locally $u=v+a$, where $v$ is psh and $a$ is smooth. As already mentioned, $u$ has analytic singularities if locally
(2.5) $u=c\log|f|^{2}+b,$
where $b$ is bounded.
Lemma 2.1.
Assume that $W$ is a connected complex manifold. If $\tau\colon W\to Y$ is holomorphic and $u$ is qpsh with (integral) analytic singularities on $Y$, then $\tau^{*}u$ is qpsh with (integral) analytic singularities in $W$. Possibly $\tau^{*}u$ is identically $\infty$- ∞.
Lemma 2.2.
If $\psi_{1}$ and $\psi_{2}$ are qpsh (with integral analytic singularities), then $\psi_{1}+\psi_{2}$ and $\log(e^{\psi_{1}}+e^{\psi_{2}})$ are qpsh (with integral analytic singularities).
Lemma 2.1 follows directly from the definitions. Lemma 2.2 is well-known for psh $\psi_{j}$ (with integral analytic singularities). It is readily checked that it also holds for qpsh (with integral analytic singularities), using the characterization that functions $\psi_{j},\ j=1,2,$ are qpsh (with integral analytic singularities) if and only if locally there is a smooth function $\phi$ such that $\psi_{j}^{\prime}:=\psi_{j}+\phi,\ j=1,2,$ are psh (with integral analytic singularities).
2.2. A regularization of locally bounded qpsh functions
In this subsection, we will consider a modification of the standard local regularization of psh functions through convolution, which will be suitable for approximating qas metrics with empty unbounded locus.
Proposition 2.3.
Let $\psi(x,y)$ be a locally bounded qpsh function on $U\subset{\mathbb{C}}^{n}_{x}\times{\mathbb{C}}^{m}_{y}$, assume that $\psi$ is smooth as a function of $y$, with $x$ fixed, and let $v(x,y)$ be a smooth psh function. Let $\rho_{\epsilon}(x)=\rho(x/{\epsilon})/{\epsilon}^{n}$ be a family of radial mollifiers with respect to the variable $x$. Define
(2.6) $\psi_{\epsilon}=\log(e^{\psi}*_{x}\rho_{\epsilon}+A{\epsilon}e^{v}),$
where $\psi*_{x}\rho_{\epsilon}$ denotes the convolution with respect to $x$. For any $V\subset\subset U$, $\psi_{\epsilon}$ is smooth on $V$ for ${\epsilon}$ small enough, and if $A\gg 1$, then $\psi_{\epsilon}$ is increasing in ${\epsilon}$. If $\psi=u+g$, where $u$ is psh and $g$ is smooth, and $dd^{c}g\geq-B\omega$, where $\omega=dd^{c}(|x|^{2}+|y|^{2})$, then $dd^{c}\psi_{\epsilon}\geq-B\omega$.
By decomposing $dd^{c}\psi_{\epsilon}=(dd^{c}(\psi_{\epsilon}+B(|x|^{2}+|y|^{2}))-B\omega$, and applying the Bedford-Taylor theory of convergence of Monge-Ampère products for locally bounded psh functions, one obtains the following convergence.
Corollary 2.4.
Let $\psi$ and $\psi_{\epsilon}$ be as in Proposition 2.3. Then, for any $k\geq 0$,
(2.7) $(dd^{c}\psi_{\epsilon})^{k}\to(dd^{c}\psi)^{k}.$
Proof of Proposition 2.3.
We first prove that if $\psi=u$ is psh, then $\psi_{\epsilon}$ is also psh. We begin by assuming that $u$ is also smooth. In that case, by approximating the convolution by Riemann sums, $u_{\epsilon}$ may be uniformly approximated by $u_{{\epsilon},N}=\log(\sum_{k}e^{u_{k}+\log c_{k}}+e^{v+\log(A{\epsilon})})$, where $c_{k}$ are positive constants, and $u_{k}(x,y)=u(x-x_{k},y)$. Since all the functions in the exponents are psh, $u_{{\epsilon},N}$ is also psh, and thus so is $u_{\epsilon}$. If $u$ is psh, we may choose a sequence of smooth functions $u_{k}$ decreasing to $u$, and then $(u_{k})_{\epsilon}$ decreases to $u_{\epsilon}$, and hence $u_{\epsilon}$ is psh.
We consider now the case when $\psi=u+g$, where $g$ is smooth, and show that $dd^{c}\psi_{\epsilon}\geq-B\omega$, or equivalently, that $\psi_{\epsilon}^{\prime}=\psi_{\epsilon}+B(|x|^{2}+|y|^{2})$ is psh. As above, it suffices to prove the statement assuming that $u$ is smooth. We may write
$\psi_{\epsilon}^{\prime}(x,y)=\log((e^{\psi}*_{x}\rho_{\epsilon})e^{B(|x|^{2}+|y|^{2})}+e^{v+\log(A{\epsilon})+B(|x|^{2}+|y|^{2}}),$
and as above, $\psi_{\epsilon}^{\prime}$ may be uniformly approximated by
$\psi_{{\epsilon},N}^{\prime}=\log(\sum e^{\psi_{k}+\log c_{k}+B(|x|^{2}+|y|^{2})}+e^{v+\log(A{\epsilon})}).$
Since $\psi_{k}$ is just a translation of $\psi$, and $dd^{c}(|x|^{2}+|y|^{2})$ is translation invariant, it follows that all the functions in the exponents are psh, and thus $\psi_{{\epsilon},N}^{\prime}$ is psh, and hence so is $\psi_{\epsilon}^{\prime}$.
Note that $\psi_{\epsilon}$ is smooth as a function of $x$ and $y$. It remains to prove that it is increasing in ${\epsilon}$ on $V$, for $A\gg 1$. To see this, we may fix $y$ and consider everything as a function of $x$. We write $\psi=u+g$, where $u$ is psh and $g$ is smooth on $X$. We then decompose $e^{\psi_{\epsilon}}$ as
[…]
Since $e^{u}$ is psh, the first term of the right hand side is increasing in $\epsilon$. It then suffices to prove that the middle term is differentiable in $\epsilon$, with a derivative less than $Ae^{v}$ to see that the whole expression is increasing in $\epsilon$. To see this, note that $(d/d\epsilon)\rho_{\epsilon}(x^{\prime})={\mathcal{O}}(1/\epsilon^{n+1})$, and has support where $|x^{\prime}|\leq\epsilon$. Since $|e^{g(x-x^{\prime})}-e^{g(x)}|\leq C\epsilon$ for $|x^{\prime}|\leq\epsilon$ (with $C$ depending on the $C^{1}$-norm of $e^{g}$), we obtain that the derivative of the middle term is bounded by a constant, which is smaller than $Ae^{v}$ on $V$ if $A$ is chosen large enough. ∎
3. qas bundles
Let $E\to X$ be a holomorphic vector bundle. As already mention in the introduction we have the following definition:
Definition 3.1.
We say that a real function is qas if it is qpsh with integral analytic singularities. A singular metric $h$ on $E\to X$ is qas if the function $\alpha\mapsto\log|\alpha|^{2}_{h}$ is qas on the total space $E$. In that case $(E,h)$ is said to be qas.
We will often drop the metric from the notation, and simply say that $E$ is a qas bundle. We will then also often denote the norm simply by $|\alpha|$.
Example 3.2.
If $L$ is a line bundle with metric $e^{\psi}$, then $L$ is qas precisely if the representation of $\psi$ in any local frame is qas. ∎
Lemma 3.3.
Let $E$ be a holomorphic vector bundle on $X$ with a smooth Hermitian metric $h$. Any point $x_{0}\in X$ has a neighborhood $U$ such that $h^{\prime}=he^{A|x|^{2}}$ is Griffiths negative if $A$ is large enough, where $x$ denotes a local coordinate system near $x_{0}$.
Proof.
If we decompose the curvature $\Theta_{h}$ of $h$ as $\sum\Theta_{jk}dx_{j}{\wedge}d\bar{x}_{k}$, then $h$ is Griffiths negative if $\sum(\Theta_{jk}s,s)v_{j}\bar{v}_{k}\leq 0$ for any section $s$ of $E$ and any tangent vector $v=\sum v_{j}\frac{\partial}{\partial x_{j}}$. Notice that $\Theta_{h^{\prime}}=\Theta_{h}-A\omega_{x}$, where $\omega_{x}$ is the standard Kähler form on $U\subset{\mathbb{C}}^{N}_{x}$. By possibly shrinking $U$, we may choose $A$ such that $|\sum(\Theta_{jk}s,s)v_{j}\bar{v}_{k}|\leq A|s|^{2}|v|^{2}$, which implies that $h^{\prime}$ is Griffiths negative. ∎
The statement means that locally there is a negative line bundle $L_{A}\to U$ such that $E\otimes L_{A}$ is Griffiths negative. It is well-known that if $(E,h)$ is Griffiths negative, then $\log|\alpha|^{2}_{h}$ is psh on $E$, see, e.g., [9]. Thus we have
Corollary 3.4.
Any vector bundle with a smooth Hermitian metric is qas.
Lemma 3.5.
(i) If $\tau\colon W\to X$ is holomorphic and $E\to X$ is qas, then $\tau^{*}E\to W$ is qas.
(ii) If $g\colon E\to F$ is a bundle morphism and $F$ is qas, then $E$ is qas if equipped with the induced metric $|\alpha|:=|g(x)\alpha|$.
(iii) If $E_{j}\to X$, $j=1,2$, are qas, then $E_{1}\oplus E_{2}\to X$ is qas.
(iv) If $E\to X$ has two qas metrics $h_{1}$ and $h_{2}$, then $h_{1}+h_{2}$ is qas. That is, $|\alpha|^{2}=|\alpha|^{2}_{h_{1}}+|\alpha|^{2}_{h_{2}}$ is qas.
(v) If $L$ is a qas line bundle, and $E$ is a qas vector bundle, then $E\otimes L$ is qas.
Proof.
If $\tau\colon W\to X$, then $|\tau^{*}\alpha|^{2}_{\tau^{*}h}=\tau^{*}|\alpha|^{2}_{h}$ so (i) follows from Lemma 2.1. To see (ii), consider the holomorphic mapping $E\to F$, $(x,\alpha)\mapsto(x,g(x)\alpha)$. Since the pullback under a holomorphic mapping preserves the property qas, (ii) follows.
Locally $E_{j}=X\times{\mathbb{C}}_{\alpha_{j}}^{r_{j}}$. If $\log|\alpha_{j}|^{2}$ is qas on $E_{j}$, for $j=1,2$, then so is its pullback to $E_{1}\oplus E_{2}$. Thus $\log|\alpha_{j}|^{2}$ is qas on $E_{1}\oplus E_{2}$, so that by Lemma 2.2, $\log(|\alpha_{1}|^{2}+|\alpha_{2}|^{2})$ is qas on $E_{1}\oplus E_{2}$. We get (iv) by applying (i) and (iii) to $(E,h_{1})$, $(E,h_{2})$ and the mapping $E\to E\oplus E;\ \ \alpha\mapsto(\alpha,\alpha).$ For (v), note that if $e_{L}$ is a local frame for $L$, and $e^{\psi_{L}}$ the corresponding representation of the metric, then $\log|\alpha\otimes e_{L}|^{2}=\log|\alpha|^{2}+\psi_{L}$, so (v) follows by Lemma 2.2. ∎
Remark 3.6.
We say that a function is ${\mathbb{Q}}$-qas or ${\mathbb{R}}$-qas if it is qpsh with analytic singularities, and the local constants $c$ in (2.5) are in ${\mathbb{Q}}_{\geq 0}$ or ${\mathbb{R}}_{\geq 0}$, and that a vector bundle is ${\mathbb{Q}}$-qas or ${\mathbb{R}}$-qas if $\alpha\mapsto\log|\alpha|^{2}$ is of the corresponding class.
Note that Lemma 2.1 holds for ${\mathbb{Q}}$-qas and ${\mathbb{R}}$-qas functions, and consequently Lemma 3.5 (i)-(ii) holds for ${\mathbb{Q}}$-qas and ${\mathbb{R}}$-qas metrics. We also have that Lemma 2.2 holds for ${\mathbb{Q}}$-qas functions, but not necessarily for ${\mathbb{R}}$-qas functions (since $c_{1}\log|f_{1}|^{2}+c_{2}\log|f_{2}|^{2}$ and $\log(|f_{1}|^{2c_{1}}+|f_{2}|^{2c_{2}})$ do not necessarily have analytic singularities if $c_{1}$ and $c_{2}$ are linearly independent over ${\mathbb{Q}}$). Consequently, the corresponding version of Lemma 3.5 (iii)-(v) holds for ${\mathbb{Q}}$-qas metrics, but not necessarily for ${\mathbb{R}}$-qas metrics. ∎
Lemma 3.7.
For any $x_{0}\in X$, there is a neighborhood $U\subset X$ and holomorphic $F_{k}(x,\alpha)=f_{k}(x)\cdot\alpha$ in $E|_{U}\simeq U\times{\mathbb{C}}^{r}_{\alpha}$, $k=1,\ldots,m$, such that
(3.1) $\sum_{k=1}^{m}|F_{k}(x,\alpha)|\sim|\alpha|_{h(x)}.$
Proof.
Assuming say that $x_{0}=0\in{\mathbb{C}}^{N}$ in a local embedding, by the assumption of analytic singularities, there are holomorphic functions $F_{k}(x,\alpha)$ in a set $|x|\leq c,|\alpha|\leq c,\ c>0$ such that (3.1) holds. Since $F_{k}(x,0)=0$ we have that $F_{k}(x,\alpha)=f_{k}(x)\cdot\alpha+F^{\prime}_{k}$, where $F^{\prime}_{k}$ is ${\mathcal{O}}(|\alpha|_{0}^{2})$ (where $|\alpha|_{0}^{2}$ denotes the trivial metric on ${\mathbb{C}}^{r}_{\alpha}$). We write the tuple as $F=f(x)\cdot\alpha+F^{\prime}$. For $x\leq c$, $|\alpha|=c$ and $|\lambda|\leq 1$, we have
$F(x,\lambda\alpha)=\lambda f(x)\cdot\alpha+{\mathcal{O}}(|\lambda|^{2}).$
Thus
$|\lambda||\alpha|_{h(x)}\leq C|F(x,\lambda\alpha)|\leq C|\lambda||f(x)\cdot\alpha|+C|\lambda|^{2},$
so that
$|\alpha|_{h(x)}\leq C|f(x)\cdot\alpha|,\quad|\alpha|=c,\ |x|\leq c.$
By homogeneity this holds for all $\alpha$. In the same way we get the opposite estimate. Thus (3.1) holds with $f(x)\cdot\alpha$ instead of $F(x,\alpha)$. However, $f(x)\cdot\alpha$ is defined in the whole set $U\times{\mathbb{C}}^{r}$ and (3.1) holds there by homogeneity. ∎
Thus at least locally in $x$, $|\cdot|_{h}^{2}$ is comparable with a metric as in Example 1.1.
Proposition 3.8.
If $(E,h)$ is qas, then $(L_{E},e^{\psi})$ is qas.
Proof.
Indeed, by cf. (2.2), $\psi$ is qas on ${\mathbb{P}}(E)$ if $\log|\alpha|_{h}^{2}$ is qas on $E$, or in fact even on $E\setminus 0_{E}$. ∎
Notice by the way that $E$ is Griffiths negative, i.e., $\psi$ is psh, if and only if $e^{\psi}$ is a negatively curved metric.
Remark 3.9.
Almost all results in this paper only depend on that $L_{E}$ is qas. At a few places, Lemma 3.5 (ii)-(iii) and Lemma 3.7, and the part of Theorem 1.4 that the multiplicities are non-negative, we need the (possibly) stronger assumption that $E$ is qas. The remainder of the results in the article which are formulated for qas vector bundles $E$ indeed hold if $L_{E}$ is assumed to be qas.
In [18], the standing assumption is that $E$ is equipped with a metric $h$ such that the induced metric on $L_{E}$ is psh with analytic singularities. Note that by [18, Proposition 5.4], this is equivalent to $\log|\alpha|_{h}^{2}$ being psh with analytic singularities on $E\setminus 0_{E}$. The same arguments yield the corresponding statement when psh is replaced by qpsh, i.e., for qas metrics and functions. For such metrics, similarly to [18, Proposition 5.2], one may show that if the metric on $L_{E}$ is qas, then $\log|\alpha|_{h}^{2}$ is in fact qpsh on all of $E$. However we do not know if $\log|\alpha|_{h}^{2}$ has analytic singularities over $0_{E}$, so it is indeed possibly a stronger assumption that $E$ is qas.
The results mentioned above where we need to assume that $E$ is qas do not appear in [18]. However, if $L_{E}$ has a psh metric with analytic singularities, then the Lelong numbers of Segre forms are non-negative. Indeed, since the Lelong numbers do not depend on the choice of reference metric, we may may locally choose a negatively curved reference metric, and then $s_{k}(E,h,h_{0})$ is a positive $(k,k)$-current, and hence it has non-negative Lelong numbers.
If the following generalization of Lemma 3.3 to the case when $L_{E}$ is qas holds, then it would imply that $E$ may be equipped with a singular metric $h^{\prime}$ comparable to $h$ which is such that induced metric on $L_{E}$ is psh with analytic singularities, and in particular, $s_{k}(E,h,h_{0})$ would have non-negative Lelong numbers also under this assumption. The generalization of Lemma 3.3 is as follows: If $E$ is equipped with a singular metric $h$ such that $L_{E}$ is qas, then $he^{A|x|^{2}}$ is a Griffiths negative singular metric in a neighborhood of $x_{0}$ for $A$ large enough, i.e., the induced metric on $L_{E}$ is psh. We do not know whether it holds. ∎
4. Generalized cycles and quasi-cycles
Let $Y$ be a complex manifold or a reduced analytic space. In [5] was introduced the ${\mathbb{Z}}$-module ${\mathcal{GZ}}(Y)$ of generalized cycles, which are currents on $Y$, and which is an extension of the module ${\mathcal{Z}}(Y)$ of analytic cycles, that is, there is an injective mapping $i\colon{\mathcal{Z}}(Y)\to{\mathcal{GZ}}(Y)$, if the elements in ${\mathcal{Z}}(Y)$ are identified with their associated Lelong currents. It was shown that many properties of ${\mathcal{Z}}(Y)$ such as unique decomposition into irreducible components, well-defined integer multiplicities at each point etc., hold for ${\mathcal{GZ}}(Y)$ as well.
Recall that if $Y$ is a reduced analytic space, then smooth forms on $Y$ are defined in terms of restrictions of smooth forms in local embeddings, cf. e.g. [2], and currents as the appropriate dual of smooth compactly supported forms.
Let ${\mathcal{QZ}}(Y)$ be the $\mathbb{Z}$-module of locally finite sums of currents of the form $\tau_{*}\gamma$, where $\tau\colon W\to Y$ is a proper holomorphic mapping, $W$ is a complex manifold, and $\gamma$ is a current on $W$ of the form
(4.1) $\gamma=dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{p},$
where $e^{b_{j}}$ are metrics on holomorphic line bundles $L_{j}$ over $W$ such that $b_{j}$ are locally bounded and qpsh. We say that the elements in ${\mathcal{QZ}}(Y)$ are quasi-cycles on $Y$.
Remark 4.1.
If we assume all $b_{j}$ are smooth we get precisely the module ${\mathcal{GZ}}(Y)$, cf.  [5, Remark 3.1], which is thus a submodule of ${\mathcal{QZ}}(Y)$. ∎
Remark 4.2.
One can just as well consider the corresponding ${\mathbb{Q}}$-module or ${\mathbb{R}}$-module that we denote by ${\mathcal{QZ}}_{\mathbb{Q}}(Y)$ and ${\mathcal{QZ}}_{\mathbb{R}}(Y)$, respectively. It is natural to refer to the elements as rational and real quasi-cycles, respectively. ∎
The product (4.1), defined by Bedford and Taylor, is a closed current of order 00. It follows that any $\mu\in{\mathcal{QZ}}(Y)$ is a closed current of order 00. In particular, such currents may be multiplied with characteristic functions ${\bf 1}_{V}$, if $V$ is an analytic subset of $W$. If $\tau\colon Y^{\prime}\to Y$ is a proper map, and $\mu\in{\mathcal{QZ}}(Y^{\prime})$, then $\tau_{*}\mu\in{\mathcal{QZ}}(Y)$.
Most results in ${\mathcal{GZ}}(Y)$ from [5] extend to ${\mathcal{QZ}}(Y)$, with the same arguments, because of the following three properties of mixed Monge-Ampère products of the form (4.1):
Lemma 4.3.
Let $e^{b_{1}},\ldots,e^{b_{k}}$ be metrics on line bundles $L_{1},\dots,L_{k}$ on a complex manifold $W$ such that $b_{1},\dots,b_{k}$ are locally bounded qpsh.
(i) If $i\colon V\to W$ is an analytic subspace of positive codimension, then
(4.2) ${\bf 1}_{V}dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{k}=0.$
(ii) If $i\colon V\to W$ is as in (ii), then
(4.3) $dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{k}\wedge[V]=i_{*}(dd^{c}i^{*}b_{1}\wedge\cdots\wedge dd^{c}i^{*}b_{k}).$
(iii) If $W$ is compact and $k=\dim W$, then
[…]
Here $c_{1}(L_{j})$ denotes the Chern class of $L_{j}$, which can be represented by the first Chern forms associated with any smooth Hermitian metric on $L_{j}$.
Proof.
It is well-known from the theory of Monge-Ampère products due to Bedford-Taylor, [8], that such products do not charge pluripolar sets, cf. also, e.g., [13, Proposition III.3.11], so in particular (4.2) holds. The equality (4.3) is a local statement and follows by a suitable regularization, see, [4, Lemma 3.4]. The last equality, (4.4), is well-known in case all $b_{j}$ are smooth. Here is a sketch of an argument in our case. If $\beta_{j}$ are smooth metrics, then $b_{j}-\beta_{j}$ are global functions. We prove by induction that (4.4) holds if $b_{1},\ldots,b_{\ell}$ are replaced by $\beta_{1},\ldots\beta_{\ell}$ for $1\leq\ell\leq k$. Assume it is true for $\ell-1$. Then by the commutativity of Monge-Ampère products it is enough to verify, with $u=b_{\ell}-\beta_{\ell}$, that
$dd^{c}(uT)=dd^{c}b_{\ell}\wedge T-dd^{c}\beta_{\ell}\wedge T$
if $T$ is any positive closed current. This is a local statement and follows from [13, Proposition III.4.9] since locally $u=\hat{b}_{\ell}-\hat{\beta}_{\ell}$, where $\hat{b}_{\ell}$ is psh and bounded, and $\hat{\beta}_{\ell}$ is smooth. Now the induction step follows from Stokes’ theorem. ∎
Using (4.2), it follows as in [5, Lemma 3.2] that if $V\hookrightarrow Y$ is an analytic subspace, and $\mu\in{\mathcal{QZ}}(Y)$, then ${\bf 1}_{V}\mu\in{\mathcal{QZ}}(Y)$. Indeed, if
(4.5) $\mu=\sum(\tau_{j})_{*}\gamma_{j},$
where $\tau_{j}:W_{j}\to Y$, then
(4.6) ${\bf 1}_{V}\mu=\sum_{\gamma_{j}(W_{j})\subset V}(\tau_{j})_{*}\gamma_{j}.$
Proposition 4.4.
Assume that $i\colon V\to Y$ is an analytic subspace. Then the image of $i_{*}\colon{\mathcal{QZ}}(V)\to{\mathcal{QZ}}(Y)$ is precisely the elements in ${\mathcal{QZ}}(Y)$ that have Zariski support on $V$.
The proof is precisely the same as for the analogous statement for ${\mathcal{GZ}}(Y)$, [5, Proposition 3.6].
Example 4.5.
If $\mu\in{\mathcal{QZ}}(Y)$ has support at $x\in Y$, then since it has order 00 it must be a number times $[x]$. In view of Proposition 4.4 this number is a finite sum of numbers $\alpha=\int\tau_{*}\gamma$, where $\tau(W)=\{x\}$. Now
$\alpha=\int_{Y}\tau_{*}\gamma=\int_{W}\gamma$
which is an intersection number, cf. (4.4), and hence an integer. ∎
4.1. Monge-Ampère type products as homomorphisms on ${\mathcal{QZ}}(Y)$
Let $u$ be a qas function on the complex manifold $Y$. Let $\pi\colon Y^{\prime}\to Y$ be a modification which principalizes the ideals associated to $\pi^{*}u$, i.e., such that in any local representation (2.5), $f$ defines a principal ideal. Such a modification indeed exists. In the case of integral analytic singularities, note first that we may always choose $c=1$ in (2.5). If we then have two different such representations with $f$ and $f^{\prime}$, then $\log(|f^{\prime}|^{2}/|f|^{2})$ is locally bounded, i.e., $|f^{\prime}|\sim|f|$. If we let ${\mathcal{J}}$ denote the integral closure of the ideal generated by $f$, we thus obtain a well-defined coherent ideal sheaf on $Y$. By blowing up ${\mathcal{J}}$ (followed by a resolution of singularities if one wants $Y^{\prime}$ to be smooth), one obtains such a modification. For the case when the analytic singularities are rational or real, see Remark 4.11.
We may thus assume that there is a section $\sigma$ of a holomorphic line bundle ${\mathcal{L}}\to Y^{\prime}$ such that if we equip ${\mathcal{L}}$ with some Hermitian norm, then
(4.7) $\pi^{*}u=\log|\sigma|^{2}+a,$
where $a$ is qpsh and locally bounded. By the Poincaré-Lelong formula, $dd^{c}\log|\sigma|^{2}=[\operatorname{div}\sigma]-c_{1}({\mathcal{L}})$ and hence
(4.8) $dd^{c}\pi^{*}u=[\operatorname{div}\sigma]+\nu,$
where $\nu=-c_{1}({\mathcal{L}})+a$ is a global current on $Y^{\prime}$ of the type that occurs in (4.1).
Remark 4.6.
Note that $\nu$ does not depend on the choice of Hermitian norm on ${\mathcal{L}}$. One can in fact incorporate $a$ into the norm on ${\mathcal{L}}$, i.e., choose a norm $\|\cdot\|^{2}$ on ${\mathcal{L}}$ (with empty unbounded locus) so that $\pi^{*}u=\log\|\sigma\|^{2}$. ∎
Lemma 4.7.
Assume that $\gamma$ is a product on $W$ as in (4.1) and $f$ is a non-trivial holomorphic function. Then $(\log|f|^{2})\gamma$ has locally finite mass, and $dd^{c}((\log|f|^{2})\gamma)=[\operatorname{div}f]\wedge\gamma.$ If $u_{j}$ are smooth and decrease to $\log|f|^{2}$, then $u_{j}\gamma\to(\log|f|)\gamma$.
Proof.
The first statement follows from [13, Theorem III.4.5] since locally in $Y^{\prime}$ all $b_{j}$ in $\gamma$ are $b_{j}=b_{j}^{\prime}+a_{j}$ where $b_{j}^{\prime}$ are psh and bounded and $a_{j}$ are smooth. The second and third one follow from [13, Proposition III.4.9]. ∎
Proposition 4.8.
Let $u$ be a qas function on $Y$ with singular set $Z$. If $\mu\in{\mathcal{QZ}}(Y)$, then $u{\bf 1}_{Y\setminus Z}\mu$ is a measure with locally finite mass and
$[dd^{c}u]\wedge\mu:=dd^{c}(u{\bf 1}_{Y\setminus Z}\mu)$
is an element in ${\mathcal{QZ}}(Y)$.
Proof.
By linearity we can assume that $\mu=\tau_{*}\gamma$,  $\tau\colon W\to Y$. If the image of $\tau$ is contained in $Z$, then ${\bf 1}_{Y\setminus Z}\mu=0$. Otherwise $\tau^{*}u\not\equiv 0$ is qas on $W$ by Lemma 2.1. Then
${\bf 1}_{Y\setminus Z}\mu=\tau_{*}({\bf 1}_{W\setminus\tau^{-1}Z}\gamma)=\tau_{*}\gamma=\mu$
since $\tau^{-1}Z$ has positive codimension on $W$, cf. (4.2). In view of the discussion above we may assume, after possibly performing a modification of $W$, that $\tau^{*}u=\log|\sigma|^{2}+a$ on $W$, cf. (4.7), where $a$ is locally bounded.
Since $u$ is upper semi-continuous there is a sequence $u_{j}$ of smooth functions that decreases to $u$. By Lemma 4.7, $\tau^{*}u_{j}\gamma$ tends to the locally finite measure $\tau^{*}u\gamma$. It follows that
$u\mu=\lim_{j}u_{j}\mu=\lim\tau_{*}(\tau^{*}u_{j}\gamma)=\tau_{*}(\tau^{*}u\gamma)$
is a locally finite measure. Moreover, $dd^{c}(u\mu)=\tau_{*}(dd^{c}(\tau^{*}u\gamma))$ so it is enough to see that $dd^{c}(\tau^{*}u\gamma)$ is in ${\mathcal{QZ}}(W)$.
By (4.8) on $W$, and Lemma 4.7,
(4.9) $dd^{c}(\tau^{*}u\gamma)=[\operatorname{div}\sigma]\wedge\gamma+\nu\wedge\gamma.$
The last term in (4.9) is of the form (4.1). Notice that $[\operatorname{div}\sigma]$ is a sum of terms $m_{\ell}[D_{\ell}]$, where $m_{\ell}$ are integers and $D_{\ell}$ are divisors in $W$. In view of Lemma 4.3 (ii), the first term on the right hand side of (4.9) is therefore a finite sum of terms $m_{\ell}(i_{\ell})_{*}(i_{\ell}^{*}\gamma)$, where $i_{\ell}\colon D_{\ell}\to W$, and each of them is in ${\mathcal{QZ}}(W)$. ∎
Recursively we define
(4.10) $[dd^{c}u]^{0}\wedge\mu=\mu,\quad[dd^{c}u]^{k+1}\wedge\mu=[dd^{c}u]\wedge[dd^{c}u]^{k}\wedge\mu,\quad k=0,1,2,....$
If we let [^3] If $u=\log|\sigma|^{2}$, then $M^{u}$ is denoted by $M^{\sigma}$ in, e.g., [5]. We hope this will not cause any confusion.
(4.11) $\langle dd^{c}u\rangle^{k}\wedge\mu:={\bf 1}_{X\setminus Z}[dd^{c}u]^{k}\wedge\mu,\quad M^{u}_{k}\wedge\mu:={\bf 1}_{Z}[dd^{c}u]^{k}\wedge\mu,$
we see that, e.g.,
(4.12) $[dd^{c}u]^{k+1}\wedge\mu=dd^{c}(u\langle dd^{c}u\rangle^{k}\wedge\mu).$
Let $M^{u}=M^{u}_{0}+M^{u}_{1}+\cdots$. Notice that if $\mu$ has support on $Z$, then $M^{u}_{0}\wedge\mu=\mu$.
Example 4.9.
With the notation in (4.9) in the proof of Proposition 4.8 we have
(4.13) $[dd^{c}\tau^{*}u]^{k}\wedge\gamma=M_{k}^{\tau^{*}u}\wedge\gamma+\langle dd^{c}\tau^{*}u\rangle^{k}\wedge\gamma=[\operatorname{div}\sigma]\wedge\nu^{k-1}\wedge\gamma+\nu^{k}\wedge\gamma.$
Proposition 4.10.
If $\tau\colon Y^{\prime}\to Y$ is a proper mapping $\mu\in{\mathcal{QZ}}(Y^{\prime})$, then
(4.14) $[dd^{c}u]^{k}\wedge\tau_{*}\mu=\tau_{*}\big{(}[dd^{c}\tau^{*}u]^{k}\wedge\mu\big{)},\quad\langle dd^{c}u\rangle^{k}\wedge\tau_{*}\mu=\tau_{*}\big{(}\langle dd^{c}\tau^{*}u\rangle^{k}\wedge\mu\big{)},\\ \quad M_{k}^{u}\wedge\tau_{*}\mu=\tau_{*}\big{(}M_{k}^{\tau^{*}u}\wedge\mu\big{)},\quad k=0,1,2,\ldots.$
Proof.
The case $k=0$ is trivial, and by induction it is enough to consider $k=1$. Then the proposition follows from the equalities ${\bf 1}_{Z}\tau_{*}\mu=\tau_{*}({\bf 1}_{\tau^{-1}Z}\mu)$ and
$dd^{c}(u_{j}{\bf 1}_{Y\setminus Z}\tau_{*}\mu)=\tau_{*}\big{(}dd^{c}(\tau^{*}u_{j}{\bf 1}_{Y^{\prime}\setminus\tau^{-1}Z}\mu)\big{)},$
where $u_{j}$ is a sequence of smooth functions as in the proof of Proposition 4.8. ∎
Remark 4.11.
If $u$ is $\mathbb{Q}$-qas or $\mathbb{R}$-qas, we claim that one may indeed also in this case find a modification $\pi:Y^{\prime}\to Y$ such that in the local representations (2.5) of $\pi^{*}u$, the ideal generated by $f$ is principal. It then follows readily that the results in this section hold for such $u$ when ${\mathcal{QZ}}$ is replaced by ${\mathcal{QZ}}_{\mathbb{Q}}$ or ${\mathcal{QZ}}_{\mathbb{R}}$.
To prove the claim, we first claim that there is a well-defined minimal modification $\pi^{\prime\prime}:Y^{\prime\prime}\to Y$, where $Y^{\prime\prime}$ is normal and such that the ideal generated by $f$ in the local representation (2.5) of $\pi^{*}u$ is principal. Given such a modification, one obtains $\pi:Y^{\prime}\to Y$ by possible taking the corresponding reduced space, and taking a resolution of singularities $\pi^{\prime}:Y^{\prime}\to Y^{\prime\prime}$.
If $u$ is ${\mathbb{Q}}$-qas, then one readily verifies that $Y^{\prime\prime}$ may be defined locally by blowing up the ideal generated by $f$ in the local representation (2.5) of $u$ followed by the normalization. Given two representations (2.5) with $c,f$ and $c^{\prime},f^{\prime}$, they will be such that $|f^{\prime}|\sim|f|^{c/c^{\prime}}$, where $c/c^{\prime}\in{\mathbb{Q}}$, and then it follows by the following well-known facts that the above procedure is well-defined: If ${\mathcal{J}}\subset{\mathcal{O}}_{Y}$ is a coherent ideal, and $\bar{{\mathcal{J}}}$ denotes its integral closure, and $\operatorname{Bl}^{+}_{\mathcal{J}}Y$ denotes the normalization of the blowup of ${\mathcal{J}}$, then $\operatorname{Bl}^{+}_{{\mathcal{J}}^{k}}Y=\operatorname{Bl}^{+}_{\mathcal{J}}Y=\operatorname{Bl}^{+}_{\bar{{\mathcal{J}}}}Y$ for $k\geq 1$.
If $u$ is ${\mathbb{R}}$-qas, then given two representations (2.5) with $c,f$ and $c^{\prime},f^{\prime}$, then if $x\in Z(f)=Z(f^{\prime})$, the Lelong number $\nu(x,u)$ is a positive integer times $c$ and an positive integer times $c^{\prime}$, and hence, $c/c^{\prime}\in{\mathbb{Q}}$, and one may reduce to the rational case. ∎
4.2. Action of smooth Segre and Chern forms
Notice that if $L\to X$ is a line bundle with a smooth Hermitian metric $e^{\psi}$, then $c_{1}(L,e^{\psi})$ defines an endomorphism ${\bf c}_{1}(L,e^{\psi})$ on currents acting as multiplication by the smooth form $c_{1}(L,e^{\psi})$. Moreover, if $\mu=\tau_{*}\gamma$, then $c_{1}(L,e^{\psi})\wedge\tau_{*}\gamma=\tau_{*}(c_{1}(\tau^{*}L,\tau^{*}e^{\psi})\wedge\gamma)$ and hence ${\bf c}_{1}(L,e^{\psi})$ yields an endomorphism on ${\mathcal{QZ}}(X)$. We also have the endomorphism ${\bf s}_{1}(L,e^{\psi})=-{\bf c}_{1}(L,e^{\psi})$.
Let $E\to X$ be a Hermitian holomorphic vector bundle. Since $p\colon{\mathbb{P}}(E)\to X$ is a submersion, there is a well-defined morphism $p^{*}$ mapping currents on $X$ to currents on ${\mathbb{P}}(E)$ such that $p_{*}\xi\wedge\mu=p_{*}(\xi\wedge p^{*}\mu)$ for smooth forms $\xi$ on ${\mathbb{P}}(E)$.
Lemma 4.12.
We have
$p^{*}\colon{\mathcal{QZ}}(X)\to{\mathcal{QZ}}({\mathbb{P}}(E)).$
Proof.
By functoriality we have the commutative diagram
(4.15) $\begin{array}[c]{cccc}{\mathbb{P}}(\tau^{*}E)&\stackrel{{\scriptstyle\tau^{\prime}}}{{\longrightarrow}}&{\mathbb{P}}(E)\\ \downarrow\scriptstyle{p^{\prime}}&&\downarrow\scriptstyle{p}\\ W&\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}&X.\end{array}$
If $\mu=\tau_{*}\gamma$, thus $p^{*}\mu=\tau^{\prime}_{*}(p^{\prime})^{*}\gamma$, which is an element in ${\mathcal{QZ}}({\mathbb{P}}(E))$. ∎
Since $s(E,h)=p_{*}s(L_{E},e^{\psi})$, for any current $\mu$ on $X$,
(4.16) $s(E,h)\wedge\mu=p_{*}s(L_{E},e^{\psi})\wedge\mu=p_{*}(s(L_{E},e^{\psi})\wedge p^{*}\mu).$
It follows from the lemma that (4.16) defines endomorphisms
(4.17) ${\bf s}_{k}(E,h)\colon{\mathcal{QZ}}(X)\to{\mathcal{QZ}}(X),\quad k=0,1,\ldots,$
where ${\bf s}_{k}(E,h)\wedge\mu:=p_{*}(s_{1}(L_{E},e^{\psi})^{k+r-1}\wedge p^{*}\mu)$. In the same way we have endomorphisms
(4.18) ${\bf c}_{k}(E,h)\colon{\mathcal{QZ}}(X)\to{\mathcal{QZ}}(X),\quad k=0,1,\ldots$
defined by suitable compositions, cf. (2.4). Since these operators are just multiplication by $s_{k}(E,h)$ and $c_{k}(E,h)$, in particular we have that ${\bf c}_{0}(E,h)={\bf s}_{0}(E,h)=I$, the identity on ${\mathcal{QZ}}(X)$.
4.3. The submodule ${\mathcal{QZ}}_{0}(Y)$
Let $0\to S\to E\to Q\to 0$ be a short exact sequence of holomorphic vector bundles over a complex manifold $W$ with the metrics induced by a smooth metric on $E$. Then, by [11], for each component $\beta=(c(E)-c(S)c(Q))_{k}$ there is smooth form $v$ on $W$ such that $dd^{c}v=\beta$. It follows from (4.18) that $\beta$ induces an endomorphism ${\mathcal{QZ}}(W)\to{\mathcal{QZ}}(W)$,  $\mu\mapsto\beta\wedge\mu$.
Let $L$ be a holomorphic line bundle over $W$ with two different metrics $e^{b}$ and $e^{b^{\prime}}$, where $b^{\prime}$ is assumed to be smooth, and $b$ is qpsh and locally bounded. Then $v=b-b^{\prime}$ is a locally bounded qpsh function, and thus $e^{v}$ may be considered as a metric on the trivial line bundle. Let $\beta:=dd^{c}b^{\prime}-dd^{c}b=dd^{c}v$. Again we have a mapping ${\mathcal{QZ}}(W)\to{\mathcal{QZ}}(W)$,  $\mu\mapsto\beta\wedge\mu$. We say that any of these two forms $\beta$ is a beta form.
Definition 4.13.
We let ${\mathcal{QZ}}_{0}(Y)$ be the submodule of ${\mathcal{QZ}}(Y)$ generated by all currents of the form $\tau_{*}\mu^{\prime}$, $\tau\colon W\to Y$, where $\mu^{\prime}$ is in the image of one of the mappings $\beta$ above.
Thus ${\mathcal{QZ}}_{0}(Y)$ is generated by all $\tau_{*}(\beta\wedge\gamma)$ where $\beta$ is a beta form and $\gamma$ is a product of the form (4.1). For each such form there is a locally bounded function $v$ such that
(4.19) $dd^{c}\tau_{*}(v\wedge\gamma)=\tau_{*}(\beta\gamma).$
Remark 4.14.
Since the pullback of a beta form is a beta form, the module ${\mathcal{QZ}}_{0}(Y)$ is actually generated by all currents $\tau_{*}\gamma$, where $\gamma$ is as in (4.1) but with at least one factor $dd^{c}b_{j}$ ’replaced’ by a beta form. ∎
If $\tau\colon Y^{\prime}\to Y$ is a proper mapping, then $\tau_{*}\colon{\mathcal{QZ}}_{0}(Y^{\prime})\to{\mathcal{QZ}}_{0}(Y)$. In view of (4.6), for any analytic subspace $V\subset Y$ we have that
(4.20) ${\bf 1}_{V}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y).$
Remark 4.15.
Assume that $\mu\in{\mathcal{QZ}}_{0}(Y)$, and $\mu$ support on the analytic subspace $V\subset Y$. It then follows from Proposition 4.4 and (4.19) that there is a current $w$ of order 00 which is the pushforward of a current on $V$ such that $dd^{c}w=\mu$.
In particular, if $\mu\in{\mathcal{QZ}}_{0}(Y)$ has support at a point $x$ therefore
[…]
and since $\mu=\alpha[x]$, cf. Example 4.5, $\mu=0$. ∎
Proposition 4.16.
(i) If $u$ is a qas function on $Y$, then
$[dd^{c}u]^{k}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y),\quad k=0,1,2,\ldots.$
(ii) If $u^{\prime}$ is qas and $u^{\prime}=u+{\mathcal{O}}(1)$ locally on $Y$, then
$[dd^{c}u]^{k}-[dd^{c}u^{\prime}]^{k}\colon{\mathcal{QZ}}(Y)\to{\mathcal{QZ}}_{0}(Y),\quad k=0,1,2,\ldots.$
(iii) The mapping $p^{*}$ in Lemma 4.12 maps ${\mathcal{QZ}}_{0}(X)\to{\mathcal{QZ}}_{0}({\mathbb{P}}(E))$.
Proof.
Recall that if $\mu$ has support in $Z$, then $[dd^{c}u]\wedge\mu=0$. Otherwise, assume that $\mu=\tau_{*}(\beta\wedge\gamma)$ where $\gamma$ is product as in (4.1) and $\beta$ is a beta form. As in the proof of Proposition 4.8 we can therefore assume, cf. (4.9), that
(4.22) $[dd^{c}u]\wedge\mu=\tau_{*}\big{(}[\operatorname{div}\sigma]\wedge\beta\wedge\gamma+\nu\wedge\beta\wedge\gamma\big{)}.$
If $i\colon D\to W$, then $i^{*}(\beta\wedge\gamma)=i^{*}\beta\wedge i^{*}\gamma$ and thus $[\operatorname{div}\sigma]\wedge\beta\wedge\gamma$ is in ${\mathcal{QZ}}_{0}(W)$. Thus (i) holds for $k=1$ and hence for a general $k$ by induction.
By (i) and a simple induction over $k$ it is enough to prove (ii) for $k=1$. Let us assume that $\mu=\tau_{*}\gamma$, $\tau\colon W\to Y$. After a suitable further modification we can assume that $\tau^{*}u=\log|\sigma|^{2}+a$ and $\tau^{*}u^{\prime}=\log|\sigma^{\prime}|^{2}+a^{\prime}$ in $W$ where $a$ and $a^{\prime}$ are locally bounded. The assumption in (ii) implies that $[\operatorname{div}\sigma^{\prime}]=[\operatorname{div}\sigma]$. If $a^{\prime\prime}$ is some smooth metric on the line bundle corresponding to those divisors, then $dd^{c}(a-a^{\prime})=dd^{c}(a-a^{\prime\prime})-dd^{c}(a^{\prime}-a^{\prime\prime})=\beta-\beta^{\prime}$ where both $\beta$ and $\beta^{\prime}$ are beta forms on $W$. Therefore,
$([dd^{c}u^{\prime}]-[dd^{c}u])\wedge\mu=\tau_{*}(dd^{c}\tau^{*}u\wedge\gamma)-\tau_{*}(dd^{c}\tau^{*}u^{\prime}\wedge\gamma)=\tau_{*}(\beta\wedge\gamma)-\tau_{*}(\beta^{\prime}\wedge\gamma)$
which is in ${\mathcal{QZ}}_{0}(Y)$.
Part (iii) follows from the proof of Lemma 4.12. In fact, if $\mu$ is in ${\mathcal{QZ}}_{0}(X)$, then we can assume that $\mu=\tau_{*}(\beta\wedge\gamma)$ where $\gamma$ is of the form (4.1) and $\beta$ is a beta form. Then $p^{*}\mu=\tau^{\prime}_{*}(p^{\prime})^{*}(\beta\wedge\gamma)$ is in ${\mathcal{QZ}}_{0}({\mathbb{P}}(E))$. ∎
Definition 4.17.
We define the quotient module ${\mathcal{QB}}(Y):={\mathcal{QZ}}(Y)/{\mathcal{QZ}}_{0}(Y)$.
The definition is analogous with the definition of the quotient module ${\mathcal{B}}(Y)$ of ${\mathcal{GZ}}(Y)$, see [5], and there is a natural mapping $\mathcal{{\mathcal{B}}}(Y)\to{\mathcal{QB}}(Y)$.
4.4. Multiplicities of quasi-cycles
Let $\xi$ be a tuple of holomorphic functions that defines the maximal ideal at $x\in Y$. If $\mu\in{\mathcal{QZ}}(Y)$ and $u=\log|\xi|^{2}$, then $M^{u}\wedge\mu$ is in ${\mathcal{QZ}}(Y)$ and has support at $x$. Thus it has the form $\alpha[x]$ for some integer $\alpha$, cf. Example 4.5. By Proposition 4.16 (ii) and (4.21) it follows that $\alpha$ is independent of the choice of tuple $\xi$ defining the maximal ideal. If $\mu$ has bidegree $(k,k)$, for degree reasons, we thus have that $M^{u}\wedge\mu=M^{u}_{n-k}\wedge\mu$.
Definition 4.18.
Let $\mu\in{\mathcal{QZ}}(Y)$ be a $(k,k)$-current and $x\in Y$. We define the multiplicity of $\mu$ at $x$, $\operatorname{mult}_{x}\mu$, as
(4.23) $\operatorname{mult}_{x}\mu[x]=\int M^{u}\wedge\mu=\int{\bf 1}_{\{x\}}(dd^{c}\log|\xi|^{2})^{n-k}\wedge\mu.$
If $\mu\in{\mathcal{QZ}}_{0}(Y)$ then $M^{u}\wedge\mu\in{\mathcal{QZ}}_{0}(Y)$ by Proposition 4.16 (i), and hence $\operatorname{mult}_{x}\mu=0$, cf. (4.21). That is,
Proposition 4.19.
The multiplicity $\operatorname{mult}_{x}\mu$ only depends on the class of $\mu$ in ${\mathcal{QB}}(Y)$.
Remark 4.20.
If $\mu$ is a positive closed current, it is clear from (4.23), that $\operatorname{mult}_{x}\mu$ is the Lelong number of $\mu$ at $x$. ∎
Theorem 4.21.
Assume that $\mu\in{\mathcal{QZ}}(Y)$. Then any multiplicities $\operatorname{mult}_{x}\mu$ is an integer. If $\mu$ has bidegree $(k,k)$, then there is a unique decomposition
(4.24) $\mu=\mu_{F}+\mu_{M},$
where $\mu_{F}$ is the Lelong current of an analytic cycle on $Y$ and the set where $\operatorname{mult}_{x}\mu_{M}$ does not vanish is contained in an analytic set of codimension $\geq k+1$.
If $\mu^{\prime}$ defines the same class in ${\mathcal{QB}}(Y)$, then, cf. Proposition 4.19, it has the same multiplicities. It follows from the theorem that $\mu^{\prime}_{F}=\mu_{F}$. Clearly $\mu^{\prime}_{M}$ and $\mu_{M}$ may differ, but they still have the same multiplicities at each point. We say that $\mu_{F}$ is the fixed part and $\mu_{M}$ is the moving part of the class of $\mu$ in ${\mathcal{QB}}(Y)$.
Proof.
Indeed, it follows from Example 4.5 that $\operatorname{mult}_{x}\mu$ is an integer. Since we know that $\operatorname{mult}_{x}\mu$ is integer valued, (4.24) follows directly from the Siu decomposition in case $\mu$ is a positive current. As in the case with $\mu\in{\mathcal{GZ}}(Y)$, if $\mu=\tau_{*}\gamma$, and $\tau=i\circ\tau^{\prime}$, where $\tau^{\prime}\colon W\to Z$ is surjective, then $\operatorname{mult}_{x}\tau^{\prime}_{*}\gamma=\operatorname{mult}_{i(x)}\tau_{*}\gamma$. Moreover, [1, Lemma 7.3] holds, with the same proof, for $\mu=\tau_{*}\gamma$ if $\gamma$ is as in (4.1). In fact, the crucial point that $M^{\tau^{*}\xi}\wedge\gamma=0$ if $M^{\tau^{*}\xi}=M^{\tau^{*}\xi}\wedge 1=0$. To see this we can assume that $\xi$ is a tuple of functions (i.e., a section of a vector bundle with trivial metric) in a neighborhood of $x$. Then $M^{\tau^{*}\xi}$ is a positive current. If $b_{\ell}$ is locally of the form $u_{\ell}+g_{\ell}$, where $u_{\ell}$ is psh, and $g_{\ell}$ is smooth, we may take a sequence $b_{\ell}^{j}=u_{\ell}^{j}+g_{\ell}$, where each $u_{\ell}^{j}$ is psh, and $u_{\ell}^{j}$ decreases to $u_{\ell}$. Then, by [13, Proposition III.4.9], $M^{\tau^{*}\xi}\wedge\gamma_{j}\to M^{\tau^{*}\xi}\wedge\gamma$. If $M^{\tau^{*}\xi}=0$, thus $M^{\tau^{*}\xi}\wedge\gamma_{j}=0$ and hence $M^{\tau^{*}\xi}\wedge\gamma=0$. Now Theorem 4.21 follows precisely as [1, Theorem 7.1]. ∎
Remark 4.22.
It follows that if $\mu\in{\mathcal{QZ}}_{\mathbb{Q}}(Y)$, then $\operatorname{mult}_{x}\mu$ is a rational number. ∎
Example 4.23.
Assume that $\mu\in{\mathcal{QZ}}(Y)$ has bidegree $(p,p)$ and has support on an analytic subspace $V\subset Y$ of pure (co)dimension. Since $\mu$ is a normal current, it follows that $\mu$ is an analytic cycle a priori by real coefficients if ${\rm codim\,}V=p$, and $\mu=0$ if ${\rm codim\,}V>p$. Since the multiplicities are integral we see that it is actually a cycle with integer coefficients, i.e., in ${\mathcal{Z}}(Y)$. If $\mu$ is in ${\mathcal{QZ}}_{0}(Y)$, then all its multiplicities vanish, and hence it is 00. Thus ${\mathcal{Z}}(Y)$ is indeed a submodule of ${\mathcal{QB}}(Y)$. ∎
5. Regularization of qas metrics on a line bundle
Let $L\to Y$ be a holomorphic vector bundle, $e^{\psi_{0}}$ a smooth reference metric on $L$, and let $\omega=s_{1}(L,e^{\psi_{0}})$. Note that a singular metric $e^{\psi}$ on $L$ is a qas if and only if the global function $\psi-\psi_{0}$ is qas. Indeed, given a local trivialization $L\cong X\times{\mathbb{C}}_{\alpha}$,
$\psi-\psi_{0}=\log|\alpha|^{2}_{e^{\psi}}-\log|\alpha|^{2}_{e^{\psi_{0}}}=\log|1|^{2}_{e^{\psi}}-\log|1|^{2}_{e^{\psi_{0}}},$
so if the metric $e^{\psi}$ is qas, then the right-hand side is qas by Lemma 2.1. The other direction follows by writing
$\log|\alpha|^{2}_{e^{\psi}}=(\psi-\psi_{0})+\log|\alpha|^{2}_{e^{\psi_{0}}},$
and using that $\log|\alpha|^{2}_{e^{\psi_{0}}}$ is qas.
We define the endomorphism
(5.1) $[dd^{c}\psi]_{\omega}\wedge\mu=[dd^{c}(\psi-\psi_{0})]\wedge\mu+\omega\wedge\mu$
on ${\mathcal{QZ}}(Y)$. It follows from Proposition 4.8 that (5.1) indeed maps ${\mathcal{QZ}}(Y)$ to ${\mathcal{QZ}}(Y)$.
Remark 5.1.
Assume that $L$ is trivial and $\psi_{0}=0$. Then the notation is consistent if $[dd^{c}\psi]=[dd^{c}\psi]_{0}$. ∎
Remark 5.2.
If $Z=Y$, i.e., $\psi\equiv-\infty$, it is natural to let $[dd^{c}\psi]_{\omega}\wedge\mu=\omega\wedge\mu$ for all $\mu$. ∎
Proposition 5.3.
(i) If $e^{\psi}$ is a qas metric on $L\to Y$, then $[dd^{c}\psi]^{k}_{\omega}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y)$.
(ii) If $e^{\psi^{\prime}}$ is another qas metric and $\psi^{\prime}=\psi+{\mathcal{O}}(1)$ then $[dd^{c}\psi^{\prime}]^{k}_{\omega}-[dd^{c}\psi]^{k}_{\omega}\colon{\mathcal{QZ}}(Y)\to{\mathcal{QZ}}_{0}(Y)$.
(iii) If $\tau\colon Y^{\prime}\to Y$ is a proper mapping and $\mu=\tau_{*}\mu^{\prime}$, then
(5.2) $[dd^{c}\psi]_{\omega}\wedge\mu=\tau_{*}\big{(}[dd^{c}\tau^{*}\psi]_{\tau^{*}\omega}\wedge\mu^{\prime}\big{)}.$
Proof.
The case $k=1$ of (i) follows directly from (5.1), Proposition 4.16 (i), and the fact that multiplication by $\omega$ maps ${\mathcal{QZ}}_{0}(Y)$ into itself. The general case follows by trivial induction. Part (ii) follows direct from Proposition 4.16 (ii) applied to $u^{\prime}=\psi^{\prime}-\psi_{0}$ and $u=\psi-\psi_{0}$. Part (iii) follows from Proposition 4.10. ∎
Although $[dd^{c}\psi]_{\omega}$ seems to be the fundamental endomorphism, it is sometimes useful to express it in terms of $[dd^{c}\psi]$ which is well-defined although $\psi$ is only a metric, not a global function. Let $Z$ be the unbounded locus of $e^{\psi}$, and let $\omega=dd^{c}\psi_{0}$. If $\mu={\bf 1}_{Y\setminus Z}\mu$, then (5.1) is
$[dd^{c}(\psi-\psi_{0})]\wedge\mu+\omega\wedge\mu=:[dd^{c}\psi]\wedge\mu,$
and if $\mu={\bf 1}_{Z}\mu$, then it is just $\omega\wedge\mu$. Therefore,
(5.3) $[dd^{c}\psi]_{\omega}\wedge\mu=[dd^{c}\psi]\wedge{\bf 1}_{Y\setminus Z}\mu+\omega\wedge{\bf 1}_{Z}\mu,$
which coincides with how the corresponding operator is defined in [18]. We also have endomorphisms
$\langle dd^{c}\psi\rangle^{k}\wedge\mu={\bf 1}_{Y\setminus Z}[dd^{c}\psi]^{k}\wedge\mu,\quad M^{\psi}_{k}\wedge\mu={\bf 1}_{Z}[dd^{c}\psi]^{k}\wedge\mu;$
on ${\mathcal{QZ}}(Y)$. It is readily verified by induction that
(5.4) $[dd^{c}\psi]^{k}_{\omega}\wedge\mu=\langle dd^{c}\psi\rangle^{k}\wedge\mu+\sum_{\ell=0}^{k}\omega^{k-\ell}\wedge M^{\psi}_{\ell}\wedge\mu,\quad k=0,1,2,\ldots.$
Theorem 5.4.
Let $\chi_{j}$ be a sequence of smooth bounded from below convex increasing functions on ${\mathbb{R}}$, which decrease to the identity function as $j\to\infty$. Let $e^{\psi}$ be a qas metric on $L\to Y$, $\psi_{0}$ a smooth metric, and let $\psi_{j}=\chi_{j}\circ(\psi-\psi_{0})+\psi_{0}$. If $\omega=dd^{c}\psi_{0}$, then for each $\mu\in{\mathcal{QZ}}(Y)$,
(5.5) $(dd^{c}\psi_{j})^{k}\wedge\mu\to[dd^{c}\psi]^{k}_{\omega}\wedge\mu,\quad j\to\infty,\quad k=0,1,2,\ldots.$
Since $\psi-\psi_{0}$ is a global function, the left hand side in (5.5) is well-defined. The smoothness assumption on $\chi_{j}$ is not necessary, but for simplicity we restrict to this case, since we will only apply the theorem to $\chi_{\epsilon}(t)=\log(e^{t}+{\epsilon})$.
The theorem is known when $\mu$ is $1$, however expressed with the right hand side in (5.4). In [7] it is proved for a more general class of (q)psh functions. Here we will rely on the presentation in [10] and show that the proof given there works even when acting on a quasi-cycle $\mu$.
Proof.
Since the statement is local, we can restrict to an open subset where we have a global frame element, and hence that $\psi$ and $\psi_{0}$ are fixed functions. For simplicity we let $Y$ denote this open set.
We first consider the case when $\psi_{0}=0$. If $\mu$ has support on $Z$, then both sides of (5.5) are equal to $\omega^{k}\mu$. Thus we can assume that $\mu=\tau^{*}\gamma$, where $\tau\colon W\to Y$, $\gamma$ is a product as in (4.1) and $\tau^{-1}Z$ has positive codimension in $W$. After possibly another modification we may assume that $\tau^{*}\psi=\log|\sigma|^{2}+a$, where $\sigma$ is a section of a Hermitian line bundle ${\mathcal{L}}\to W$ and $a$ is qpsh and locally bounded. We claim that
(5.6) $(dd^{c}\chi_{j}\circ\tau^{*}\psi)^{k}\wedge\gamma\to[dd^{c}\tau^{*}\psi]^{k}\wedge\gamma.$
Following [10] there are unique convex increasing functions $\gamma_{j}$ such that $\gamma_{j}(-1)=\chi_{j}(-1)$ and $\gamma^{\prime}_{j}=(\chi^{\prime}_{j})^{k}.$ It follows that also $\gamma_{j}(t)$ decreases to $t$. A direct computation yields
$(dd^{c}(\chi_{j}\circ\tau^{*}\psi))^{k}\wedge\gamma=dd^{c}((\gamma_{j}\circ\tau^{*}\psi)\nu^{k-1}\wedge\gamma),$
where $\nu=dd^{c}a-c_{1}({\mathcal{L}})$. Since $\gamma_{j}\circ\tau^{*}\psi\to\tau^{*}\psi$ in $L^{1}$ with respect to the (total variation of the) measure $\nu^{k-1}\wedge\gamma$, we see that by definition,
$(dd^{c}\chi_{j}\circ\tau^{*}\psi)^{k}\wedge\gamma\to dd^{c}(\tau^{*}\psi\nu^{k-1}\wedge\gamma)=([\operatorname{div}\sigma]+\nu)\wedge\nu^{k-1}\wedge\gamma=[dd^{c}\tau^{*}\psi]^{k}\wedge\gamma;$
Here the last equality follows from Example 4.9 and $\nu=dd^{c}a-c_{1}({\mathcal{L}})$. Thus (5.6) holds. Applying $\tau_{*}$ and using Proposition 4.10, now (5.5) follows for $\psi_{0}=0$.
Let us now assume that $\psi_{0}$ is an arbitrary smooth function. Using the case $\psi_{0}=0$,
(5.7) $(dd^{c}\psi_{j})^{k}\wedge\mu=\sum_{\ell=0}^{k}{k\choose\ell}\omega^{k-\ell}\wedge(dd^{c}\chi_{j}\circ(\psi-\psi_{0}))^{\ell}\wedge\mu\to\\ \sum_{\ell=0}^{k}{k\choose\ell}\omega^{k-\ell}\wedge[dd^{c}(\psi-\psi_{0})]^{\ell}\wedge\mu=([dd^{c}(\psi-\psi_{0})]+\omega)^{k}\wedge\mu,$
where the last power is composition of the endomorphism $\mu\mapsto([dd^{c}(\psi-\psi_{0})]+\omega)\wedge\mu$ on ${\mathcal{QZ}}(Y)$. Now (5.5) follows from (5.1) and (5.7). ∎
6. Segre forms and endomorphisms of qas bundles
Assume that $(E,h)$ is a qas bundle over $X$. If $e^{\psi}$ denotes the induced metric on $L_{E}$, then $\psi=\log|\alpha|^{2}_{h}$ is qas, i.e., for any local frame $a$ of $L_{E}$, i.e., the well-defined function $\log|\alpha a|^{2}$ is qas. Note that on the set where $\alpha_{j}\neq 0$ we may take $a=1/\alpha_{j}$.
Let $h_{0}$ be a smooth reference metric on $E$, let $e^{\psi_{0}}$ denote the induced metric on $L_{E}$, and $\omega=dd^{c}\psi_{0}$. We thus have that
$s(L,e^{\psi_{0}})=\sum_{\ell=0}^{\infty}\omega^{\ell}=\frac{1}{1-\omega}.$
If $Z^{\prime}\subset{\mathbb{P}}(E)$ denotes the set where $\psi$ is not locally bounded, then we have outside of $Z^{\prime}$ that $s_{1}(L_{E},e^{\psi})=dd^{c}\psi$. Recall, cf. (5.1) and (5.3), that
(6.1) $[dd^{c}\psi]_{\omega}\wedge\mu=dd^{c}(\psi{\bf 1}_{X\setminus Z^{\prime}}\mu)+\omega\wedge{\bf 1}_{Z^{\prime}}\mu$
defines an endomorphism on ${\mathcal{QZ}}({\mathbb{P}}(E))$. In view of Lemma 4.12 and Proposition 4.16 (iii) the following definition makes sense.
Definition 6.1.
Given a qas bundle $(E,h)$ over $X$ we have the Segre endomorphisms ${\mathcal{QZ}}(X)\to{\mathcal{QZ}}(X)$ and ${\mathcal{QZ}}_{0}(X)\to{\mathcal{QZ}}_{0}(X)$ of degree $k$, defined by the formula
(6.2) ${\bf s}_{k}(E,h,h_{0})\wedge\mu=p_{*}([dd^{c}\psi]_{\omega}^{k+r-1}\wedge p^{*}\mu),$
and ${\bf s}(E,h,h_{0}):=\sum_{k}{\bf s}_{k}(E,h,h_{0})$.
If $h$ is smooth, $[dd^{c}\psi]_{\omega}$ is just multiplication with the smooth form $dd^{c}\psi$, so ${\bf s}_{k}(E,h,h_{0})$ is just multiplication by the smooth form $s_{k}(E,h)$, see (4.16).
We define the Segre form as $s(E,h,h_{0})={\bf s}(E,h,h_{0})1$ on $X$, so that
(6.3) $s(E,h,h_{0}):=p_{*}\sum_{k=0}^{\infty}p_{*}[dd^{c}\psi]^{k}_{\omega}.$
Note that if $L$ is a line bundle, then $p_{L}:{\mathbb{P}}(L)\to X$ is the identity. Hence,
(6.4) $s(E,h,h_{0}):=p_{*}s(L_{E},e^{\psi},e^{\psi_{0}}).$
Remark 6.2.
The definition of $s(E,h,h_{0})$ and ${\bf s}(E,h,h_{0})$ here coincides with the definition in [18] except that in those articles is used the convention that ${\mathbb{P}}(E)_{x}$ denotes the hyperplanes in $E_{x}$ so one has to make some necessary adaptions. Indeed, in [18], the main results are formulated for singular metrics $h$ which are positively curved, in contrast to the negatively curved metrics we consider here. This difference is however essentially artificial. For a singular metric $h$, there is an induced dual metric $h^{*}$ on the dual bundle $E^{*}$, see e.g., [17, Lemma 3.1]. By definition, a singular metric $h$ on $E$ is Griffiths positive if $h^{*}$ is Griffiths negative. The definition of Segre form in [18] is such that if $h$ is a Griffiths positive singular metric on $E$, then it holds just as in the smooth case that
$s_{k}(E,h,h_{0})=(-1)^{k}s_{k}(E^{*},h^{*},h_{0}^{*}),$
where now the left-hand side is the Segre form defined in [18], and the right-hand side is the Segre form defined in this article. ∎
Below, we will sometimes drop $h$ and $h_{0}$ from the notation of the Segre forms and operators, and consider $E$ to be a qas bundle, i.e., equipped with a qas metric $h$, and furthermore a smooth reference metric $h_{0}$.
By (5.4), with $Y={\mathbb{P}}(E)$ and $Z=Z^{\prime}$, we have
(6.5) $\sum_{k=0}^{\infty}[dd^{c}\psi]^{k}_{\omega}\wedge\mu=\sum_{k=0}^{\infty}\langle dd^{c}\psi\rangle^{k}\wedge\mu+\sum_{\ell=0}^{\infty}\sum_{k=\ell}^{\infty}M^{\psi}_{\ell}\wedge(dd^{c}\psi_{0})^{k-\ell}\wedge\mu=\\ \left(\sum_{k=0}^{\infty}\langle dd^{c}\psi\rangle^{k}+s(L_{E},e^{\psi_{0}})\wedge M^{\psi}\right)\wedge\mu.$
Thus
(6.6) ${\bf s}(E,h,h_{0})\wedge\mu=\sum_{\ell=0}^{\infty}p_{*}(\langle dd^{c}\psi\rangle^{\ell}\wedge p^{*}\mu)+p_{*}\big{(}s(L_{E},e^{\psi_{0}})\wedge M^{\psi}\wedge p^{*}\mu\big{)}.$
We may decompose ${\bf s}_{k}(E,h,h_{0})\wedge\mu$ as
${\bf s}_{k}^{\prime}(E,h,h_{0})\wedge\mu+M^{E,h,h_{0}}\wedge\mu:={\bf 1}_{X\setminus Z}{\bf s}_{k}(E,h,h_{0})\wedge\mu+{\bf 1}_{Z}{\bf s}_{k}(E,h,h_{0})\wedge\mu.$
By (6.6), it follows that
$s_{k}^{\prime}(E,h,h_{0})\wedge\mu=p_{*}(\langle dd^{c}\psi\rangle^{\ell}\wedge p^{*}\mu)\text{ and }M^{E,h,h_{0}}_{k}\wedge\mu=s(L_{E},e^{\psi_{0}})\wedge M^{\psi}\wedge p^{*}\mu.$
In particular, it follows that $s^{\prime}(E,h,h_{0})$ is actually independent of $h_{0}$, and may hence be denoted simply by $s^{\prime}(E,h)$.
Example 6.3.
Let $(E,h_{0})$ and $(F,h_{F})$ be Hermitian vector bundles over $X$, and let $g\colon E\to F$ be a holomorphic morphism. Let $h$ be the singular metric on $E$ defined by $|\alpha|_{h}:=|g\alpha|_{h_{F}}$. Recall the representation $\psi=\log|\alpha|_{h}^{2}/|\alpha_{j}|^{2}$ of the induced metric on $L_{E}$, (2.2). In [1], the corresponding representation is denoted by $\log|\alpha|^{2}_{\circ}$. It follows from (6.6), with $\mu=1$, that the definition given in [1, Example 9.2] of $s(E)$ coincides with (6.3). ∎
Proposition 6.4.
Assume that $E\to X$ is a holomorphic vector bundle equipped with a locally bounded qas metric $h$. The Segre endomorphism induced by $h$ on ${\mathcal{QZ}}(X)$ is independent of the choice of reference metric, and may thus be denoted by ${\bf s}(E,h)$.
Locally, there exists a sequence of smooth metrics $h_{\epsilon}$ decreasing to $h$ when $\epsilon\to 0$ such that
$s(E,h_{\epsilon})\wedge\mu\to{\bf s}(E,h)\wedge\mu,\quad{\epsilon}\to 0,\quad\mu\in{\mathcal{QZ}}(X).$
If $h$ is of the form $h_{1}\oplus h_{2}$, then $h_{\epsilon}=(h_{1})_{\epsilon}\oplus(h_{2})_{\epsilon}$, where $(\ )_{\epsilon}$ denotes such a regularization of the corresponding metric.
In case $h$ is Griffiths negative, then such a regularization may be obtained from [9, Proposition 3.1] by representing $h$ as a matrix in a given trivialization of $E$, and letting $h_{\epsilon}=h\ast\rho_{\epsilon}$, where $\ast$ denotes the convolution with respect to a choice of local coordinates $x$, and $\rho_{\epsilon}(x)$ is a family of radial mollifiers, and it will follow by the Bedford-Taylor theory and the definition of ${\bf s}(E,h)$ that the corresponding currents converge. If $h$ is just qas, then one may instead take
(6.7) $h_{\epsilon}=h\circ\rho_{\epsilon}+A{\epsilon}I,$
where $A$ is chosen to be a large enough constant to make $h_{\epsilon}$ decreasing in ${\epsilon}$ when ${\epsilon}\to 0$, and $I$ denotes the identity matrix.
Proof.
As remarked above, $s^{\prime}(E,h,h_{0})$ is independent of the choice of reference metric $h_{0}$. If $Z=\emptyset$, then $s(E,h,h_{0})=s^{\prime}(E,h,h_{0})$, and hence $s(E,h,h_{0})$ is also independent of $h_{0}$.
Fix a trivialization of $E$, and a local coordinate system $x$ on $X$ such that locally, $E\cong{\mathbb{C}}^{r}_{\alpha}\times{\mathbb{C}}^{n}_{x}$. Define $h_{\epsilon}$ by (6.7). Note that $h_{\epsilon}$ depends on both the trivialization of $E$ and the local coordinate system on $X$. It is readily verified that a convolution of positive definite matrices is positive definite, and the entries are smooth, so $h_{\epsilon}$ indeed defines locally a smooth metric on $E$.
To prove the convergence result, assume that we are on the subset where $\alpha_{j}\neq 0$. With respect to the frame induced by $\alpha/\alpha_{j}$ of $L_{E}$, the metric $e^{\psi_{\epsilon}}$ induced by $h_{\epsilon}$ is represented by the function $\psi_{\epsilon}=\log|\alpha/\alpha_{j}|_{h_{\epsilon}(x)}$. Then $\psi_{\epsilon}$ is exactly of the form (2.6) with $\psi=\log|\alpha/\alpha_{j}|^{2}_{h(x)}$, and $v=\log(|\alpha/\alpha_{j}|^{2})$. By Proposition 2.3, if $A$ is chosen large enough, $\psi_{\epsilon}$ may locally be decomposed as $u_{\epsilon}+g$, where $u_{\epsilon}$ and $g$ are smooth and $u_{\epsilon}$ decrease pointwise to a locally bounded function $u$. It then follows by the Bedford-Taylor theory that
$s(L_{E},e^{\psi_{\epsilon}})\wedge\pi^{*}\mu=\sum_{k}(dd^{c}\psi_{\epsilon})^{k+r-1}\wedge p^{*}\mu\to\sum_{k}(dd^{c}\psi)^{k+r-1}\wedge p^{*}\mu.$
Taking the pushforward to $X$ finishes the proof of the convergence.
The statement about the case when $h=h_{1}\oplus h_{2}$ follows readily by noting that a convolution preserves a direct sum. ∎
Proof of Proposition 1.3.
Proposition 1.3 follows immediately from Proposition 6.4 by taking $\mu=1$. ∎
Here is our main theorem on the Segre endomorphism ${\bf s}(E)$.
Theorem 6.5.
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$ and a smooth reference metric $h_{0}$, and let $Z$ be the unbounded locus of $h$. Let ${\bf s}(E,h,h_{0})$ be the associated Segre endomorphism on ${\mathcal{QZ}}(X)$.
(i) If $h_{\epsilon}$ is the metric defined by (1.1), then
(6.8) ${\bf s}(E,h_{\epsilon})\wedge\mu\to{\bf s}(E,h,h_{0})\wedge\mu,\quad{\epsilon}\to 0,\quad\mu\in{\mathcal{QZ}}(X).$
(ii) The endomorphisms ${\bf s}_{k}(E,h,h_{0})$ vanish for $k<0$ and ${\bf s}_{0}(E,h,h_{0})=I$, the identity mapping.
(iii) If $\tau\colon X^{\prime}\to X$ is a proper mapping, then $\tau^{*}h$ is qas and,
(6.9) $\tau_{*}({\bf s}(\tau^{*}E,\tau^{*}h,\tau^{*}h_{0})\wedge\mu)={\bf s}(E,h,h_{0})\wedge\tau_{*}\mu,\quad\mu\in{\mathcal{QZ}}(X^{\prime}).$
(iv) For any $\mu\in{\mathcal{QZ}}(X)$ there is a current $w$ on $X$ such that
(6.10) $({\bf s}(E,h,h_{0})-{\bf s}(E,h_{0}))\wedge\mu=dd^{c}w.$
(v) If $E^{\prime}\to X$ is another holomorphic vector bundle, equipped with a smooth metric $h^{\prime}$, then
(6.11) ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})\wedge\mu=s(E^{\prime},h^{\prime})\wedge{\bf s}(E,h,h_{0})\wedge\mu,\quad\mu\in{\mathcal{QZ}}(X).$
(vi) For each $k$, ${\bf s}_{k}(E,h,h_{0})$ yields a well-defined endomorphism on ${\mathcal{QB}}(X)$, independent of $h_{0}$, and if $h^{\prime}$ is a qas metric on $E$ comparable to $h$, then ${\bf s}_{k}(E,h^{\prime},h_{0})$ yields the same endomorphism as ${\bf s}_{k}(E,h,h_{0})$ on ${\mathcal{QB}}(X)$.
(vii) Assume that $E_{1}$ and $E_{2}$ are qas bundles, and let $Z_{E_{j}}$ denotes the unbounded locus of $Z_{E_{j}}$. For any $k,\ell$, there is a normal current $w$, which is the push-forward of a current on $Z_{E_{1}}\cap Z_{E_{2}}$ such that
$({\bf s}_{k}(E_{1})\wedge{\bf s}_{\ell}(E_{2})-{\bf s}_{\ell}(E_{2})\wedge{\bf s}_{k}(E_{1}))\wedge\mu=dd^{c}w.$
Remark 6.6.
(i) If $h_{0}$ and $h_{0}^{\prime}$ are two smooth reference metrics, then by the first part of (vi), $({\bf s}(E,h,h_{0})\wedge\mu-{\bf s}(E,h,h_{0}^{\prime}))\wedge\mu$ belongs to ${\mathcal{QZ}}_{0}(X)$. Since ${\bf s}(E,h,h_{0})$ and ${\bf s}(E,h,h_{0}^{\prime})$ coincide outside of $Z$, it follows by Remark 4.15 that there is a current $w$, which is the pushforward of a current on $Z$ such that
(6.12) $\big{(}{\bf s}_{k}(E,h,h_{0})-{\bf s}_{k}(E,h,h_{0}^{\prime}))\wedge\mu=dd^{c}w.$
In particular, if $\mu$ has bidegree $(q,q)$, then $w$ has bidegree $(k+q-1,k+q-1)$, and hence it vanishes if $k+q\leq{\rm codim\,}Z$.
(ii) Note that if $\mu$ has bidegree $(q,q)$, then the current $w$ in (vii) has bidegree $(k+\ell+q-1,k+\ell+q-1)$. Since $w$ is the push-forward of a current on $Z$, if $k+\ell+q\leq{\rm codim\,}Z$, then $w$ vanishes. ∎
Proof.
Notice that if $\chi_{\epsilon}(t)=\log(e^{t}+{\epsilon})$, then
$\psi_{\epsilon}:=\log(|\alpha|_{h}^{2}+{\epsilon}|\alpha|^{2}_{h_{0}})=\log(|\alpha|_{h}^{2}/|\alpha|^{2}_{h_{0}}+{\epsilon})+\log|\alpha|^{2}_{h_{0}}=\log(e^{\psi-\psi_{0}}+{\epsilon})+\psi_{0}=\chi_{\epsilon}\circ(\psi-\psi_{0})+\psi_{0},$
so it follows from Theorem 5.4 that
$(dd^{c}\psi_{\epsilon})^{k}\wedge p^{*}\mu\to[dd^{c}\psi]^{k}_{\omega}\wedge p^{*}\mu.$
Applying $p_{*}$ we get (6.8) and hence (i) is proved.
Now (ii) follows since it holds when the metric is smooth, and hence also when it is locally bounded qas by Proposition 6.4.
Notice that a holomorphic map $\tau\colon X^{\prime}\to X$ induces a holomorphic map $\tau\colon{\mathbb{P}}(\tau^{*}E)\to{\mathbb{P}}(E)$. Part (iii) now follows from the definition (6.2) and Proposition 5.3 (iii).
In (iv) we can assume that $\mu=\tau_{*}\gamma$ where $\gamma$ is a product as in (4.1) on $W$ and $\tau\colon W\to X$. In view of (iii) and by changing notation we have to prove (iv) on $X$ but with $\mu=\gamma$.
If the metric on $E$ is identically 00, i.e., $Z^{\prime}={\mathbb{P}}(E)$, then ${\bf s}(E,h,h_{0})={\bf s}(E,h_{0})$ and then the statement is trivial. Let us thus assume that $Z^{\prime}$ is not equal to ${\mathbb{P}}(E)$. It may very well happen anyhow that $Z=X$. Notice that $p^{*}\gamma$ is a product as in (4.1) on ${\mathbb{P}}(E)$. Let $\psi$ and $\psi_{0}$ be as above. Then $\psi-\psi_{0}$ is a global function on ${\mathbb{P}}(E)$ and, cf. Proposition 4.8,
(6.13) $w^{\prime}=\frac{\psi-\psi_{0}}{(1-dd^{c}\psi_{0})(1-\langle dd^{c}\psi\rangle)}\wedge p^{*}\gamma$
is a locally finite measure. In view of (4.11) and (4.12) we have
(6.14) ${\bf 1}_{{\mathbb{P}}(E)\setminus Z^{\prime}}dd^{c}w^{\prime}=\left(\frac{1}{1-\langle dd^{c}\psi\rangle}-\frac{1}{1-dd^{c}\psi_{0}}\right)\wedge p^{*}\gamma$
and
(6.15) ${\bf 1}_{Z^{\prime}}dd^{c}w^{\prime}=\frac{1}{1-dd^{c}\psi_{0}}{\bf 1}_{Z^{\prime}}dd^{c}\psi\frac{1}{1-dd^{c}\psi}\wedge p^{*}\gamma=s(L_{E},e^{\psi_{0}})\wedge\sum_{\ell=1}^{\infty}M^{\psi}_{\ell}\wedge p^{*}\gamma.$
Notice that $Z^{\prime}$ has positive codimension so that $M^{\psi}_{0}\wedge p^{*}\gamma={\bf 1}_{Z^{\prime}}p^{*}\gamma=0$. Letting $w=p_{*}w^{\prime}$ we conclude from (6.6), (6.14) and (6.15) that $dd^{c}w=({\bf s}(E,h,h_{0})-{\bf s}(E,h_{0}))\wedge\mu.$ Thus (iv) is proved.
To prove (v), consider the metric $h_{\epsilon}^{\prime\prime}:=(h^{\prime}\oplus h)+{\epsilon}(h^{\prime}\oplus h_{0})$, which has empty unbounded locus. By Proposition 6.4, we may locally find smooth metrics $h_{{\epsilon},{\epsilon}^{\prime}}$ such that ${\bf s}(E,h_{{\epsilon},{\epsilon}^{\prime}})\to{\bf s}(E,h_{\epsilon}^{\prime\prime})$ when ${\epsilon}^{\prime}\to 0$ and $h_{{\epsilon},{\epsilon}^{\prime}}=((1+{\epsilon})h^{\prime})_{{\epsilon}^{\prime}}\oplus(h+{\epsilon}h_{0})_{{\epsilon}^{\prime}}$, where $(\ )_{{\epsilon}^{\prime}}$ denotes the regularization given by Proposition 6.4 for the corresponding metric. For each ${\epsilon}^{\prime}>0$ we have
$s(E,h_{{\epsilon},{\epsilon}^{\prime}})=s(E,((1+{\epsilon})h^{\prime})_{{\epsilon}^{\prime}})\wedge s(E,(h_{\epsilon})_{{\epsilon}^{\prime}}).$
Multiplying this equality of smooth forms with $\mu$ and letting ${\epsilon}^{\prime}\to 0$, it follows that
(6.16) ${\bf s}(E,h_{\epsilon})\wedge\mu=s(E,(1+{\epsilon})h^{\prime})\wedge s(E,h_{\epsilon})\wedge\mu.$
By (i) the left hand side tends to ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})\wedge\mu$. Notice that $s(E,(1+{\epsilon})h^{\prime})=s(E,h^{\prime})$. The right hand side of (6.16) is therefore equal to ${s(E,h^{\prime})\wedge s(E,h_{\epsilon})\wedge\mu}$ that tends to $s(E,h^{\prime})\wedge{\bf s}(E,h,h_{0})\wedge\mu$ by (i).
Let $h_{0}^{\prime}$ denote another smooth reference metric. By (6.6) we see that ${\bf s}(E,h,h_{0}^{\prime})-{\bf s}(E,h,h_{0})$ is the image under $p_{*}$ of $(s(L_{E},e^{\psi_{0}^{\prime}})-s(L_{E},e^{\psi_{0}}))\wedge M^{\psi}$, which belongs to ${\mathcal{QZ}}_{0}({\mathbb{P}}(E))$. Now the first statement follows. The second one follows from Proposition 5.3 (ii) and (6.2).
It remains to prove (vii). By (6.4), it suffices to prove the result when $(E_{1},E_{2},k,\ell)$ is replaced by $(L_{E_{1}},L_{E_{2}},k+{\rm rank\,}E_{1}-1,\ell+{\rm rank\,}E_{2}-1)$. We may thus simply assume that ${\rm rank\,}E_{1}={\rm rank\,}E_{2}=1$. In that case, ${\bf s}_{k}={\bf s}_{1}^{k}$, so
$\big{(}{\bf s}_{k}(E_{1})\wedge{\bf s}_{\ell}(E_{2})-{\bf s}_{\ell}(E_{2})\wedge{\bf s}_{k}(E_{1})\big{)}\wedge\mu=({\bf s}_{1}(E_{1})^{k}\wedge{\bf s}_{1}(E_{2})^{\ell}-{\bf s}_{1}(E_{2})^{\ell}\wedge{\bf s}_{1}(E_{1})^{k})\wedge\mu.$
We may rewrite the right-hand side as a sum of terms of the form
$\mathbf{S}^{+}\wedge\big{(}{\bf s}_{1}(E_{1})\wedge{\bf s}_{1}(E_{2})-{\bf s}_{1}(E_{2})\wedge{\bf s}_{1}(E_{1})\big{)}\wedge\mathbf{S}^{-}\wedge\mu,$
where $\mathbf{S}^{+}$ and $\mathbf{S}^{-}$ are products of ${\bf s}_{1}(E_{1})$ and ${\bf s}_{1}(E_{2})$. Let $Z=Z_{E_{1}}\cap Z_{E_{2}}$. It suffices to prove that each such term is of the form $dd^{c}w$, where $w$ is a current that is the push-forward of a current on $Z$. Since $\mathbf{S}^{-}\wedge\mu$ is in ${\mathcal{QZ}}$, we obtain by Lemma 6.7 below a current $w_{0}$, which is the push-forward of a current on $Z$ such that
$\big{(}{\bf s}_{1}(E_{1})\wedge{\bf s}_{1}(E_{2})-{\bf s}_{1}(E_{2})\wedge{\bf s}_{1}(E_{1})\big{)}\wedge\mathbf{S}^{-}\wedge\mu=dd^{c}w_{0},$
where $w_{0}$ is current that is the push-forward of a current on $Z$. Note that $w_{0}$ has support on $Z$, and both ${\bf s}_{1}(E_{1})$ and ${\bf s}_{1}(E_{2})$ act by multiplication with a closed smooth form on currents in ${\mathcal{QZ}}$ with support on $Z$, and hence $\mathbf{S}^{+}$ acts as multiplication with a closed smooth form $S^{+}_{0}$, so we may take $w=S^{+}_{0}\wedge w_{0}$. ∎
Lemma 6.7.
If $E_{1},E_{2}$ are qas line bundles, $Z=Z_{E_{1}}\cap Z_{E_{2}}$ denote the intersection of the unbounded loci of $E_{1}$ and $E_{2}$. For any $\mu\in{\mathcal{QZ}}(X)$, there exists a current $w$, which is the push-forward of a current on $Z$ such that
$({\bf s}_{1}(E_{1})\wedge{\bf s}_{1}(E_{2})-{\bf s}_{1}(E_{2})\wedge{\bf s}_{1}(E_{1}))\wedge\mu=dd^{c}w,$
Proof.
By assuming that $\mu=\tau_{*}\gamma$, $\tau:W\to X$, and where $\gamma$ is of the form (4.1), we may assume by (6.9), after replacing $X$ by $W$, that $\mu$ is of the form (4.1). Let $\psi_{j}$ denote the metric on $E_{j}$, for $j=1,2$, and let $\omega_{j}$ denote the first Segre form of the smooth reference metric on $E_{j}$. Either some of $\psi_{j}\equiv-\infty$, say it is the case for $\psi_{1}$. Then ${\bf s}_{1}(E_{1})$ acts by multiplication with the smooth form $\omega_{1}$, which commutes with ${\bf s}_{1}(E_{2})$, and we may take $w_{0}=0$. Hence, we may assume that $\psi_{j}\not\equiv-\infty$.
As in the beginning of Section 4.1, we may perform a principalization of the singularities of $\psi_{j}$, and after a decomposition, we may assume that
$dd^{c}\psi_{j}=[D_{j}]+[W_{j}]+\omega_{j}^{\prime},$
where $D_{j},W_{j}$ are divisors for $j=1,2$, $|D_{1}|=|D_{2}|$, and $W_{1},W_{2},D_{1}$ intersect pairwise properly.
A calculation yields that
$({\bf s}_{1}(E_{1})\wedge{\bf s}_{1}(E_{2})-{\bf s}_{1}(E_{2})\wedge{\bf s}_{1}(E_{1}))\wedge\mu\\ =(([D_{1}]+[W_{1}]+\omega_{1}^{\prime})\wedge([W_{2}]+\omega_{2}^{\prime})+\omega_{1}\wedge[D_{2}]-([D_{2}]+[W_{2}]+\omega_{2}^{\prime})\wedge([W_{1}]+\omega_{1}^{\prime})+\omega_{2}\wedge[D_{1}])\wedge\mu\\ =([D_{1}]\wedge([W_{2}]+\omega_{2}^{\prime}-\omega_{2})-[D_{2}]\wedge([W_{1}]+\omega_{1}^{\prime}-\omega_{1}))\wedge\mu.$
We may choose qas metrics $h_{j}$ on ${\mathcal{O}}(D_{j})$ such that $\omega_{j}-[W_{j}]-\omega_{j}^{\prime}=s_{1}({\mathcal{O}}(D_{j}),h_{j})$, where the unbounded locus of $h_{j}$ equals $|W_{j}|$. Decompose $[D_{j}]=\sum d_{j,k}[D^{k}]$ for irreducible divisors $D_{k}$, and choose smooth reference metrics on each ${\mathcal{O}}(D^{k})$ with corresponding first Segre forms $\omega_{D^{k}}$. Then there is a qas function $u_{j}$ such that
$dd^{c}u_{j}=[W_{j}]+\omega_{j}^{\prime}+\sum_{k}d_{j,k}\omega_{D_{k}}-\omega_{j},$
and the set where $u_{j}$ is not locally bounded equals $|W_{j}|$. We may then take
$w=(u_{1}[D_{2}]-u_{2}[D_{1}])\wedge\mu,$
cf. the proof of Lemma 4.7. Note that indeed $w$ is the push-forward of a current of order 00 on $|D_{1}|=|D_{2}|=Z$. ∎
Proof of Theorem 1.4.
Items (i) to (v) are covered by Theorem 6.5 with $\mu=1$.
The decomposition (1.8) holds since $M^{E,h,h_{0}}$ is in ${\mathcal{QZ}}(X)$. The statement about the codimension of the non-vanishing multiplicities of $s_{k}^{\prime}$ and $N_{k}$ follows from Theorem 4.21. From Theorem 6.5 (vi) we know that $s(E,h,h_{0})$ and $s(E,h^{\prime},h_{0}^{\prime})$ in (vii) are in the same class in ${\mathcal{QB}}(X)$. Hence $M^{E,h,h_{0}}$ and $M^{E,h^{\prime},h_{0}^{\prime}}$ are as well by (4.20). By Proposition 4.19 they therefore have the same multiplicities.
We now prove that these numbers are positive. First assume that $\psi$ is psh. In a small open neighborhood ${\mathcal{U}}$ of a given point on $X$, $E={\mathcal{U}}\times{\mathbb{C}}^{r}_{\alpha}$, and therefore $p^{-1}{\mathcal{U}}={\mathcal{U}}\times{\mathbb{P}}({\mathbb{C}}^{r}_{\alpha})$. Assume we choose a trivial metric $h_{0}$ on $E$, which implies that $\omega_{\alpha}=s_{1}(L_{E},h_{0})$ is positive. Assume now that $\psi$ is psh. In particular this holds if the metric is inherited from a morphism $g\colon E\to F$ and $F$ has a trivial metric, since then $\psi=\log|g(x)\alpha|^{2}$, which is psh. In that case, $s(L_{E},e^{\psi},e^{\psi_{0}})$ will be a positive current, and hence, so will $s(E,h,h_{0})=p_{*}(s(L_{E},e^{\psi},e^{\psi_{0}}))$. Furthermore, all components in the decomposition $s(E,h,h_{0})=s^{\prime}(E,h)+S^{E,h}+N^{E,h,h_{0}}$ will be positive as well, and in particular, all their components will have non-negative Lelong numbers.
By Lemma 3.7 any qas metric $h$ on $E$ is locally equivalent to a psh metric as above, and therefore all the concerned currents for $E$ with the metric $h$ have non-negative multiplicities. ∎
Remark 6.8.
The reference to Lemma 3.7 in last argument means that for the positivity of the multiplicities in part (vi), we need $E$ to be qas and not only that $L_{E}$ is qas, cf. Remark 3.9. The remainder of the theorem holds assuming only that $L_{E}$ is qas. ∎
Remark 6.9.
Almost all of Theorem 1.4 and Theorem 6.5 has natural extensions to ${\mathbb{Q}}$-qas or ${\mathbb{R}}$-qas metrics. Indeed, if $h$ is ${\mathbb{Q}}$-qas or ${\mathbb{R}}$-qas, it follows readily that ${\bf s}(E)$ yields an endomorphism of ${\mathcal{QZ}}_{\mathbb{Q}}(X)$ or ${\mathcal{QZ}}_{\mathbb{R}}(X)$, cf. Remark 4.11. In that case, the multiplicities referred to in the version of Theorem 1.4 would be non-negative rational or real numbers, cf. Remark 4.22. ∎
Example 6.10.
With the notation in the proof of Lemma 4.12, i.e., with $\mu=\tau_{*}\gamma$, we have
(6.17) ${\bf s}_{k}(E,h,h_{0})\wedge\mu=\Pi_{*}(([dd^{c}(\tau^{\prime})^{*}\psi]_{(\tau^{\prime})^{*}\omega}^{k+r-1}\wedge(p^{\prime})^{*}\gamma),$
where the product on the right takes place on ${\mathbb{P}}(\tau^{*}E)$ and $\Pi=p\circ\tau^{\prime}=\tau\circ p^{\prime}$. The order of the factors in the product on ${\mathbb{P}}(\tau^{*}E)$ is important in general. To see (6.17) just notice that
$p_{*}([dd^{c}\psi]_{\omega}^{k+r-1})\wedge\tau^{\prime}_{*}(p^{\prime})^{*}\gamma)=p_{*}\tau^{\prime}_{*}([dd^{c}\psi]_{\omega}^{k+r-1})\wedge(p^{\prime})^{*}\gamma),$
and this holds because $\Gamma=(p^{\prime})^{*}\gamma$ is a product as in (4.1) and from (5.2) therefore
$[dd^{c}\psi]^{k+r-1}_{\omega}\wedge\tau^{\prime}_{*}\Gamma=\tau^{\prime}_{*}\big{(}[dd^{c}(\tau^{\prime})^{*}\psi]^{k+r-1}_{(\tau^{\prime})^{*}\omega}\wedge\Gamma\big{)}.$
Notice that
${\mathbb{P}}(\tau^{*}E)={\mathbb{P}}(E)\times_{X}W.$
Consider now (4.15) with $W={\mathbb{P}}(E)$ and $\tau=p$. Then we get the fiber diagram
(6.18) $\begin{array}[c]{cccc}Y&\stackrel{{\scriptstyle\pi_{1}}}{{\longrightarrow}}&{\mathbb{P}}(E_{1})\\ \downarrow\scriptstyle{\pi_{2}}&&\downarrow\scriptstyle{p_{1}}\\ {\mathbb{P}}(E_{2})&\stackrel{{\scriptstyle p_{2}}}{{\longrightarrow}}&X\end{array}$
with
(6.19) $Y={\mathbb{P}}(E_{1})\times_{X}{\mathbb{P}}(E_{2}),\quad\Pi=p_{1}\circ\pi_{1}=p_{1}\circ\pi_{2}\colon Y\to X,$
where $E_{1}$ and $E_{2}$ are distinct copies of $E\to X$. It follows from (6.2) that
${\bf s}_{\ell_{2}}(E,h,h_{0})\wedge s_{\ell_{1}}(E,h,h_{0})=\Pi_{*}\big{(}[dd^{c}\pi_{2}^{*}\psi_{2}]^{\ell_{2}+r-1}_{\pi_{2}^{*}\omega_{2}}\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r-1}_{\pi_{1}^{*}\omega_{1}}\big{)}=\\ (p_{2})_{*}([dd^{c}\psi_{2}]^{\ell_{2}+r-1}_{\omega_{2}}\wedge(\pi_{2})_{*}[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r-1}_{\pi_{1}^{*}\omega_{1}})=(p_{2})_{*}([dd^{c}\psi_{2}]^{\ell_{2}+r-1}_{\omega_{2}}\wedge p_{2}^{*}s_{\ell_{1}}(E,h,h_{0})),$
where $e^{\psi_{j}}$ and $e^{\psi_{j,0}}$ denote the metrics on $L_{E_{j}}$ induced by $h$ and $h_{0}$, and $\omega_{j}=s_{1}(L_{E_{j}},e^{\psi_{j,0}})$. Consider now $\mu=\tau_{*}\gamma$, where $\tau:W\to X$ and $\gamma\in{\mathcal{QZ}}(W)$. We may just as well consider $E_{1},\ldots,E_{k}$ to be $k$ copies of $E\to X$, $Y=W\times_{X}{\mathbb{P}}(E_{1})\times_{X}\cdots\times_{X}{\mathbb{P}}(E_{k})$, $\pi_{j}\colon Y\to{\mathbb{P}}(E_{j})$ and $\pi_{0}\colon Y\to W$ the natural projections. Then, similarly to above,
(6.20) ${\bf s}_{\ell_{k}}(E,h,h_{0})\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h,h_{0})\wedge\mu=\Pi_{*}([dd^{c}\pi_{k}^{*}\psi_{k}]_{\pi_{k}^{*}\omega_{k}}^{\ell_{k}+r-1}\wedge\cdots\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]_{\pi_{1}^{*}\omega_{1}}^{\ell_{1}+r-1}\wedge\pi_{0}^{*}\gamma\big{)}.$
In particular, our definition of product of Segre forms coincides with the product defined in [18, Section 6.1], cf. also Remark 6.2. ∎
Proposition 6.11.
Assume that $h$ is a qas metric on $E$ with empty unbounded locus. Let $h_{\epsilon}$ be the local regularization of $h$ from Proposition 6.4. If $\mu$ is in ${\mathcal{QZ}}$, then
$s_{\ell_{k}}(E,h_{\epsilon})\wedge\cdots\wedge s_{\ell_{1}}(E,h_{\epsilon})\wedge\mu\to{\bf s}_{\ell_{k}}(E,h)\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h)\wedge\mu.$
Proof.
Using notation from the previous example, and arguing as in the beginning of the proof of Proposition 6.4, it follows that if $\mu$ is in ${\mathcal{QZ}}$, then
${\bf s}_{\ell_{k}}(E,h)\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h)\wedge\mu=\Pi_{*}((dd^{c}\pi_{k}^{*}\psi_{k})^{\ell_{k}+r-1}\wedge\cdots\wedge(dd^{c}\pi_{1}^{*}\psi_{1})^{\ell_{1}+r-1}\wedge\pi_{0}^{*}\gamma\big{)}.$
Let $e^{\psi_{j,{\epsilon}}}$ denote the metrics corresponding to $h_{\epsilon}$. Then,
$s_{\ell_{k}}(E,h_{\epsilon})\wedge\cdots\wedge s_{\ell_{1}}(E,h_{\epsilon})\wedge\mu=\Pi_{*}((dd^{c}\pi_{k}^{*}\psi_{k,{\epsilon}})^{\ell_{k}+r-1}\wedge\cdots\wedge(dd^{c}\pi_{1}^{*}\psi_{1,{\epsilon}})^{\ell_{1}+r-1}\wedge\pi_{0}^{*}\gamma\big{)}.$
Letting ${\epsilon}\to 0$ yields the desired statement, just as in the proof of Proposition 6.4, see, e.g., [13, Theorem III.3.7]. ∎
7. Chern forms of qas bundles
Let $(E,h)$ be a qas bundle, $h_{0}$ a smooth reference metric, and let $\hat{\bf s}(E,h,h_{0})={\bf s}(E,h,h_{0})-I$. In this section, we will mostly drop $h$ and $h_{0}$ from the notation. In view of Theorem 6.5 (ii), $\hat{\bf s}(E):={\bf s}(E)-I={\bf s}_{1}(E)+{\bf s}_{2}(E)+\cdots$ is an endomorphism on ${\mathcal{QZ}}(X)$ of positive bidegree. Let $\hat{\bf s}(E)^{k}$ the composition of it $k$ times. We define the Chern endomorphism
(7.1) ${\bf c}(E,h,h_{0}):=\sum_{k=0}^{\infty}(-1)^{k}\hat{\bf s}(E,h,h_{0})^{k}=I-\hat{\bf s}(E,h,h_{0})+\hat{\bf s}(E,h,h_{0})^{2}+\cdots.$
Just as for the Segre form, the Chern endomorphism yields a Chern form $c(E,h,h_{0}):={\bf c}(E,h,h_{0})1\in{\mathcal{QZ}}(X)$. If $h$ is smooth, then $c(E,h,h_{0})$ equals the usual Chern form of $h$, cf. (2.4).
Proposition 7.1.
Let $E\to X$ be an qas bundle. Then
(i) ${\bf c}(E)\wedge{\bf s}(E)={\bf s}(E)\wedge{\bf c}(E)=I$.
The statements in Proposition 6.4 and (ii) to (vi) of Theorem 6.5 hold if ${\bf s}(E)$ is replaced by ${\bf c}(E)$.
Proof.
Clearly ${\bf s}(E)$ and $\hat{\bf s}(E)$ commute, and hence the computation
${\bf s}(E)\wedge{\bf c}(E)=(I+\hat{\bf s}(E))\sum_{k=0}^{\infty}(-1)^{k}\hat{\bf s}(E)^{k}=I,$
is legitimate, so that ${\bf s}(E)\wedge{\bf c}(E)={\bf c}(E)\wedge{\bf s}(E)=I$. The other equality in (i) follows in the same way.
The corresponding version of Proposition 6.4 with ${\bf s}(E)$ replaced by ${\bf c}(E)$ follows from Proposition 6.4 and Proposition 6.11.
It remains to prove the corresponding statements of Theorem 6.5 when ${\bf s}(E)$ is replaced by ${\bf c}(E)$. Part (ii) is obvious, and part (iii) for ${\bf c}(E)$ follows from (iii) for ${\bf s}(E)$.
To see (iv), notice that ${\bf s}(E,h,h_{0})-{\bf s}(E,h_{0})=\hat{\bf s}(E,h,h_{0})-\hat{\bf s}(E,h_{0})$ and that $\hat{\bf s}(E,h,h_{0})$ and $\hat{\bf s}(E,h_{0})$ commute. By Theorem 6.5 (iv), for each $k$, there are $w_{k}$ such that
$dd^{c}w_{k}=({\bf s}(E,h,h_{0})-{\bf s}(E,h_{0}))\wedge\sum_{j=0}^{k-1}\hat{\bf s}(E,h,h_{0})^{j}\wedge\hat{\bf s}(E,h_{0})^{k-1-j}\wedge\mu.$
If $w=-w_{1}+w_{2}-\cdots$, then $dd^{c}w=(s(E,h,h_{0})-s(E,h_{0}))\wedge\mu$.
We get (v) for ${\bf c}$ from the statement for ${\bf s}$ by a formal computation. For instance, by (i) and (v) for ${\bf s}$, $I={\bf s}(E^{\prime}\oplus E)\wedge{\bf c}(E^{\prime}\oplus E)=s(E^{\prime})\wedge{\bf s}(E)\wedge{\bf c}(E^{\prime}\oplus E)$. Multiplying by ${\bf c}(E)$ and ${\bf c}(E^{\prime})$ from the left we get ${\bf c}(E^{\prime})\wedge{\bf c}(E)={\bf c}(E^{\prime}\oplus E)$. Finally, we get (vi) for ${\bf c}$ by repeated use of (vi) for ${\bf s}$. ∎
We do not know if in general, $c_{k}(E)=0$ for $k>{\rm rank\,}E$, as it is in the smooth case.
7.1. Chern forms for qas metrics with empty unbounded locus
Let us now assume that $h$ is a qas metric with empty unbounded locus $Z$. If $Y$ is as in the end of Example 6.10, then just as in the proof of Proposition 6.11, one has a well-defined operator
(7.2) ${\bf s}_{\ell_{k}}(E,h)\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h)1=\Pi_{*}\big{(}(dd^{c}\pi_{k}^{*}\psi_{k})^{\ell_{k}+r-1}\wedge\cdots\wedge\pi_{1}^{*}(dd^{c}\psi_{1})^{\ell_{1}+r-1}\big{)},$
which is commuting in the factors, since it is the direct image of a Bedford-Taylor product on ${\mathbb{P}}(E)\times_{X}\cdots\times_{X}{\mathbb{P}}(E)$. In particular, one may define products of currents like $s_{\ell_{k}}(E,h)\wedge\cdots\wedge s_{\ell_{1}}(E,h)$. Therefore, there is a well-defined current
$c(E,h)=\frac{1}{s(E,h)}=1+(s_{1}(E,h)^{2}-s_{2}(E,h))+\cdots$
just as is in the smooth case. If $h_{\epsilon}$ is the regularization from Proposition 6.4, then by Proposition 6.11, $c(E,h)=\lim c(E,h_{\epsilon})$. In particular,
(7.3) $c_{k}(E,h)=0,\quad k>{\rm rank\,}E,$
since it holds for $c_{k}(E,h_{\epsilon})$.
We do not know whether it holds for a general qas metric $h$, i.e., when $h$ is not assumed to be locally bounded, that $c_{k}(E,h,h_{0})=0$ for $k>{\rm rank\,}E$.
7.2. An alternative endomorphism
As will be discussed in the final section, there does not appear to exist a canonical choice of Segre or Chern forms associated to a qas metric, except for in low degrees. In this subsection, we study another choice of an extension of a Segre form as an endomorphism of ${\mathcal{QZ}}(X)$, which in turn leads to a different choice of Chern form. This Chern form is the one which appears in [1].
Assume that $E$ is equipped with a qas metric $h$, and a reference metric $h_{0}$. In this section, we will mostly drop $h,h_{0}$ from the notation of $s(E,h,h_{0})$, and we will let $E_{0}$ denote the bundle $E$ equipped with the metric $h_{0}$, and hence write $s(E_{0}):=s(E,h_{0})$.
Let us introduce the endomorphism
(7.4) ${\bf s}(E,Z)\wedge\mu={\bf s}(E)\wedge{\bf 1}_{X\setminus Z}\mu+{\bf s}(E_{0})\wedge{\bf 1}_{Z}\mu,\quad\mu\in{\mathcal{QZ}}(X).$
If $Z=X$, then it is just ${\bf s}(E_{0})$ so it is of interest only when $Z$ has positive codimension, which we will assume from now on. Note that if we let $s(E,Z):={\bf s}(E,Z)1$, then
(7.5) $s(E,Z)=s(E)$
since ${\bf 1}_{Z}1=0$ and ${\bf 1}_{X\setminus Z}1=1$ as long as $Z$ has positive codimension.
Proposition 7.2.
Let $E\to X$ be an qas bundle and $\mu\in{\mathcal{QZ}}(X)$.
(7.6) $({\bf s}(E,Z)-{\bf s}(E))\wedge\mu=({\bf s}(E)-{\bf s}(E_{0}))\wedge{\bf 1}_{Z}\mu.$
(ii) There is a current $w$, which is the pushforward of a current on $Z$, such that
(7.7) $({\bf s}(E,Z)-{\bf s}(E))\wedge\mu=dd^{c}w.$
Proof.
The equation in (i) follows readily by definition of ${\bf s}(E,Z)$.
To prove (ii), note first that since ${\bf 1}_{Z}\mu$ is in ${\mathcal{QZ}}(X)$ and has support on $Z$, by Proposition 4.4, there is a current $\mu_{Z}\in{\mathcal{QZ}}(Z)$ such that ${\bf 1}_{Z}\mu=i_{*}\mu_{Z}$, where $i:Z\to X$ denotes the inclusion. By Theorem 6.5 (iv), there is $w^{\prime}\in{\mathcal{QZ}}(Z)$ such that
$({\bf s}(i^{*}E)-{\bf s}(i^{*}E_{0}))\wedge\mu_{Z}=dd^{c}w^{\prime}.$
To be precise, we have formulated Theorem 6.5 only for a complex manifold, but when want to apply it here on $Z$ which might be singular, i.e., a reduced analytic space. However, one may note either that the same result holds on reduced analytic spaces, with the same proof, or one might reduce to the case of manifolds by taking a resolution of singularities $\tau\colon Z^{\prime}\to Z$, finding the appropriate current on $Z^{\prime}$ and taking the push-forward to $Z$. Hence, (7.7) follows by (i) together with Theorem 6.5 (iii) by taking $w=i_{*}w^{\prime}$. ∎
It follows from (7.4) that ${\bf s}_{0}(E,Z)\wedge\mu=\mu$ so that ${\bf s}_{0}(E,Z)=I$, the identity on ${\mathcal{QZ}}(X)$. We can thus define
(7.8) ${\bf c}(E,Z)\wedge\mu=\sum_{k=0}^{\infty}(-1)^{k}\hat{\bf s}(E,Z)^{k}\wedge\mu,$
where $\hat{\bf s}(X,Z)={\bf s}(X,Z)-I$.
Proposition 7.3.
Let $E\to X$ be an qas bundle and $\mu\in{\mathcal{QZ}}(X)$.
(7.9) ${\bf c}(E,Z)\wedge{\bf s}(E,Z)={\bf s}(E,Z)\wedge{\bf c}(E,Z)={I}.$
(ii) For each $\mu\in{\mathcal{QZ}}(X)$ there is a current $w$, which is the pushforward of a current on $Z$, such that
$dd^{c}w=\big{(}{\bf c}(E,Z)-{\bf c}(E)\big{)}\wedge\mu.$
(iii) If $Z$ has positive codimension in $X$, then, if ${\bf c}^{\prime}(E):={\bf 1}_{X\setminus Z}{\bf c}(E)$,
(7.10) ${\bf c}(E,Z)=\big{(}{\bf c}^{\prime}(E)-{\bf c}(E_{0})\wedge M^{E}\wedge{\bf c}^{\prime}(E)\big{)}\wedge{\bf 1}_{X\setminus Z}\mu+{\bf c}(E_{0})\wedge{\bf 1}_{Z}\mu.$
By definition both $c(E)$ and $c(E,Z)$ are equal to the usual Chern form in $X\setminus Z$ if the metric on $E$ is smooth there. However, in general is $c(E,Z)\neq c(E)$ across $Z$, see Example 8.2.
Remark 7.4.
In view of (iii), $c(E,Z)$ is precisely the Chern form that appeared in [1] in case the metric on $E$ is obtained from $F$ via a generically injective $g\colon E\to F$, cf. Example 6.3
Proof of Proposition 7.3.
Part (i) follows in the same way as (i) in Proposition 7.1.
We may decompose
$\hat{\bf s}(E)^{k}-\hat{\bf s}(E,Z)^{k}=\sum_{\ell=0}^{k-1}{\bf s}(E,Z)^{k-\ell-1}\wedge({\bf s}(E)-{\bf s}(E,Z))\wedge{\bf s}(E)^{\ell}.$
By (7.7), for each $\ell$, there exists $w_{\ell}$, which is the push-forward of a current on $Z$, such that
$({\bf s}(E)-{\bf s}(E,Z))\wedge{\bf s}(E)^{\ell}\wedge\mu=dd^{c}w_{\ell}.$
Since $w_{\ell}$ has support on $Z$,
${\bf s}(E,Z)^{k-\ell-1}\wedge({\bf s}(E)-{\bf s}(E,Z))\wedge{\bf s}(E)^{\ell}\wedge\mu=s(E_{0})^{k-\ell-1}\wedge dd^{c}w_{\ell}=dd^{c}(s(E_{0})^{k-\ell-1}w_{\ell}),$
so we may take
$w=\sum s(E_{0})^{k-\ell-1}\wedge w_{\ell}$
to prove (ii).
For (iii), recall that $M^{E}\wedge\mu:={\bf 1}_{Z}{\bf s}(E)\wedge\mu$ and ${\bf s}^{\prime}(E)\wedge\mu={\bf 1}_{X\setminus Z}{\bf s}(E)\wedge\mu.$ We will prove by induction that
(7.11) $\hat{\bf s}(E,Z)^{k}=\hat{\bf s}^{\prime}(E)^{k}+\sum_{\ell=0}^{k-1}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell},\quad k=1,2,\ldots,$
when acting on ${\bf 1}_{X\setminus Z}\mu$. By definition it holds for $k=1$. Assume that it holds for $k$. When applying ${\bf s}(E)\wedge{\bf 1}_{X\setminus Z}+{\bf s}(E_{0})\wedge{\bf 1}_{Z}$ to the right hand side of (7.11) the first mapping only acts on the first term and gives $\hat{\bf s}(E)\wedge\hat{\bf s}^{\prime}(E)^{k}=\hat{\bf s}^{\prime}(E)^{k+1}+M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k}$. The second mapping neglects the first term and just gives another factor $s(E_{0})$ on the terms in the sum. Thus (7.11) holds. Thus
${\bf c}(E,Z)=I+\sum_{k=1}^{\infty}(-1)^{k}\Big{(}\hat{\bf s}^{\prime}(E)^{k}+\sum_{\ell=0}^{k-1}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell}\Big{)}=\\ {\bf c}^{\prime}(E)+\sum_{\ell=0}^{\infty}\sum_{k=\ell+1}^{\infty}(-1)^{k}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell}=\\ {\bf c}^{\prime}(E)-\sum_{\ell=0}^{\infty}(-1)^{\ell}{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf c}^{\prime}(E)={\bf c}^{\prime}(E)-{\bf c}(E_{0})\wedge M^{E}\wedge{\bf c}^{\prime}(E).$
The action on ${\bf 1}_{Z}\mu$ is just $\hat{\bf s}(E_{0})^{k}\wedge{\bf 1}_{Z}\mu$. Thus (iii) follows. ∎
Remark 7.5.
One can check directly that $c(E,Z)={\bf c}(E,Z)1$ is cohomologous with $c(E_{0})$. In fact, accepting that the operations are legitimate we have, if
$dd^{c}u=s(E)-s(E_{0})=M^{E}+s^{\prime}(E)-s(E_{0}),$
that
$-dd^{c}(c(E_{0})\wedge u\wedge c^{\prime}(E))=-c(E_{0})\wedge(s(E)-s(E_{0}))\wedge c^{\prime}(E)=c^{\prime}(E)-c(E_{0})\wedge M^{E}\wedge c^{\prime}(E)-c(E_{0}).$
8. Examples and remarks
In this final section, we discuss to which extent there is a canonical choice of $s(E,h)$ (and other characteristic forms, i.e., Chern forms $c(E,h)$) when $h$ is a qas metric. In case the unbounded locus $Z$ of $h$ is empty, there seems to be a canonical choice of Segre form $s(E,h)$, given, e.g., by Segre forms $s(E,h_{\epsilon})$ of appropriate decreasing regularizations $h_{\epsilon}$ of $h$ as ${\epsilon}\to 0$, cf. Proposition 6.4. When $Z\neq\emptyset$, the form $s(E,h,h_{0})$ will typically depend on $h_{0}$, see [17, Example 8.2]. In particular, it follows by Theorem 1.4 (i) that limits $s(E,h_{\epsilon})$ as ${\epsilon}\to 0$ might depend on the sequence $h_{\epsilon}$, even when $h_{\epsilon}$ is a sequence of Griffiths negative smooth metrics decreasing to $h$. Hence, there is in general no canonical choice of a Segre form $s(E,h)$ depending only on $h$.
To be a bit more precise, it follows from our results that in low enough degrees, there is a natural canonical choice of $s_{k}(E,h)$. Let $Z$ denote the set where $h$ is not locally bounded, and let $p={\rm codim\,}Z$. As long as $k\leq p$, $s_{k}(E,h,h_{0})$ is independent of $h_{0}$ by Remark 6.6 (i). On the other hand, [17, Example 8.2] mentioned above shows that $s_{k}(E,h,h_{0})$ might depend on $h_{0}$ for any $k>p$.
In case $k<p$, any definition of $s_{k}(E,h)$ as a closed current of order 00 is in fact uniquely determined by its restriction to $X\setminus Z$. Indeed, because the difference between two such choices is a normal $(k,k)$-current with support on an analytic subset of codimension $>k$, it must vanish, cf. e.g., [13, Corollary III.2.11]. In particular, $s_{k}(E,h)=s_{k}^{\prime}(E,h)$, which coincides with the currents considered by Xia (for more general classes of metrics), [21], based on using the non-pluripolar Monge-Ampère product.
In the case when $k=p$, one may by Theorem 1.4 and Proposition 6.4 find locally an appropriate decreasing sequence of smooth metrics $h_{\epsilon}$ such that $s_{p}(E,h_{\epsilon})\to s_{p}(E,h)$, independent of the choice of such a sequence. If the metric $h$ is furthermore Griffiths negative, then by [17], $s_{k}(E,h)$ may be obtained as $\lim_{{\epsilon}\to 0}s_{k}(E,h_{\epsilon})$ for any sequence of smooth Griffiths negative metrics $h_{\epsilon}$ decreasing to $h$ as ${\epsilon}\to 0$ for any $k\leq p$.
Similar statements as above will hold also for products
$s_{k_{t}}(E,h)\wedge\dots\wedge s_{k_{1}}(E,h):={\bf s}_{k_{t}}(E,h)\wedge\dots\wedge{\bf s}_{k_{1}}(E,h)1,$
depending on how $k=k_{1}+\dots+k_{t}$ compares to $p$. Such a product is furthermore robust in the sense that it is commutative in the factors provided $k\leq p$, which follows from Theorem 6.5 (vii), cf. Remark 6.6 (ii). Similarly, it follows by Proposition 7.3 (ii) that replacing any $s_{k_{j}}(E)$ by $s_{k_{j}}(E,Z)$ does not change the resulting current provided $k\leq p$. In particular, $c_{k}(E)=c_{k}(E,Z)$ for $k\leq p$. For $k>p$, we will see below that such an equality does typically not hold.
8.1. Difference between $c_{k}(E)$ and $c_{k}(E,Z)$
Example 8.1.
In general $c(E,Z)$ and $c(E)$ are different. The simplest example is when $Z=X$ and $Z^{\prime}\neq Z$. Then in general $s(E)$ may have a residue term, whereas $s(E,Z)=\omega$ and hence $c(E,Z)=c(E_{0})$. ∎
Example 8.2.
Let $g$ be the $2\times 2$-diagonal matrix, with entries $x_{1}$ and $x_{2}$, considered as a morphism between the trivial bundles $E=X\times{\mathbb{C}}^{2}_{\alpha}=F$, and $F$ is equipped with the trivial metric. With the trivial metric on $E$ as reference metric $h_{0}$ and the singular metric $h$ induced by $g$, we have
(8.1) $s_{1}(E)=[x_{1}=0]+[x_{2}=0].$
This follows by direct computation, or, e.g., from [1, Proposition 11.1]. From [1, Example 11.2] we have that $s_{2}(E)=[x_{1}=x_{2}=0]$.
Let us now compute $c(E)$ and $c(E,Z)$ for the same qas bundle $E$. To begin with $c_{1}(E)=c_{1}(E,Z)=-s_{1}(E)=-([x_{1}=0]+[x_{2}=0])$. A calculation yields that
(8.2) ${\bf s}_{1}(E)^{2}={\bf s}_{1}(E)\wedge([x_{1}=0]+[x_{2}=0])=2[x_{1}=x_{2}=0].$
For instance, if $\tau\colon{\mathbb{C}}\to{\mathbb{C}}^{2}$, $t\mapsto(t,0)$, then ${\bf s}_{1}(\tau^{*}E)=[t=0]$. Hence, by Theorem 6.5 (iii), ${\bf s}_{1}(E)\wedge[x_{2}=0]=\tau_{*}(s_{1}(\tau^{*}E))=\tau_{*}[t=0]=[x_{1}=x_{2}=0]$. From [1, Example 11.2] we have that $s_{2}(E)=[x_{1}=x_{2}=0]$. Thus
$c_{2}(E)=s_{1}(E)^{2}-s_{2}(E)=[x_{1}=x_{2}=0].$
Since $Z=\{x_{1}x_{2}=0\}$, and $s_{1}(E)=s_{1}(E,Z)$ has support on $Z$, cf. above and (7.5),
$s_{1}(E,Z)^{2}={\bf s}_{1}(E,Z)\wedge([x_{1}=0]+[x_{2}=0])=0,$
since ${\bf s}_{1}(E_{0})=0$. We conclude that
$c_{2}(E,Z)=-[x_{1}=x_{2}=0],$
which we notice is different from $c_{2}(E)$. ∎