### v2

# Segre forms of singular metrics on vector bundles and Lelong numbers

Abstract.
Let $E\to X$ be a holomorphic vector bundle. We consider a class of a singular Hermitian metrics on $E$ with analytic singularities that contains all Griffiths negative such metrics. One can define, given a smooth reference metric $h_{0}$, a current $s(E,h,h_{0})$ called the associated Segre form, which defines the expected Bott-Chern class and coincides with the usual Segre form of $h$ where it is smooth. We prove that $s(E,h,h_{0})$ is the limit of the Segre forms of a sequence of smooth metrics if the metric is smooth outside the degeneracy locus, and in general as a limit of Segre forms of metrics with empty degeneracy locus. One can also define an associated Chern form $c(E,h,h_{0})$. We prove that the Lelong numbers of $s(E,h,h_{0})$ and $c(E,h,h_{0})$ are integers if the singularities are integral, and non-negative for $s(E,h,h_{0})$.
1. Introduction
Inspired by the importance, during several decades, of singular metrics on line bundles in algebraic geometry, with origins in [12], several efforts have been done to study singular metrics on higher rank holomorphic vector bundles. In particular, such metrics appeared in the work of Berndtsson-Păun, [9], in connection with pseudoeffectivity of twisted relative canonical bundles. Contrary to the line bundle case, it is not clear what the curvature and connection matrix for a singular metric should be, cf. [20, Theorem 1.5]. This would appear to be an obstacle for defining associated Chern forms (or other characteristic forms, like Segre forms), since a standard definition is in terms of the curvature matrix. However, for appropriate metrics, it has still been possible to introduce such associated forms, or more precisely currents of order 0, see, e.g., [15], [16], [17], [18] and [21]. In this article, we further the study of such associated forms, and in slightly more general situations.
Assume that $E\to X$ is a holomorphic vector bundle over a complex manifold $X$ with a Griffiths negative singular metric $h$ with analytic singularities; this means that $\psi=\log|\cdot|^{2}_{h}$ is plurisubharmonic (psh) with analytic singularities on $E\setminus 0_{E}$, see Section 3. The latter condition means that locally $\psi=c\log|f|^{2}+b,$ where $f$ is a tuple of holomorphic functions, $c>0$ and $b$ is bounded. We say that the singularities are integral (rational/real) if $c$ is a positive integer (rational/real number). In [18] was introduced an associated Segre form [^1] We keep the notion ’Segre form’ although in the singular case $s(E,h,h_{0})$ is in general a current (of order 0). $s(E,h,h_{0})$, slightly depending on a smooth reference metric $h_{0}$ on $E$, that represents the expected Bott-Chern class, and coincides with the usual Segre form where the metric is smooth.
We will allow bundles with metrics such that $\psi=\log|\cdot|^{2}_{h}$ is quasi-plurisubharmonic (qpsh) with analytic singularities on $E$. If in addition $\psi$ has integral singularities, we say that the metric, and the bundle equipped with such a metric, is qas. If the singularities are rational or real we say that the bundle/metric is $\mathbb{Q}$}-qas or $\mathbb{R}$}-qas. We mainly consider (integral) qas metrics, which is practical in order to keep track on arithmetic information. There are analogous results for $\mathbb{Q}$}-qas and $\mathbb{R}$}-qas metrics, see Remark 6.10.
The definition of Segre form $s(E,h,h_{0})$ in [18] extends to the case when $h$ is a qas metric on $E\to X$ and $h_{0}$ is a smooth reference metric. Let $h_{\epsilon}=h+{\epsilon}h_{0}$, i.e.,
(1.1) $|\cdot|^{2}_{h_{\epsilon}}=|\cdot|^{2}_{h}+{\epsilon}|\cdot|^{2}_{h_{0}}.$
If $h$ is smooth outside the degeneracy locus, then $h_{\epsilon}$ are smooth and so the classical associated Segre forms $s(E,h_{\epsilon})$ are well-defined. In general $h_{\epsilon}$ is qas with empty degeneracy locus, and the same definition makes sense using Bedford-Taylor products, see (6.18). Our first main result, Theorem 1.3, in particular states that
(1.2) $s(E,h,h_{0})=\lim_{{\epsilon}\to 0}s(E,h_{\epsilon}).$
Example 1.1.
Assume that $g\colon E\to F$ is a holomorphic vector bundle morphism and $F$ has a smooth metric, or a qas metric with empty degeneracy locus. Then the metric $|\alpha|^{2}=|g(x)\alpha|^{2}$ on $E$ inherited from $F$ is qas, see Corollary 3.6. ∎
Example 1.2.
An example of a singular metric $e^{\psi}$ on a line bundle $L$ is given by $\psi=\log(|\sigma_{1}|^{2}+\cdots+|\sigma_{m}|^{2})$, where $\sigma_{j}$ are sections of $L^{-1}$. It is covered by Example 1.1 via the mapping $L\to X\times\oplus_{1}^{m}{\mathbb{C}}$, defined by $(\sigma_{1},\ldots,\sigma_{m})$. Such metrics often occur in applications to complex algebraic geometry. ∎
We have that $s(E,h,h_{0})=s_{0}(E,h,h_{0})+s_{1}(E,h_{0},h)+\cdots$, where $s_{k}(E,h,h_{0})$ are closed $(k,k)$-currents of order 0. They are not necessarily positive unless both $h$ and $h_{0}$ are Griffiths negative. There are anyway well-defined multiplicities $\operatorname{mult}_{x}s_{k}(E,h,h_{0})$ at each point $x$, defined in the same way as Lelong numbers, see Section 4. Here is our first main theorem:
Theorem 1.3.
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$, and let $Z$ be the degeneracy locus of $h$. Let $h_{0}$ be a smooth reference metric, and let $s(E,h,h_{0})$ denote the associated Segre form.
(i) If $h_{\epsilon}$ is the metric defined by (1.1), then (1.2) holds.
(ii) There is a current $w$ on $X$ such that
(1.3) $dd^{c}w=s(E,h,h_{0})-s(E,h_{0}).$
(iii) We have the decomposition
(1.4) $s(E,h,h_{0})=s^{\prime}(E,h)+M^{E,h,h_{0}},$
where $M^{E,h,h_{0}}$ is closed and has support on $Z$ and $s^{\prime}(E,h)={\bf 1}_{X\setminus Z}s(E,h,h_{0})$ is independent of $h_{0}$. If $h$ is smooth on $X\setminus Z$, then $s^{\prime}(E,h)$ coincides with the usual Segre form of $h$ there.
(iv) If $\pi\colon X^{\prime}\to X$ is a modification, then $\pi^{*}h$ is a qas metric on $\pi^{*}E$, and
(1.5) $\pi_{*}s(\pi^{*}E,\pi^{*}h,\pi^{*}h_{0})=s(E,h,h_{0}).$
(v) If $E^{\prime}\to X$ is another holomorphic vector bundle, equipped with a smooth metric $h^{\prime}$, then
(1.6) $s(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})=s(E^{\prime},h^{\prime})\wedge s(E,h,h_{0}).$
(vi) For each integer $k\geq 0$ there is a decomposition
(1.7) $M_{k}^{E,h,h_{0}}=S^{E,h}_{k}+N^{E,h,h_{0}}_{k},$
where $S_{k}^{E,h}$ is the Lelong current of an analytic cycle of codimension $k$, and the set of $x$ such that $\operatorname{mult}_{x}s_{k}^{\prime}(E,h)$ or $\operatorname{mult}_{x}N_{k}^{E,h,h_{0}}$ is nonvanishing, is contained in an analytic set of codimension $\geq k+1$. The multiplicities of $s_{k}^{\prime}(E,h)$, $S_{k}^{E,h}$ and $N_{k}^{E,h,h_{0}}$ are non-negative integers that are independent of $h_{0}$.
(vii) If $h^{\prime}$ is a qas metric on $E\to X$ comparable to $h$, then $\operatorname{mult}_{x}M_{k}^{E,h,h_{0}}=\operatorname{mult}_{x}M_{k}^{E,h^{\prime},h_{0}}$ and $\operatorname{mult}_{x}s^{\prime}_{k}(E,h)=\operatorname{mult}_{x}s_{k}^{\prime}(E,h^{\prime})$ for $x\in X$ and $k=0,1,2,\ldots$. In particular, $S_{k}^{E,h}=S_{k}^{E,h^{\prime}}$, $k=0,1,2,\ldots$.
If the metric is $\mathbb{Q}$}-qas ($\mathbb{R}$}-qas), then the same statements hold, except that the multiplicities in part (vi) are rational (real).
Remark 1.4.
(i) Equation (1.3) means that $s(E,h,h_{0})$ represents the Bott-Chern class of $E$.
(ii) The properties (ii) to (iv) are (at least implicitly) proved in [18].
(iii) Equality (1.6) implies that
$s^{\prime}(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})$ $=s(E^{\prime},h^{\prime})\wedge s^{\prime}(E,h,h_{0})\text{ and }$
$M^{E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0}}$ $=s(E^{\prime},h^{\prime})\wedge M^{E,h,h_{0}}.$
(iv) If $M_{k}^{E,h,h_{0}}$ is a positive current, (1.7) is precisely the Siu decomposition. ∎
The definition of a Griffiths negative singular metric was introduced in [9], and its dual metric is by definition Griffiths positive. Although most attention in the literature is paid to positive metrics, we focus in this article on qas metrics which are slightly more general than negative metrics with integral analytic singularities. This leads to simpler formulations, in particular less signs. By considering duals of such metrics, one obtains corresponding statements concerning Griffiths positive metrics, cf. Remark 6.2.
Recall that in algebraic geometry, characteristic classes of vector bundles such as the Chern class, are defined as endomorphisms on Chow groups, i.e., the $\mathbb{Z}$}-module of analytic cycles ${\mathcal{Z}}(X)$ modulo rational equivalence.
Inspired by [5] we introduce a $\mathbb{Z}$}-module ${\mathcal{QZ}}(X)$ of currents of order 0 that contains the module ${\mathcal{Z}}(X)$ of analytic cycles as a submodule, if cycles are identified with their associated Lelong currents. In particular the $(0,0)$-current 11 is an element in ${\mathcal{QZ}}(X)$. Any element ${\mathcal{QZ}}(X)$ of a fixed dimension has a well-defined multiplicity (Lelong number) that is an integer, at each point. We also introduce certain quotient classes ${\mathcal{QB}}(X)$ that, contrary to the Chow classes, keep local information such as multiplicities; this fact is crucial in this paper.
A second main result is that the Segre form $s(E,h,h_{0})$ extends to an endomorphism ${\bf s}(E,h,h_{0})$ on ${\mathcal{QZ}}(X)$ such that ${\bf s}_{0}(E,h,h_{0})=I$, the identity, and ${\bf s}(E,h,h_{0})1=s(E,h,h_{0})$. An important fact is that ${\bf s}(E,h,h_{0})$ induces a well-defined endomorphism ${\bf s}(E,h)$ of ${\mathcal{QB}}(X)$ that is independent of $h_{0}$. There is an analogue of Theorem 1.3 for ${\bf s}(E,h,h_{0})$, see Theorem 6.7. Since ${\bf s}(E,h,h_{0})-I$ has positive degree,
(1.8) ${\bf c}(E,h,h_{0})=\sum_{\ell=0}^{\infty}(-1)^{\ell}({\bf s}(E,h,h_{0})-{I})^{\ell}$
is a well-defined endomorphism such that
(1.9) ${\bf c}(E,h,h_{0}){\bf s}(E,h,h_{0})={\bf s}(E,h,h_{0}){\bf c}(E,h,h_{0})={I}.$
It is natural, cf. (2.4), to say that $c(E,h,h_{0}):={\bf c}(E,h,h_{0})1$ is the Chern form associated with $(E,h,h_{0})$. The same definition of $c(E,h,h_{0})$ appeared in [18] but expressed in a different way.
Theorem 1.5.
Let $(E,h,h_{0})$ be as in Theorem 1.3. Properties (ii)-(vii) of Theorem 1.3 hold if $s(E,h,h_{0})$, $s^{\prime}(E,h)$, and $M^{E,h,h_{0}}$ are replaced by $c(E,h,h_{0})$, $c^{\prime}(E,h)$, and ${\bf 1}_{Z}s(E,h,h_{0})$, except that the multiplicities are not non-negative.
The analogue of (1.2) is not true for $c(E,h,h_{0})$, see Section 8.
As already mentioned, the starting point for this paper is [18] and the ideas there. The regularization (1.1) relies on results in [6, 10]. The arithmetic analysis of the Lelong numbers relies on ideas in [1], and the $\mathbb{Z}$}-modules ${\mathcal{QZ}}(X)$ and ${\mathcal{QB}}(X)$ that are modelled after the somewhat smaller modules ${\mathcal{GZ}}(X)$ and ${\mathcal{B}}(X)$ in [5].
Acknowledgements
We are grateful to Håkan Samuelsson Kalm, Antonio Trusiani and David Witt Nyström for valuable discussions.
2. Preliminaries
Throughout this article $X$ is assumed to be a complex manifold. Given a holomorphic vector bundle $E\to X$ we let $p\colon{\mathbb{P}}(E)\to X$ denote the projectivization, so that the fiber above $x$ is the set of all lines [^2] There is another common convention in the literature, where the fiber instead is all hyperplanes through the origin. through the origin in $E_{x}$. Then the pullback $p^{*}E\to{\mathbb{P}}(E)$ has the tautological sub(line)bundle $L_{E}\to{\mathbb{P}}(E)$ whose fiber over $(x,[\alpha])\in{\mathbb{P}}(E)$ is the line in $E_{x}$ determined by $\alpha$.
If $E$ has a smooth Hermitian metric $h$, then $p^{*}E$ and hence $L_{E}$ inherit smooth Hermitian metrics. Notice that a local section $s$ of $L_{E}$ has the form $s=\tilde{s}(x,\alpha)\alpha$, where $\tilde{s}$ is a holomorphic function on $E\setminus 0_{E}$ which is 1-1-homogeneous in $\alpha$. Thus any local section can be identified with such an $\tilde{s}$. The induced norm is
(2.1) $|s|^{2}_{e^{\psi}}:=|\tilde{s}|^{2}|\alpha|^{2}_{h}.$
In a local frame for $L_{E}$, say $1/\alpha_{j}$, the section $s$ is represented by the function $\tilde{s}\alpha_{j}$, and from (2.1) we have that
$|s|^{2}_{e^{\psi}}=|\tilde{s}\alpha_{j}|^{2}e^{\log(|\alpha|_{h}^{2}/|\alpha_{j}|^{2})}.$
Thus the metric $e^{\psi}$ on $L_{E}$ with respect to the local frame $1/\alpha_{j}$ is given by
(2.2) $\psi=\log(|\alpha|^{2}_{h}/|\alpha_{j}|^{2}).$
Therefore $dd^{c}\log|\alpha|_{h}^{2}$ is a well-defined form on ${\mathbb{P}}(E)$ and $c_{1}(L_{E},e^{\psi})=-dd^{c}\log|\alpha|^{2}_{h}$ is the first Chern form. Moreover,
$s(L_{E},e^{\psi})=\frac{1}{c(L_{E},e^{\psi})}=\frac{1}{1+c_{1}(L_{E},e^{\psi})}=\sum_{\ell=0}^{\infty}(dd^{c}\log|\alpha|_{h}^{2})^{\ell}.$
The Segre form of $(E,h)$ is defined as
(2.3) $s(E,h)=p_{*}s(L_{E},e^{\psi})=p_{*}\sum_{\ell=0}^{\infty}(dd^{c}\log|\alpha|_{h}^{2})^{\ell},$
and it is a smooth form since $p$ is a proper submersion. We write $s(E,h)=\sum s_{k}(E,h)$, where $s_{k}(E,h)$ has bidegree $(k,k)$. It turns out that $s_{0}(E,h)=1$ and thus s(E,h)1s(E,h)-1 has positive bidegree. We get the Chern form
(2.4) $c(E,h)=\frac{1}{s(E,h)}=\sum_{\ell=0}^{\infty}(-1)^{\ell}(s(E,h)-1)^{\ell}.$
By [19], see also [14], this definition of $c(E,h)$ coincides with the definition as the sum of the elementary symmetric polynomials acting on ($i/2\pi$ times) the curvature tensor $\Theta(h)=D_{h}^{2}$ of $(E,h)$, where $D_{h}$ denotes the Chern connection associated with $h$. The Segre form is then defined as $s(E,h)=1/c(E,h)$.
2.1. Qpsh functions with analytic singularities
Let $Y$ be a complex manifold. Recall that a real-valued function $u$ is qpsh (quasi-plurisubharmonic) if locally $u=v+a$, where $v$ is psh and $a$ is smooth. As already mentioned, $u$ has (integral) analytic singularities if locally
(2.5) $u=c\log|f|^{2}+b,$
where $b$ is bounded and $c$ is a positive (integral) number.
Lemma 2.1.
Assume that $W$ is a connected complex manifold. If $\tau\colon W\to Y$ is holomorphic and $u$ is qpsh with (integral) analytic singularities on $Y$, then $\tau^{*}u$ is qpsh with (integral) analytic singularities in $W$. Possibly $\tau^{*}u$ is identically -$\infty$.
Lemma 2.2.
If $\psi_{1}$ and $\psi_{2}$ are qpsh (with integral analytic singularities), then $\psi_{1}+\psi_{2}$ and $\log(e^{\psi_{1}}+e^{\psi_{2}})$ are qpsh (with integral analytic singularities).
Lemma 2.1 follows directly from the definitions. Lemma 2.2 is well-known for psh $\psi_{j}$ (with integral analytic singularities). It is readily checked that it also holds for qpsh (with integral analytic singularities), using the characterization that functions $\psi_{j},\ j=1,2,$ are qpsh (with integral analytic singularities) if and only if locally there is a smooth function $\phi$ such that $\psi_{j}^{\prime}:=\psi_{j}+\phi,\ j=1,2,$ are psh (with integral analytic singularities).
3. qas bundles
We say that a function is qas if it is qpsh and has integral analytic singularities. We can assume then that $c=1$ in (2.5).
Definition 3.1.
A singular metric $h$ on a holomorphic vector bundle $E\to X$ is qas if the function $\alpha\mapsto\log|\alpha|^{2}_{h}$ is qas on the total space $E$. In that case $(E,h)$ is said to be qas. The degeneracy locus of $h$ is the set $Z\subset X$ of points $x$ such that $h(x)$ is degenerate.
Notice that a qas metric $h$ is non-degenerate if and only if $\alpha\mapsto\log|\alpha|^{2}_{h}$ is locally bounded on $E\setminus 0_{E}$.
Example 3.2.
Let $L$ be a line bundle with metric $h=e^{\psi}$. We claim that $(L,e^{\psi})$ is qas if and only if $\psi$ is qas on $X$. Given a local frame, $L=U\times{\mathbb{C}}_{\alpha}$ and $\log|\alpha|^{2}_{h(x)}=\log|\alpha|^{2}+\psi(x)$. If this function is qas on $L$, then by Lemma 2.1 its pullback to $U=U\times\{1\}$, i.e., $\psi$, is qas. Conversely, if $\psi(x)$ is qas  then $\log|\alpha|^{2}+\psi(x)$ is qas on $U\times{\mathbb{C}}_{\alpha}$, cf. Lemmas 2.1 and $2.$ ∎
Sometimes we suppress $h$ and say that $E$ is a qas bundle, and write $|\alpha|$ rather than $|\alpha|_{h}$.
Lemma 3.3.
(i) If $\tau\colon W\to X$ is holomorphic and $E\to X$ is qas, then $\tau^{*}E\to W$ is qas.
(ii) If $g\colon E\to F$ is a bundle morphism and $F$ is qas, then $E$ is qas if equipped with the induced metric $|\alpha|:=|g(x)\alpha|$.
(iii) If $(E_{j},h_{j})$ are qas bundles over $X$, $j=1,2$, then $(E_{1}\oplus E_{2},h_{1}\oplus h_{2})$ is qas.
(iv) If $E\to X$ has two qas metrics $h_{1}$ and $h_{2}$, then $h_{1}+h_{2}$ is a qas metric on $E$.
(v) If $L$ is a qas line bundle, and $E$ is a qas vector bundle, then $E\otimes L$ is qas.
Proof.
If $\tau\colon W\to X$, then $|\tau^{*}\alpha|^{2}_{\tau^{*}h}=\tau^{*}|\alpha|^{2}_{h}$ so (i) follows from Lemma 2.1. Part (ii) also follows from Lemma 2.1 by considering the holomorphic mapping $E\to F$, $(x,\alpha)\mapsto(x,g(x)\alpha)$.
Locally $E_{j}=X\times{\mathbb{C}}_{\alpha_{j}}^{r_{j}}$. If $\log|\alpha_{j}|^{2}$ is qas on $E_{j}$, $j=1,2$, then so is its pullback to $E_{1}\oplus E_{2}$. By Lemma 2.2, $\log(|\alpha_{1}|^{2}+|\alpha_{2}|^{2})$ is qas on $E_{1}\oplus E_{2}$. Thus (iii) holds. We get (iv) by applying (ii) and (iii) to $(E,h_{1})$, $(E,h_{2})$ and the mapping $E\to E\oplus E,\ \alpha\mapsto(\alpha,\alpha).$ For (v), note that if $e_{L}$ is a local frame for $L$ and $e^{\psi_{L}}$ the corresponding representation of the metric, then $\log|\alpha\otimes e_{L}|^{2}=\log|\alpha|^{2}+\psi_{L}$, so (v) follows from Lemma 2.2. ∎
Remark 3.4.
We say that a function is $\mathbb{Q}$${\mathbb{R}}-qas$ if it is qpsh with analytic singularities, and the local positive constants $c$ in (2.5) are rational/real. We say that a vector bundle is $\mathbb{Q}$${\mathbb{R}}-qas$ if $\alpha\mapsto\log|\alpha|^{2}$ is $\mathbb{Q}$${\mathbb{R}}-qas$. Lemma 2.1 holds for $\mathbb{Q}$}-qas and $\mathbb{R}$}-qas functions, and consequently Lemma 3.3 (i)-(ii) holds for $\mathbb{Q}$}-qas and $\mathbb{R}$}-qas bundles. Lemma 2.2 also holds for $\mathbb{Q}$}-qas functions, but not for $\mathbb{R}$}-qas functions in general. In fact, $\log(|f_{1}|^{2c_{1}}+|f_{2}|^{2c_{2}})$ does not have analytic singularities if $c_{1},c_{2}$ are arbitrary real numbers. Therefore, the analogues of Lemma 3.3 (iii)-(v) hold for $\mathbb{Q}$}-qas bundles, but not necessarily for $\mathbb{R}$}-qas bundles. ∎
Lemma 3.5.
Let $E$ be a holomorphic vector bundle on $X$ with a smooth Hermitian metric $h$. Any point $x_{0}\in X$ has a neighborhood $U$ such that $h^{\prime}=he^{A|x|^{2}}$ is Griffiths negative if $A$ is large enough, where $x$ denotes a local coordinate system near $x_{0}$.
Proof.
If the curvature $\Theta_{h}$ of $h$ is $\sum\Theta_{jk}dx_{j}{\wedge}d\bar{x}_{k}$, then $h$ is Griffiths negative if $\sum(\Theta_{jk}s,s)v_{j}\bar{v}_{k}\leq 0$ for any section $s$ of $E$ and tangent vector $v=\sum v_{j}\frac{\partial}{\partial x_{j}}$. Notice that $\Theta_{h^{\prime}}=\Theta_{h}-I_{E}A\omega_{x}$, where $\omega_{x}$ is the standard Kähler form on $U\subset{\mathbb{C}}^{N}_{x}$ and $I_{E}$ is the identity endomorphism on $E$. By possibly shrinking $U$, we may choose $A$ such that $|\sum(\Theta_{jk}s,s)v_{j}\bar{v}_{k}|\leq A|s|^{2}|v|^{2}$. Then $h^{\prime}$ is Griffiths negative. ∎
The statement means that locally there is a negative trivial line bundle $L_{A}\to U$ such that $E\otimes L_{A}$ is Griffiths negative. It is well-known that if $(E,h)$ is Griffiths negative, then $\log|\alpha|^{2}_{h}$ is psh on $E$, see, e.g., [9]. Thus we have
Corollary 3.6.
Any vector bundle with a smooth Hermitian metric is qas.
Lemma 3.7.
Assume that $h$ is a qas metric. For any $x_{0}\in X$ there is a neighborhood $U\subset X$, holomorphic $r$-tuples $f_{k}(x)$, $k=1,\ldots,m$, and $C>0$, such that in $E|_{U}\simeq U\times{\mathbb{C}}^{r}_{\alpha}$,
(3.1) $(1/C)|\alpha|_{h(x)}\leq\sum_{k=1}^{m}|f_{k}(x)\cdot\alpha)|\leq C|\alpha|_{h(x)}.$
Proof.
Let $x_{0}=0\in{\mathbb{C}}^{n}$ and let $|\alpha|$ be the standard norm on ${\mathbb{C}}^{r}_{\alpha}$. The assumption that $\log|\alpha|^{2}_{h(x)}$ has analytic singularities means that there are holomorphic functions $F_{k}(x,\alpha)$ in a set $|x|\leq c,|\alpha|\leq c,\ c>0$, such that (3.1) holds there with $F_{k}(x,\alpha)$ instead of $f_{k}(x)\cdot\alpha$. Since $F_{k}(x,0)=0$ we have that $F_{k}(x,\alpha)=f_{k}(x)\cdot\alpha+F^{\prime}_{k}$, where $F^{\prime}_{k}$ is ${\mathcal{O}}(|\alpha|^{2})$. We write the tuple as $F=f(x)\cdot\alpha+F^{\prime}$. For $x\leq c$, $|\alpha|=c$ and $|\lambda|\leq 1$, $F(x,\lambda\alpha)=\lambda f(x)\cdot\alpha+{\mathcal{O}}(|\lambda|^{2}).$ Thus
$|\lambda||\alpha|_{h(x)}\leq C|F(x,\lambda\alpha)|\leq C|\lambda||f(x)\cdot\alpha|+C|\lambda|^{2}.$
Letting $\lambda\to 0$ we get
$|\alpha|_{h(x)}\leq C|f(x)\cdot\alpha|,\quad|\alpha|=c,\ |x|\leq c.$
By homogeneity this holds for all $\alpha$. In the same way we get the opposite inequality. Thus (3.1) holds with $f(x)\cdot\alpha$ instead of $F(x,\alpha)$. Moreover, $f(x)\cdot\alpha$ is defined in the whole set $U\times{\mathbb{C}}^{r}$ and (3.1) holds there by homogeneity. ∎
In $U$ thus $|\cdot|_{h}^{2}$ is comparable to the metric in Example 1.1, if $F=U\times{\mathbb{C}}^{m}$ with trivial metric, and $g=(f_{1},\ldots,f_{m})$.
Proposition 3.8.
If $E\to X$ is qas, then $L_{E}\to{\mathbb{P}}(E)$ is qas.
Proof.
By Lemma 3.3 (i), $p^{*}E\to{\mathbb{P}}(E)$ is qas, and by (ii) its subbundle $L_{E}$ is qas. ∎
Remark 3.9.
Almost all results in this paper assuming that $E$ is qas only depend on that $L_{E}$ is qas. At a few places, Lemma 3.3 (ii)-(iii) and Lemma 3.7, and the non-negativity of multiplicities in Theorem 1.3, we use the (possibly) stronger assumption that $E$ is qas. Also if $L_{E}$ has a psh metric with analytic singularities, then the multiplicities are non-negative. Indeed, locally we can choose a negative reference metric; then $s_{k}(E,h,h_{0})$ is a positive $(k,k)$-current, and hence it has non-negative Lelong numbers. In [18], the standing assumption is that $E$ is equipped with a metric $h$ such that $L_{E}$ is psh with analytic singularities. By the arguments in the proof of [18, Proposition 5.4] it follows that $L_{E}$ is qas if and only if $\log|\alpha|_{h}^{2}$ is qas on $E\setminus 0_{E}$. Following the proof of [18, Proposition 5.2], one can show that if $L_{E}$ is qas, then $\log|\alpha|_{h}^{2}$ is qpsh on all of $E$. However we do not know if $\log|\alpha|_{h}^{2}$ has analytic singularities over $0_{E}$, so $E$ being qas is possibly a stronger assumption, unless ${\rm rank\,}E=1$, cf. Example 3.2. ∎
4. Generalized cycles and quasi-cycles
Let $Y$ be a complex manifold or a reduced analytic space. In [5] was introduced the $\mathbb{Z}$}-module ${\mathcal{GZ}}(Y)$ of generalized cycles, which are closed currents on $Y$. This module is an extension of the module ${\mathcal{Z}}(Y)$ of analytic cycles, i.e., there is an injective mapping $i\colon{\mathcal{Z}}(Y)\to{\mathcal{GZ}}(Y)$, if the elements in ${\mathcal{Z}}(Y)$ are identified with their associated Lelong currents. It was shown that many properties of ${\mathcal{Z}}(Y)$ such as unique decomposition into irreducible components, well-defined integer multiplicities at each point etc., hold for ${\mathcal{GZ}}(Y)$ as well.
Recall that if $Y$ is a reduced analytic space, then smooth forms on $Y$ are defined in terms of restrictions of smooth forms in local embeddings, see, e.g., [2], and currents are dual elements as in the smooth case.
Let ${\mathcal{QZ}}(Y)$ be the $\mathbb{Z}$-module of locally finite sums of currents of the form $\tau_{*}\gamma$, where $\tau\colon W\to Y$ is a proper holomorphic mapping, $W$ is a complex manifold, and $\gamma$ is a Bedford-Taylor product, [8],
(4.1) $\gamma=dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{p},$
on $W$, where $e^{b_{j}}$ are metrics on holomorphic line bundles $L_{j}$ over $W$ such that $b_{j}$ are locally bounded and qpsh. We say that the elements in ${\mathcal{QZ}}(Y)$ are quasi-cycles on $Y$.
Remark 4.1.
If we assume all $b_{j}$ be smooth we get precisely the module ${\mathcal{GZ}}(Y)$, cf. [5, Remark 3.1], which is thus a submodule of ${\mathcal{QZ}}(Y)$. ∎
Remark 4.2.
One can just as well consider the corresponding $\mathbb{Q}$}-module or $\mathbb{R}$}-module that we denote by ${\mathcal{QZ}}_{\mathbb{Q}}(Y)$ and ${\mathcal{QZ}}_{\mathbb{R}}(Y)$, respectively. ∎
The product (4.1) is a closed current of order 0 so it follows that any $\mu\in{\mathcal{QZ}}(Y)$ is a closed current of order 0. If $\tau\colon Y^{\prime}\to Y$ is a proper map and $\mu\in{\mathcal{QZ}}(Y^{\prime})$, then $\tau_{*}\mu\in{\mathcal{QZ}}(Y)$.
Most results about ${\mathcal{GZ}}(Y)$ in [5] extend to ${\mathcal{QZ}}(Y)$, with the same arguments, because of the following three properties of mixed Monge-Ampère products of the form (4.1):
Lemma 4.3.
Let $e^{b_{1}},\ldots,e^{b_{k}}$ be metrics on line bundles $L_{1},\dots,L_{k}$ on a complex manifold $W$ such that $b_{1},\dots,b_{k}$ are locally bounded and qpsh.
(i) If $i\colon V\to W$ is an analytic subspace of positive codimension, then
(4.2) ${\bf 1}_{V}dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{k}=0.$
(ii) If $i\colon V\to W$ is as in (i), then
(4.3) $dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{k}\wedge[V]=i_{*}(dd^{c}i^{*}b_{1}\wedge\cdots\wedge dd^{c}i^{*}b_{k}).$
(iii) If $W$ is compact and $k=\dim W$, then
(4.4) $\int_{W}dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{k}=\int_{W}c_{1}(L_{1})\wedge\cdots\wedge c_{1}(L_{k}).$
Here $c_{1}(L_{j})$ denotes the Chern class of $L_{j}$, which can be represented by the first Chern form associated with any smooth Hermitian metric on $L_{j}$.
Proof.
It is well-known that products like (4.1), do not charge pluripolar sets, see, e.g., [13, Proposition III.3.11]. In particular (4.2) holds. The equality (4.3) is a local statement and follows by a suitable regularization, see, e.g., [4, Lemma 3.4]. The last equality, (4.4), is well-known in case all $b_{j}$ are smooth. Here is a sketch of an argument in our case. If $\beta_{j}$ are smooth metrics, then $b_{j}-\beta_{j}$ are global functions. We prove by induction that (4.4) holds if $b_{1},\ldots,b_{\ell}$ are replaced by $\beta_{1},\ldots\beta_{\ell}$ for $1\leq\ell\leq k$. Assume it is true for $\ell-1$. Then by the commutativity of Monge-Ampère products it is enough to verify, with $u=b_{\ell}-\beta_{\ell}$, that
$dd^{c}(uT)=dd^{c}b_{\ell}\wedge T-dd^{c}\beta_{\ell}\wedge T$
if $T$ is any positive closed current. This is a local statement and follows from [13, Proposition III.4.9] since locally $u=\hat{b}_{\ell}-\hat{\beta}_{\ell}$, where $\hat{b}_{\ell}$ is psh and bounded, and $\hat{\beta}_{\ell}$ is smooth. Now the induction step follows from Stokes’ theorem. ∎
Using (4.2), it follows, cf. [5, Lemma 3.2], that if $V\hookrightarrow Y$ is an analytic subspace and $\mu\in{\mathcal{QZ}}(Y)$, then ${\bf 1}_{V}\mu\in{\mathcal{QZ}}(Y)$. Indeed, if
(4.5) $\mu=\sum(\tau_{j})_{*}\gamma_{j},$
where $\tau_{j}\colon W_{j}\to Y$, then
(4.6) ${\bf 1}_{V}\mu=\sum_{\tau_{j}(W_{j})\subset V}(\tau_{j})_{*}\gamma_{j}.$
Proposition 4.4.
Assume that $i\colon V\to Y$ is an analytic subspace. Then the image of $i_{*}\colon{\mathcal{QZ}}(V)\to{\mathcal{QZ}}(Y)$ is precisely the elements in ${\mathcal{QZ}}(Y)$ that have Zariski support on $V$.
The proof is the same as for the analogous statement for ${\mathcal{GZ}}(Y)$, [5, Proposition 3.6].
Example 4.5.
If $\mu\in{\mathcal{QZ}}(Y)$ has support at $x\in Y$, then since it has order 0 it must be a number times $[x]$, the point evaluation current at $x$. In view of Proposition 4.4 this number is a finite sum of numbers $\alpha=\int\tau_{*}\gamma$, where $\tau(W)=\{x\}$. Now
$\alpha=\int_{Y}\tau_{*}\gamma=\int_{W}\gamma,$
which is an intersection number, cf. (4.4), and hence an integer. ∎
4.1. Monge-Ampère type products as endomorphisms on ${\mathcal{QZ}}(Y)$
Assume that $u$ is a qas function on $Y$. Let $f$ be the tuple in a local representation (2.5) in $U\subset Y$ (with $c=1$) and let $\pi\colon U^{\prime}\to U$ be the normalization of the blowup of $Y$ along (the ideal generated by) $f$. If $\sigma$ is a section of the line bundle ${\mathcal{L}}\to U^{\prime}$ defining the exceptional divisor in $U^{\prime}$, then $\pi^{*}f=\sigma f^{\prime}$ where $f^{\prime}$ is a non-vanishing tuple of sections of ${\mathcal{L}}^{-1}$. If $\tilde{f}$ is the tuple from another representation (2.5) in $U$, then $|\tilde{f}|\sim|f|$ and therefore this modification is also the normalization of the blowup along $\tilde{f}$. This follows from the minimality property of such a modification. Therefore we get a global normal modification $p\colon Y^{\prime}\to Y$  such that any $f$ occurring in some local (2.5) is principal. (If wanted, by taking a further modification, one can assume that $Y^{\prime}$ is smooth but this is not necessary.) We thus get a global section $\sigma$ of a holomorphic line bundle ${\mathcal{L}}\to Y^{\prime}$ such that, if we equip $\mathcal{L}$} with some Hermitian norm,
(4.7) $\pi^{*}u=\log|\sigma|^{2}+a,$
where $a$ is qpsh and locally bounded. By the Poincaré-Lelong formula, $dd^{c}\log|\sigma|^{2}=[\operatorname{div}\sigma]-c_{1}({\mathcal{L}})$ and hence
(4.8) $dd^{c}\pi^{*}u=[\operatorname{div}\sigma]+\nu,$
where $\nu=-c_{1}({\mathcal{L}})+a$ is a global current on $Y^{\prime}$ of the form (4.1).
Remark 4.6.
Note that $\nu$ does not depend on the choice of Hermitian norm on $\mathcal{L}$}. One can in fact incorporate $a$ into the norm on $\mathcal{L}$}, i.e., choose a qas norm $\|\cdot\|^{2}$ on $\mathcal{L}$} (with empty degeneracy locus) so that $\pi^{*}u=\log\|\sigma\|^{2}$. ∎
Lemma 4.7.
Assume that $\gamma$ is a product on $W$ as in (4.1) and $f$ is a non-trivial holomorphic function. Then $(\log|f|^{2})\gamma$ has locally finite mass, and $dd^{c}((\log|f|^{2})\gamma)=[\operatorname{div}f]\wedge\gamma.$ If $u_{j}$ are smooth and decrease to $\log|f|^{2}$, then $u_{j}\gamma\to(\log|f|^{2})\gamma$.
Proof.
The first statement follows from [13, Theorem III.4.5] since locally in $Y^{\prime}$ all $b_{j}$ in $\gamma$ are $b_{j}=b_{j}^{\prime}+a_{j}$ where $b_{j}^{\prime}$ are psh and bounded and $a_{j}$ are smooth. The second statement follows from [13, Proposition III.4.9] and the third one from dominated convergence. ∎
Proposition 4.8.
Let $u$ be a qas function on $Y$ with degeneracy locus $Z$. If $\mu\in{\mathcal{QZ}}(Y)$, then $u{\bf 1}_{Y\setminus Z}\mu$ has order zero and locally finite mass and
(4.9) $dd^{c}u]\wedge\mu:=dd^{c}(u{\bf 1}_{Y\setminus Z}\mu)$
is an element in ${\mathcal{QZ}}(Y)$.
We interpret the right hand side of (4.9) as 0 if $\mu$ has support in $Z$. In particular, $dd^{c}u]\wedge\mu=0$ for all $\mu$ if $u\equiv$-$\infty$.
Proof.
By linearity we can assume that $\mu=\tau_{*}\gamma$, $\tau\colon W\to Y$. If the image of $\tau$ is contained in $Z$, then ${\bf 1}_{Y\setminus Z}\mu=0$. Otherwise $\tau^{*}u\not\equiv$-$\infty$ is qas on $W$ by Lemma 2.1. Then
${\bf 1}_{Y\setminus Z}\mu=\tau_{*}({\bf 1}_{W\setminus\tau^{-1}Z}\gamma)=\tau_{*}\gamma=\mu$
since $\tau^{-1}Z$ has positive codimension on $W$, cf. (4.2). In view of the discussion above we may assume, after possibly taking a modification of $W$, that $\tau^{*}u=\log|\sigma|^{2}+a$ on $W$, cf. (4.7), where $a$ is locally bounded.
Since $u$ is upper semi-continuous there is a sequence $u_{j}$ of smooth functions that decreases to $u$. By Lemma 4.7, $\tau^{*}u_{j}\gamma$ tends to the $\tau^{*}u\gamma$ that has locally finite mass. It follows that
$u\mu=\lim_{j}u_{j}\mu=\lim\tau_{*}(\tau^{*}u_{j}\gamma)=\tau_{*}(\tau^{*}u\gamma)$
is a locally finite measure. Moreover, $dd^{c}(u\mu)=\tau_{*}(dd^{c}(\tau^{*}u\gamma))$ so it is enough to see that $dd^{c}(\tau^{*}u\gamma)$ is in ${\mathcal{QZ}}(W)$. By (4.8) on $W$, and Lemma 4.7,
(4.10) $dd^{c}(\tau^{*}u\gamma)=[\operatorname{div}\sigma]\wedge\gamma+\nu\wedge\gamma.$
The last term in (4.10) is of the form (4.1). Notice that $\operatorname{div}\sigma]$ is a sum of terms $m_{\ell}[D_{\ell}]$, where $m_{\ell}$ are integers and $D_{\ell}$ are divisors in $W$. In view of Lemma 4.3 (ii), the first term on the right hand side of (4.10) is therefore a finite sum of terms $m_{\ell}(i_{\ell})_{*}(i_{\ell}^{*}\gamma)$, where $i_{\ell}\colon D_{\ell}\to W$, and each of them is in ${\mathcal{QZ}}(W)$. ∎
Recursively we define
(4.11) $dd^{c}u]^{0}\wedge\mu=\mu,\quad[dd^{c}u]^{k+1}\wedge\mu=[dd^{c}u]\wedge[dd^{c}u]^{k}\wedge\mu,\quad k=0,1,2,.$...
Let [^3] If $u=\log|\sigma|^{2}$, then $M^{u}$ is denoted by $M^{\sigma}$ in, e.g., [5]. We hope this will not cause any confusion.
(4.12) $\langle dd^{c}u\rangle^{k}\wedge\mu:={\bf 1}_{X\setminus Z}[dd^{c}u]^{k}\wedge\mu,\quad M^{u}_{k}\wedge\mu:={\bf 1}_{Z}[dd^{c}u]^{k}\wedge\mu.$
We have, cf. (4.9),
(4.13) $dd^{c}u]^{k+1}\wedge\mu=dd^{c}(u\langle dd^{c}u\rangle^{k}\wedge\mu).$
Let $M^{u}=M^{u}_{0}+M^{u}_{1}+\cdots$. Notice that if $\mu$ has support on $Z$, then $M^{u}_{0}\wedge\mu=\mu$.
Example 4.9.
With the notation in (4.10) in the proof of Proposition 4.8 we have
(4.14) $dd^{c}\tau^{*}u]^{k}\wedge\gamma=M_{k}^{\tau^{*}u}\wedge\gamma+\langle dd^{c}\tau^{*}u\rangle^{k}\wedge\gamma=[\operatorname{div}\sigma]\wedge\nu^{k-1}\wedge\gamma+\nu^{k}\wedge\gamma.$
Proposition 4.10.
If $\tau\colon Y^{\prime}\to Y$ is a proper mapping and $\mu\in{\mathcal{QZ}}(Y^{\prime})$, then
(4.15) $dd^{c}u]^{k}\wedge\tau_{*}\mu=\tau_{*}\big([dd^{c}\tau^{*}u]^{k}\wedge\mu\big),\quad\langle dd^{c}u\rangle^{k}\wedge\tau_{*}\mu=\tau_{*}\big(\langle dd^{c}\tau^{*}u\rangle^{k}\wedge\mu\big),\\ \quad M_{k}^{u}\wedge\tau_{*}\mu=\tau_{*}\big(M_{k}^{\tau^{*}u}\wedge\mu\big),\quad k=0,1,2,\ldots.$
Proof.
The case $k=0$ is trivial, and by induction it is enough to consider $k=1$. Then the proposition follows from the equalities ${\bf 1}_{Z}\tau_{*}\mu=\tau_{*}({\bf 1}_{\tau^{-1}Z}\mu)$ and
$dd^{c}(u_{j}{\bf 1}_{Y\setminus Z}\tau_{*}\mu)=\tau_{*}\big(dd^{c}(\tau^{*}u_{j}{\bf 1}_{Y^{\prime}\setminus\tau^{-1}Z}\mu)\big),$
where $u_{j}$ is a sequence of smooth functions as in the proof of Proposition 4.8. ∎
Remark 4.11.
We claim that if $u$ is $\mathbb{Q}$-$\mathbb{R}-qas$  then there is a modification $\pi\colon Y^{\prime}\to Y$ and a section $\sigma$ of a line bundle ${\mathcal{L}}\to Y^{\prime}$ such that locally on $Y^{\prime}$, $\pi^{*}u=c\log|\sigma|^{2}+a$, where $a$ is qpsh and $c>0$ is rational/real. Notice that $c$ cannot be globally constant in general. For instance, let $u$ be a potential to $\sum_{k}(1/k)[z=k]$ in the complex plane.
To see the claim, first recall that the normalization $p\colon U^{\prime}\to U$ of the blowup of $U$ along the ideal $\mathcal{J}$} generated by $f$ is the same as the blow up of the ideal ${\mathcal{J}}^{k}$ for any $k\geq 1$. If $f$ and $\hat{f}$ occur in two local representations (2.5) then $c^{\prime}\log|f^{\prime}|=c\log|f|+{\mathcal{O}}(1)$. Therefore $|f^{\prime}|^{k^{\prime}}\sim|f|^{k}$ for some integers $k,k^{\prime}\geq 1$. This is obvious if $u$ is $\mathbb{Q}$-qas so that $c,c^{\prime}\in\mathbb{Q}$. In the real case this is clear if $f$ and $f^{\prime}$ are functions, and one is reduced to this case in a suitable modification where $f$ and $f^{\prime}$ are principal. Thus $Y^{\prime}$ is well-defined locally and hence globally. Now the claim follows if $\sigma$ is a global section that defines the exceptional divisor. ∎
4.2. Action of Segre and Chern forms associated with a smooth metric
Notice that if $L\to X$ is a line bundle with a smooth Hermitian metric $e^{\psi}$, then $c_{1}(L,e^{\psi})$ defines an endomorphism ${\bf c}_{1}(L,e^{\psi})$ on currents acting as multiplication by the smooth form $c_{1}(L,e^{\psi})$. Moreover, if $\mu=\tau_{*}\gamma$, then $c_{1}(L,e^{\psi})\wedge\tau_{*}\gamma=\tau_{*}(c_{1}(\tau^{*}L,\tau^{*}e^{\psi})\wedge\gamma)$ and hence ${\bf c}_{1}(L,e^{\psi})$ yields an endomorphism on ${\mathcal{QZ}}(X)$. We also have the endomorphism ${\bf s}_{1}(L,e^{\psi})=-{\bf c}_{1}(L,e^{\psi})$.
Let $E\to X$ be a Hermitian holomorphic vector bundle. Since $p\colon{\mathbb{P}}(E)\to X$ is a submersion, there is a well-defined morphism $p^{*}$ mapping currents on $X$ to currents on ${\mathbb{P}}(E)$ such that $p_{*}\xi\wedge\mu=p_{*}(\xi\wedge p^{*}\mu)$ for smooth forms $\xi$ on ${\mathbb{P}}(E)$.
Lemma 4.12.
We have
$p^{*}\colon{\mathcal{QZ}}(X)\to{\mathcal{QZ}}({\mathbb{P}}(E)).$
Proof.
By functoriality we have the commutative diagram
(4.16) $\begin{array}[c]{cccc}{\mathbb{P}}(\tau^{*}E)&\stackrel{{\scriptstyle\tau^{\prime}}}{{\longrightarrow}}&{\mathbb{P}}(E)\\ \downarrow\scriptstyle{p^{\prime}}&&\downarrow\scriptstyle{p}\\ W&\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}&X.\end{array}$
If $\mu=\tau_{*}\gamma$, thus $p^{*}\mu=\tau^{\prime}_{*}(p^{\prime})^{*}\gamma$, which is an element in ${\mathcal{QZ}}({\mathbb{P}}(E))$. ∎
Since $s(E,h)=p_{*}s(L_{E},e^{\psi})$, for any current $\mu$ on $X$,
(4.17) $s(E,h)\wedge\mu=p_{*}s(L_{E},e^{\psi})\wedge\mu=p_{*}(s(L_{E},e^{\psi})\wedge p^{*}\mu).$
It follows from the lemma that (4.17) defines endomorphisms
(4.18) ${\bf s}_{k}(E,h)\colon{\mathcal{QZ}}(X)\to{\mathcal{QZ}}(X),\quad k=0,1,\ldots,$
where ${\bf s}_{k}(E,h)\wedge\mu:=p_{*}(s_{1}(L_{E},e^{\psi})^{k+r-1}\wedge p^{*}\mu)$. In the same way we have endomorphisms
(4.19) ${\bf c}_{k}(E,h)\colon{\mathcal{QZ}}(X)\to{\mathcal{QZ}}(X),\quad k=0,1,\ldots$
defined by suitable compositions, cf. (2.4). Since these operators are just multiplication by $s_{k}(E,h)$ and $c_{k}(E,h)$, in particular ${\bf c}_{0}(E,h)={\bf s}_{0}(E,h)=I$, the identity on ${\mathcal{QZ}}(X)$.
4.3. The submodule ${\mathcal{QZ}}_{0}(Y)$
Let $0\to S\to E\to Q\to 0$ be a short exact sequence of holomorphic vector bundles over a complex manifold $W$ with the metrics induced by a smooth metric on $E$. Then, by [11], for each component $\beta=(c(E)-c(S)c(Q))_{k}$ there is smooth form $v$ on $W$ such that $dd^{c}v=\beta$. It follows from (4.19) that $\beta$ induces an endomorphism ${\mathcal{QZ}}(W)\to{\mathcal{QZ}}(W)$, $\mu\mapsto\beta\wedge\mu$.
Let $L$ be a holomorphic line bundle over $W$ with two different metrics $e^{b}$ and $e^{b^{\prime}}$, where $b^{\prime}$ is assumed to be smooth, and $b$ is qpsh and locally bounded. Then $v=b-b^{\prime}$ is a locally bounded qpsh function. Let $\beta:=dd^{c}b^{\prime}-dd^{c}b=dd^{c}v$. Again we have a mapping ${\mathcal{QZ}}(W)\to{\mathcal{QZ}}(W)$, $\mu\mapsto\beta\wedge\mu$. We say that any of these two forms $\beta$ is a beta form.
Definition 4.13.
We let ${\mathcal{QZ}}_{0}(Y)$ be the submodule of ${\mathcal{QZ}}(Y)$ generated by all currents of the form $\tau_{*}\mu^{\prime}$, $\tau\colon W\to Y$, where $\mu^{\prime}$ is in the image of one of the mappings above.
Thus ${\mathcal{QZ}}_{0}(Y)$ is generated by all $\tau_{*}(\beta\wedge\gamma)$ where $\beta$ is a beta form and $\gamma$ is a product as in (4.1). For each such form there is a locally bounded function $v$ such that
(4.20) $dd^{c}\tau_{*}(v\gamma)=\tau_{*}(\beta\wedge\gamma).$
Remark 4.14.
Since the pullback of a beta form is a beta form, the module ${\mathcal{QZ}}_{0}(Y)$ is actually generated by all currents $\tau_{*}\gamma$, where $\gamma$ is as in (4.1) but with at least one factor $dd^{c}b_{j}$ ’replaced’ by a beta form. ∎
If $\tau\colon Y^{\prime}\to Y$ is a proper mapping, then $\tau_{*}\colon{\mathcal{QZ}}_{0}(Y^{\prime})\to{\mathcal{QZ}}_{0}(Y)$. In view of (4.6), for any analytic subspace $V\subset Y$,
(4.21) ${\bf 1}_{V}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y).$
Proposition 4.15.
If $\mu\in{\mathcal{QZ}}_{0}(Y)$ has bidegree $(k,k)$ and support on the analytic subspace $V\subset Y$, then $\mu=0$ if $k-1<{\rm codim\,}V$.
In particular, if $\mu\in{\mathcal{QZ}}_{0}(Y)$ has support at a point, then $\mu=0$.
Proof.
It follows from Proposition 4.4 and (4.20) that there is a current $w=i_{*}w^{\prime}$, where $i\colon V\to X$, such that $dd^{c}w=\mu$. If $\mu$ has bidegree $(k,k)$, then $w$ has bidegree (k1,k1)(k-1,k-1). Thus $w^{\prime}$ has bidegree $(k-1-{\rm codim\,}V,k-1-{\rm codim\,}V)$ and hence it vanishes if $k-1<{\rm codim\,}V$. ∎
Proposition 4.16.
(i) If $u$ is a qas function on $Y$, then
$dd^{c}u]^{k}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y),\quad k=0,1,2,\ldots.$
(ii) If $u^{\prime}$ is qas and $u^{\prime}=u+{\mathcal{O}}(1)$ locally on $Y$, then
$dd^{c}u]^{k}-[dd^{c}u^{\prime}]^{k}\colon{\mathcal{QZ}}(Y)\to{\mathcal{QZ}}_{0}(Y),\quad k=0,1,2,\ldots.$
(iii) The mapping $p^{*}$ in Lemma 4.12 maps ${\mathcal{QZ}}_{0}(X)\to{\mathcal{QZ}}_{0}({\mathbb{P}}(E))$.
Proof.
Recall that if $\mu$ has support in $Z$, then $dd^{c}u]\wedge\mu=0$. Otherwise, assume that $\mu=\tau_{*}(\beta\wedge\gamma)$ where $\gamma$ is product as in (4.1) and $\beta$ is a beta form. As in the proof of Proposition 4.8 we can therefore assume, cf. (4.10), that
(4.22) $dd^{c}u]\wedge\mu=\tau_{*}\big([\operatorname{div}\sigma]\wedge\beta\wedge\gamma+\nu\wedge\beta\wedge\gamma\big).$
If $i\colon D\to W$, then $i^{*}(\beta\wedge\gamma)=i^{*}\beta\wedge i^{*}\gamma$ and thus $\operatorname{div}\sigma]\wedge\beta\wedge\gamma$ is in ${\mathcal{QZ}}_{0}(W)$. Thus (i) holds for $k=1$ and hence for a general $k$ by induction.
By (i) and a simple induction over $k$ it is enough to prove (ii) for $k=1$. Let us assume that $\mu=\tau_{*}\gamma$, $\tau\colon W\to Y$. After a suitable further modification we can assume that $\tau^{*}u=\log|\sigma|^{2}+a$ and $\tau^{*}u^{\prime}=\log|\sigma^{\prime}|^{2}+a^{\prime}$ in $W$ where $a$ and $a^{\prime}$ are locally bounded. The assumption in (ii) implies that $\operatorname{div}\sigma^{\prime}]=[\operatorname{div}\sigma]$. If $a^{\prime\prime}$ is some smooth metric on the dual of the line bundle corresponding to this divisor, then $dd^{c}(a-a^{\prime})=dd^{c}(a-a^{\prime\prime})-dd^{c}(a^{\prime}-a^{\prime\prime})=\beta-\beta^{\prime}$ where both $\beta$ and $\beta^{\prime}$ are beta forms on $W$. Therefore,
$([dd^{c}u^{\prime}]-[dd^{c}u])\wedge\mu=\tau_{*}(dd^{c}\tau^{*}u\wedge\gamma)-\tau_{*}(dd^{c}\tau^{*}u^{\prime}\wedge\gamma)=\tau_{*}(\beta\wedge\gamma)-\tau_{*}(\beta^{\prime}\wedge\gamma)$
which is in ${\mathcal{QZ}}_{0}(Y)$.
Part (iii) follows from the proof of Lemma 4.12. In fact, if $\mu$ is in ${\mathcal{QZ}}_{0}(X)$, then we can assume that $\mu=\tau_{*}(\beta\wedge\gamma)$ where $\gamma$ is of the form (4.1) and $\beta$ is a beta form. Then $p^{*}\mu=\tau^{\prime}_{*}(p^{\prime})^{*}(\beta\wedge\gamma)$ is in ${\mathcal{QZ}}_{0}({\mathbb{P}}(E))$. ∎
Definition 4.17.
We define the quotient module ${\mathcal{QB}}(Y):={\mathcal{QZ}}(Y)/{\mathcal{QZ}}_{0}(Y)$.
The definition is analogous with the definition of the quotient module ${\mathcal{B}}(Y)$ of ${\mathcal{GZ}}(Y)$, see [5], and there is a natural mapping $\mathcal{{\mathcal{B}}}(Y)\to{\mathcal{QB}}(Y)$.
4.4. Multiplicities of quasi-cycles
Let $\xi$ be a tuple of holomorphic functions that defines the maximal ideal at $x\in Y$. If $\mu\in{\mathcal{QZ}}(Y)$ and $u=\log|\xi|^{2}$, then $M^{u}\wedge\mu$ is in ${\mathcal{QZ}}(Y)$ and has support at $x$, so it is $\alpha[x]$ for some integer $\alpha$, cf. Example 4.5. By Proposition 4.16 (ii) and Proposition 4.15, $\alpha$ is independent of the choice of tuple $\xi$ defining the maximal ideal. If $\mu$ has bidegree $(k,k)$, for degree reasons, we have that $M^{u}\wedge\mu=M^{u}_{n-k}\wedge\mu$.
Definition 4.18.
Let $\mu\in{\mathcal{QZ}}(Y)$ be a $(k,k)$-current and $x\in Y$. We define the multiplicity of $\mu$ at $x$, $\operatorname{mult}_{x}\mu$, as the integer
(4.23) $\operatorname{mult}_{x}\mu[x]=\int M^{u}\wedge\mu=\int{\bf 1}_{\{x\}}(dd^{c}\log|\xi|^{2})^{n-k}\wedge\mu.$
If $\mu$ is a positive closed current, (4.23) means that $\operatorname{mult}_{x}\mu$ is the Lelong number of $\mu$ at $x$.
If $\mu\in{\mathcal{QZ}}_{0}(Y)$, then $M^{u}\wedge\mu\in{\mathcal{QZ}}_{0}(Y)$ by Proposition 4.16 (i), and hence $\operatorname{mult}_{x}\mu=0$, cf. Proposition 4.15. That is,
Proposition 4.19.
The multiplicity $\operatorname{mult}_{x}\mu$ only depends on the class of $\mu$ in ${\mathcal{QB}}(Y)$.
Theorem 4.20.
Assume that $\mu\in{\mathcal{QZ}}(Y)$. If $\mu$ has bidegree $(k,k)$, then there is a unique decomposition
(4.24) $\mu=\mu_{F}+\mu_{M},$
where $\mu_{F}$ is the Lelong current of an analytic cycle on $Y$ of codimension $k$, and the set where $\operatorname{mult}_{x}\mu_{M}$ does not vanish is contained in an analytic set of codimension $\geq k+1$.
If $\mu^{\prime}$ defines the same class in ${\mathcal{QB}}(Y)$, then, cf. Proposition 4.19, it has the same multiplicities. It follows from the theorem that $\mu^{\prime}_{F}=\mu_{F}$. Clearly $\mu^{\prime}_{M}$ and $\mu_{M}$ may differ, but they still have the same multiplicities at each point. We say that $\mu_{F}$ is the fixed part and $\mu_{M}$ is the moving part of the class of $\mu$ in ${\mathcal{QB}}(Y)$.
Proof.
Since we know that $\operatorname{mult}_{x}\mu$ is integer valued, (4.24) follows directly from the Siu decomposition in case $\mu$ is a positive current. As in the case with $\mu\in{\mathcal{GZ}}(Y)$, if $\mu=\tau_{*}\gamma$, and $\tau=i\circ\tau^{\prime}$, where $\tau^{\prime}\colon W\to Z$ is surjective, then $\operatorname{mult}_{x}\tau^{\prime}_{*}\gamma=\operatorname{mult}_{i(x)}\tau_{*}\gamma$. Moreover, [1, Lemma 7.3] holds, with the same proof, for $\mu=\tau_{*}\gamma$ if $\gamma$ is as in (4.1). In fact, the crucial point is that $M^{\tau^{*}\xi}\wedge\gamma=0$ if $M^{\tau^{*}\xi}=M^{\tau^{*}\xi}\wedge 1=0$. To see this we can assume that $\xi$ is a tuple of functions (i.e., a section of a vector bundle with trivial metric) in a neighborhood of $x$. Then $M^{\tau^{*}\xi}$ is a positive current. If $b_{\ell}$ is locally of the form $u_{\ell}+g_{\ell}$, where $u_{\ell}$ is psh, and $g_{\ell}$ is smooth, we may take a sequence $b_{\ell}^{j}=u_{\ell}^{j}+g_{\ell}$, where each $u_{\ell}^{j}$ is psh, and $u_{\ell}^{j}$ decreases to $u_{\ell}$. Then, by [13, Proposition III.4.9], $M^{\tau^{*}\xi}\wedge\gamma_{j}\to M^{\tau^{*}\xi}\wedge\gamma$. If $M^{\tau^{*}\xi}=0$, thus $M^{\tau^{*}\xi}\wedge\gamma_{j}=0$ and hence $M^{\tau^{*}\xi}\wedge\gamma=0$. Now Theorem 4.20 follows precisely as [1, Theorem 7.1]. ∎
Remark 4.21.
It follows that if $\mu\in{\mathcal{QZ}}_{\mathbb{Q}}(Y)$, then $\operatorname{mult}_{x}\mu$ is a rational number. ∎
Example 4.22.
Assume that $\mu\in{\mathcal{QZ}}(Y)$ has bidegree $(k,k)$ and support on an analytic subspace $i\colon V\to Y$ of pure codimension $q$. In view of Proposition 4.4, $\mu=i_{*}\tau$ for a $\tau\in{\mathcal{QZ}}(V)$ of bidegree (kq,kq)(k-q,k-q). Thus $\mu=0$ if $k<q$. If $k=q$, then $\tau$ is a locally constant function which is integer-valued since its multiplicities are integers. It follows that $\mu$ is a cycle, i.e., in ${\mathcal{Z}}(Y)$. If $\mu$ is in ${\mathcal{QZ}}_{0}(Y)$, then it vanishes by Proposition 4.15. Thus ${\mathcal{Z}}(Y)$ is a submodule of ${\mathcal{QB}}(Y)$. ∎
5. Regularization of qas metrics on a line bundle
Let $L\to Y$ be a holomorphic vector bundle, let $e^{\psi}$ be a qas metric on $L$, and let $e^{\psi_{0}}$ be a smooth reference metric. Then, cf. Example 3.2, the global function $\psi-\psi_{0}$ is qas on $Y$. In view of Proposition 4.8 the following definition makes sense.
Definition 5.1.
Consider a qas metric $e^{\psi}$ on $L\to X$, a smooth reference metric $e^{\psi_{0}}$, and $\omega=dd^{c}\psi_{0}$. We introduce the endomorphism
(5.1) $dd^{c}\psi]_{\omega}\wedge\mu=[dd^{c}(\psi-\psi_{0})]\wedge\mu+\omega\wedge\mu$
on ${\mathcal{QZ}}(Y)$.
Recall that if $\psi\equiv$-$\infty$, then $dd^{c}(\psi-\psi_{0})]\wedge\mu=0$, so that $dd^{c}\psi]_{\omega}=\omega\wedge\mu$.
Proposition 5.2.
(i) If $e^{\psi}$ is a qas metric on $L\to Y$, then $dd^{c}\psi]^{k}_{\omega}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y)$.
(ii) If $e^{\psi^{\prime}}$ is another qas metric and $\psi^{\prime}=\psi+{\mathcal{O}}(1)$ then $dd^{c}\psi^{\prime}]^{k}_{\omega}-[dd^{c}\psi]^{k}_{\omega}\colon{\mathcal{QZ}}(Y)\to{\mathcal{QZ}}_{0}(Y)$.
(iii) If $\tau\colon Y^{\prime}\to Y$ is a proper mapping and $\mu=\tau_{*}\mu^{\prime}$, then
(5.2) $dd^{c}\psi]_{\omega}\wedge\mu=\tau_{*}\big([dd^{c}\tau^{*}\psi]_{\tau^{*}\omega}\wedge\mu^{\prime}\big).$
Proof.
The case $k=1$ of (i) follows directly from (5.1), Proposition 4.16 (i), and the fact that multiplication by $\omega$ maps ${\mathcal{QZ}}_{0}(Y)$ into itself. The general case follows by trivial induction. Part (ii) follows direct from Proposition 4.16 (ii) applied to $u^{\prime}=\psi^{\prime}-\psi_{0}$ and $u=\psi-\psi_{0}$. Part (iii) follows from Proposition 4.10. ∎
It is sometimes useful to express $dd^{c}\psi]_{\omega}$ in terms of $dd^{c}\psi]$ which is well-defined even when $\psi$ is a metric. In fact, (4.9) is local and not affected if $u$ is replaced by $u+h$, where $h$ is pluriharmonic. Let $Z$ be the degeneracy locus of $e^{\psi}$. If $\mu={\bf 1}_{Y\setminus Z}\mu$, then (5.1) is $dd^{c}(\psi-\psi_{0})]\wedge\mu+\omega\wedge\mu=[dd^{c}\psi]\wedge\mu,$ and if $\mu={\bf 1}_{Z}\mu$, then it is just $\omega\wedge\mu$. Therefore,
(5.3) $dd^{c}\psi]_{\omega}\wedge\mu=[dd^{c}\psi]\wedge{\bf 1}_{Y\setminus Z}\mu+\omega\wedge{\bf 1}_{Z}\mu;$
this is precisely the definition of $dd^{c}\psi]_{\omega}$ in [18]. Besides $dd^{c}\psi]^{k}$ we have endomorphisms
$\langle dd^{c}\psi\rangle^{k}\wedge\mu={\bf 1}_{Y\setminus Z}[dd^{c}\psi]^{k}\wedge\mu,\quad M^{\psi}_{k}\wedge\mu={\bf 1}_{Z}[dd^{c}\psi]^{k}\wedge\mu;$
on ${\mathcal{QZ}}(Y)$. It is readily verified by induction that
(5.4) $dd^{c}\psi]^{k}_{\omega}\wedge\mu=\langle dd^{c}\psi\rangle^{k}\wedge\mu+\sum_{\ell=0}^{k}\omega^{k-\ell}\wedge M^{\psi}_{\ell}\wedge\mu,\quad k=0,1,2,\ldots.$
Theorem 5.3.
Let $\chi_{j}$ be a sequence of smooth bounded from below convex increasing functions on $\mathbb{R}$}, which decrease to the identity function as $j\to\infty$. Let $e^{\psi}$ be a qas metric on $L\to Y$, $\psi_{0}$ a smooth metric, and let $\psi_{j}=\chi_{j}\circ(\psi-\psi_{0})+\psi_{0}$. If $\omega=dd^{c}\psi_{0}$, then for each $\mu\in{\mathcal{QZ}}(Y)$,
(5.5) $(dd^{c}\psi_{j})^{k}\wedge\mu\to[dd^{c}\psi]^{k}_{\omega}\wedge\mu,\quad j\to\infty,\quad k=0,1,2,\ldots.$
Since $\psi-\psi_{0}$ is a global function $\psi_{j}$ are well-defined. The smoothness assumption on $\chi_{j}$ is not necessary, but for simplicity we restrict to this case, since we will only apply the theorem to $\chi_{\epsilon}(t)=\log(e^{t}+{\epsilon})$.
This theorem is known when $\mu$ is $1$ and with $dd^{c}\psi]^{k}_{\omega}$ expressed as in (5.4). In [7] it is proved for a more general class of (q)psh functions. Here we will rely on the presentation in [10] and show that the proof given there works even when acting on a quasi-cycle $\mu$.
Proof.
First assume that $\psi\equiv$-$\infty$. Then $\psi_{j}=\chi_{j}(-\infty)+\psi_{0}$ so that $(dd^{c}\psi_{j})^{k}\wedge\mu=\omega^{k}\wedge\mu=[dd^{c}\psi]^{k}_{\omega}\wedge\mu$. Let us thus assume that $\psi\not\equiv$-$\infty$. Since the statement is local, we can restrict to an open subset where we have a global frame element, and hence that $\psi$ and $\psi_{0}$ are fixed functions. For simplicity we let $Y$ denote this open set.
We begin with the case when $\psi_{0}=0$. If $\mu$ has support on $Z$, then both sides of (5.5) are equal to $\omega^{k}\wedge\mu$. Thus we can assume that $\mu=\tau^{*}\gamma$, where $\tau\colon W\to Y$, $\gamma$ is a product as in (4.1) and $\tau^{-1}Z$ has positive codimension in $W$. After possibly another modification we may assume that $\tau^{*}\psi=\log|\sigma|^{2}+a$, where $\sigma$ is a section of a Hermitian line bundle ${\mathcal{L}}\to W$ and $a$ is qpsh and locally bounded. We claim that
(5.6) $(dd^{c}\chi_{j}\circ\tau^{*}\psi)^{k}\wedge\gamma\to[dd^{c}\tau^{*}\psi]^{k}\wedge\gamma.$
Following [10] there are unique convex increasing functions $\gamma_{j}$ such that $\gamma_{j}(-1)=\chi_{j}(-1)$ and $\gamma^{\prime}_{j}=(\chi^{\prime}_{j})^{k}.$ It follows that also $\gamma_{j}(t)$ decreases to $t$. A direct computation yields
$(dd^{c}(\chi_{j}\circ\tau^{*}\psi))^{k}\wedge\gamma=dd^{c}((\gamma_{j}\circ\tau^{*}\psi)\nu^{k-1}\wedge\gamma),$
where $\nu=dd^{c}a-c_{1}({\mathcal{L}})$. Since $\gamma_{j}\circ\tau^{*}\psi\to\tau^{*}\psi$ in $L^{1}$ with respect to the (total variation of the) measure $\nu^{k-1}\wedge\gamma$, we see that by definition,
$(dd^{c}\chi_{j}\circ\tau^{*}\psi)^{k}\wedge\gamma\to dd^{c}(\tau^{*}\psi\nu^{k-1}\wedge\gamma)=([\operatorname{div}\sigma]+\nu)\wedge\nu^{k-1}\wedge\gamma=[dd^{c}\tau^{*}\psi]^{k}\wedge\gamma;$
Here the last equality follows from Example 4.9 and $\nu=dd^{c}a-c_{1}({\mathcal{L}})$. Thus (5.6) holds. Applying $\tau_{*}$ and using Proposition 4.10, now (5.5) follows for $\psi_{0}=0$.
Let us now assume that $\psi_{0}$ is an arbitrary smooth function. Using the case $\psi_{0}=0$,
(5.7) $(dd^{c}\psi_{j})^{k}\wedge\mu=\sum_{\ell=0}^{k}{k\choose\ell}\omega^{k-\ell}\wedge(dd^{c}\chi_{j}\circ(\psi-\psi_{0}))^{\ell}\wedge\mu\to\\ \sum_{\ell=0}^{k}{k\choose\ell}\omega^{k-\ell}\wedge[dd^{c}(\psi-\psi_{0})]^{\ell}\wedge\mu=([dd^{c}(\psi-\psi_{0})]+\omega)^{k}\wedge\mu,$
where the last power is composition of the endomorphism $\mu\mapsto([dd^{c}(\psi-\psi_{0})]+\omega)\wedge\mu$ on ${\mathcal{QZ}}(Y)$. Now (5.5) follows from (5.1) and (5.7). ∎
6. Segre forms and endomorphisms of qas bundles
Assume that $(E,h)$ is a qas bundle over $X$. If $e^{\psi}$ denotes the induced metric on $L_{E}$, then $\psi=\log|\alpha|^{2}_{h}$ is qas. Thus for any local frame $a$ of $L_{E}$ the function $\log|\alpha a|^{2}_{h}$ is qas on ${\mathbb{P}}(E)$. On the set where $\alpha_{j}\neq 0$ we may take $a=1/\alpha_{j}$.
Let $h_{0}$ be a smooth reference metric on $E$, let $e^{\psi_{0}}$ denote the induced metric on $L_{E}$, and $\omega=dd^{c}\psi_{0}$. We thus have that
$s(L,e^{\psi_{0}})=\sum_{\ell=0}^{\infty}\omega^{\ell}=\frac{1}{1-\omega}.$
If $Z^{\prime}\subset{\mathbb{P}}(E)$ denotes the set where $\psi$ is not locally bounded, then outside of $Z^{\prime}$ we have that $s_{1}(L_{E},e^{\psi})=dd^{c}\psi$. Recall, cf. (4.9) and (5.3), that
(6.1) $dd^{c}\psi]_{\omega}\wedge\mu=dd^{c}(\psi{\bf 1}_{X\setminus Z^{\prime}}\mu)+\omega\wedge{\bf 1}_{Z^{\prime}}\mu$
defines an endomorphism on ${\mathcal{QZ}}({\mathbb{P}}(E))$. In view of Lemma 4.12 and Proposition 5.2 (i) the following definition makes sense.
Definition 6.1.
Given a qas bundle $(E,h)$ over $X$, and a smooth metric $h_{0}$, we have the Segre endomorphisms ${\mathcal{QZ}}(X)\to{\mathcal{QZ}}(X)$ and ${\mathcal{QZ}}_{0}(X)\to{\mathcal{QZ}}_{0}(X)$ of bidegree $(k,k)$,
(6.2) ${\bf s}_{k}(E,h,h_{0})\wedge\mu=p_{*}([dd^{c}\psi]_{\omega}^{k+r-1}\wedge p^{*}\mu).$
We let
${\bf s}(E,h,h_{0}):=\sum_{k}{\bf s}_{k}(E,h,h_{0}).$
We define the Segre form $s(E,h,h_{0}):={\bf s}(E,h,h_{0})1$ on $X$, i.e.,
(6.3) $s(E,h,h_{0}):=p_{*}\sum_{k=0}^{\infty}[dd^{c}\psi]^{k}_{\omega}.$
Note that if $L$ is a line bundle, then $p_{L}\colon{\mathbb{P}}(L)\to X$ is the identity. Hence,
(6.4) $s(E,h,h_{0}):=p_{*}s(L_{E},e^{\psi},e^{\psi_{0}}).$
Remark 6.2.
The definition of $s(E,h,h_{0})$ here coincides with the definition in [18]. However, in that article is used the convention that ${\mathbb{P}}(E)_{x}$ denotes the hyperplanes in $E_{x}$ so one has to make some necessary adaptions. Moreover, the definition is made on the associated (Griffiths positive) dual bundle $(E^{*},h^{*})$, see, e.g., [17, Lemma 3.1]. As in the smooth case one have the relations
$s_{k}(E,h,h_{0})=(-1)^{k}s_{k}(E^{*},h^{*},h_{0}^{*}),$
which connect the definitions in this article and in [18]. ∎
By (5.4), with $Y={\mathbb{P}}(E)$ and $Z=Z^{\prime}$, we have
(6.5) $\sum_{k=0}^{\infty}[dd^{c}\psi]^{k}_{\omega}\wedge\mu=\sum_{k=0}^{\infty}\langle dd^{c}\psi\rangle^{k}\wedge\mu+\sum_{\ell=0}^{\infty}\sum_{k=\ell}^{\infty}M^{\psi}_{\ell}\wedge(dd^{c}\psi_{0})^{k-\ell}\wedge\mu=\\ \left(\sum_{k=0}^{\infty}\langle dd^{c}\psi\rangle^{k}+s(L_{E},e^{\psi_{0}})\wedge M^{\psi}\right)\wedge\mu.$
Thus
(6.6) ${\bf s}(E,h,h_{0})\wedge\mu=p_{*}\big(\sum_{\ell=0}^{\infty}\langle dd^{c}\psi\rangle^{\ell}\wedge p^{*}\mu+s(L_{E},e^{\psi_{0}})\wedge M^{\psi}\wedge p^{*}\mu\big).$
Since the support of $M^{\psi}\wedge p^{*}\mu$ is contained in $p^{-1}Z$,
(6.7) ${\bf 1}_{X\setminus Z}{\bf s}(E,h,h_{0})\wedge\mu=p_{*}\big({\bf 1}_{{\mathbb{P}}(E)\setminus p^{-1}Z}\sum_{\ell=0}^{\infty}\langle dd^{c}\psi\rangle^{\ell}\wedge p^{*}\mu\big).$
It is independent of choice of reference metric $h_{0}$ and we denote it by ${\bf s}^{\prime}(E,h)\wedge\mu$. Letting
$M^{E,h,h_{0}}\wedge\mu:={\bf 1}_{Z}{\bf s}_{k}(E,h,h_{0})\wedge\mu$
we have the decomposition
(6.8) ${\bf s}(E,h,h_{0})\wedge\mu={\bf s}_{k}^{\prime}(E,h)\wedge\mu+M^{E,h,h_{0}}\wedge\mu.$
Note that by (4.6), both terms on the right-hand side belong to ${\mathcal{QZ}}(X)$.
Example 6.3.
Let $(E,h_{0})$ and $(F,h_{F})$ be Hermitian vector bundles over $X$, and let $g\colon E\to F$ be a holomorphic morphism. Let $h$ be the singular metric on $E$ defined by $|\alpha|_{h}:=|g\alpha|_{h_{F}}$. Recall the representation $\psi=\log|\alpha|_{h}^{2}/|\alpha_{j}|^{2}$ of the induced metric on $L_{E}$, (2.2). In [1], the corresponding representation is denoted by $\log|\alpha|^{2}_{\circ}$. It follows from (6.6), with $\mu=1$, that the definition given in [1, Example 9.2] of $s(E)$ coincides with (6.3). ∎
6.1. The case when $h$ is qas and non-degenerate
We need the following technical result.
Proposition 6.4.
Let $\psi(x,y)$ be a bounded qpsh function on $U\subset{\mathbb{C}}^{n}_{x}\times{\mathbb{C}}^{m}_{y}$, such that $\psi$ is smooth in $y$ for each fixed $x$, let $v(x,y)$ be a smooth psh function, and let $\rho_{\epsilon}(x)=\rho(x/{\epsilon})/{\epsilon}^{n}$ be a family of radial mollifiers with respect to the variable $x$. Define
(6.9) $\psi_{\epsilon}=\log(e^{\psi}*_{x}\rho_{\epsilon}+A{\epsilon}e^{v}),$
where $\psi*_{x}\rho_{\epsilon}$ denotes the convolution with respect to $x$. Then for any $V\subset\subset U$, $\psi_{\epsilon}$ is smooth on $V$ for small ${\epsilon}>0$, and if $A\gg 1$, then $\psi_{\epsilon}$ decreases to $\psi$ when ${\epsilon}\to 0$. Moreover,
(6.10) $(dd^{c}\psi_{\epsilon})^{k}\wedge\mu\to(dd^{c}\psi)^{k}\wedge\mu,\quad\mu\in{\mathcal{QZ}}(U).$
Since $\psi$ is locally bounded we write $(dd^{c}\psi)^{k}\wedge\mu$ rather than $dd^{c}\psi]^{k}\wedge\mu$.
Proof.
We can choose a constant $B>0$ such that $\psi+g$ is psh on $V$ if $g=B(|x|^{2}+|y|^{2})$. For ${\epsilon}>0$,
$\psi_{\epsilon}=\log\big((e^{\psi}*_{x}\rho_{\epsilon})e^{g}+e^{v+\log({\epsilon}A)+g}\big)-g.$
We claim that the first term $\psi^{\prime}$ on the right is psh. To see this first assume that $\psi$ is smooth. If we approximate the convolution with a Riemann sum we get
(6.11) $\log\big(\sum_{k}e^{\psi(\cdot-x_{k},\cdot)+c_{k}+g}+e^{v+\log({\epsilon}A)+g}\big).$
Since $dd^{c}g$ is translation invariant each $\psi(\cdot-x_{k},\cdot)+g$ is psh. Also $v+g$ is psh so the function in (6.11) is psh. Since such functions tend uniformly to $\psi^{\prime}$ we conclude that $\psi^{\prime}$ is psh. If $\psi$ is not smooth we approximate $u=\psi+g$ by a decreasing sequence $u_{j}$ of smooth psh functions and let $\psi_{j}=u_{j}-g$. By the same argument, each $\psi^{\prime}_{j}$ is psh, and since $\psi_{j}$ decreases to $\psi$, $\psi^{\prime}$ is psh. Thus the claim is proved.
Note that $\psi_{\epsilon}$ is smooth as a function of $x$ and $y$. Clearly $\psi_{\epsilon}\to\psi$ when ${\epsilon}\to 0$. We now prove that $\psi_{\epsilon}$ decreases with $\epsilon$} on $V$ if $A$ is large enough. To see this, we may fix $y$ and consider everything as a function of $x$, as long as the choice of $A$ is uniform in $y$. We write $\psi=u+g$, where $u$ is psh and $g$ is smooth on $X$. Then
$e^{\psi_{\epsilon}}(x)=e^{g(x)}(e^{u}\ast\rho_{\epsilon})(x)+\int(e^{g(x-x^{\prime})}-e^{g(x)})\rho_{\epsilon}(x^{\prime})dV(x^{\prime})+A\epsilon e^{v}.$
Since $e^{u}$ is psh, the first term of the right hand side decreases when $\epsilon$ decreases. To see that the whole expression is decreasing when $\epsilon$ decreases it thus suffices to prove that the middle term is differentiable in $\epsilon$, with a derivative less than $Ae^{v}$. In fact, $(d/d\epsilon)\rho_{\epsilon}(x^{\prime})={\mathcal{O}}(1/\epsilon^{n+1})$, and has support where $|x^{\prime}|\leq\epsilon$. Since $|e^{g(x-x^{\prime})}-e^{g(x)}|\leq C\epsilon$ for $|x^{\prime}|\leq\epsilon$ (with $C$ depending on the $C^{1}$-norm of $e^{g}$), the derivative of the middle term is bounded by a constant that is smaller than $Ae^{v}$ on $V$ if $A$ is chosen large enough. This choice can also be made uniformly in $y$.
It remains to prove (6.10). As before we have that $\psi_{\epsilon}=\psi_{\epsilon}+g-g=u_{\epsilon}+g$, where $u_{\epsilon}$ are psh and decrease to $u$ when ${\epsilon}\to 0$, and $g$ is smooth. Applying the Bedford-Taylor theory of convergence of Monge-Ampère products for bounded psh functions, one obtains (6.10) in case $\mu$ has the form (4.1). If $\mu=\tau_{*}\gamma$, where $\tau\colon W\to U$ is proper, then $\tau^{*}\psi_{\epsilon}+\tau^{*}g$ is psh and smooth on $\tau^{-1}V$ and decreases to $\tau^{*}\psi+\tau^{*}g$ when ${\epsilon}\to 0$. We conclude that $(dd^{c}\tau^{*}\psi_{\epsilon})^{k}\wedge\gamma\to(dd^{c}\tau^{*}\psi)^{k}\wedge\gamma$ and taking $\tau_{*}$ we get (6.10). ∎
We will now see that, as long as qas bundles have empty degeneracy loci, their Segre forms in many ways behave as in the smooth case.
Theorem 6.5.
Assume that $E\to X$ is a holomorphic vector bundle equipped with a non-degenerate qas metric $h$.
(i) The Segre endomorphism on ${\mathcal{QZ}}(X)$ induced by $h$ is independent of the choice of reference metric, and is denoted by ${\bf s}(E,h)$. On the (possibly empty) open subset where $h$ is smooth, ${\bf s}(E,h)$ coincides with multiplication by the Segre form of $h$ there.
(ii) Locally there exist smooth metrics $h_{\epsilon}$, decreasing to $h$ when $\epsilon\to 0$, such that
(6.12) $s(E,h_{\epsilon})\wedge\mu\to{\bf s}(E,h)\wedge\mu,\quad{\epsilon}\to 0,\quad\mu\in{\mathcal{QZ}}(X).$
(iii) If $(E^{\prime},h^{\prime})$ is another qas bundle and $h^{\prime}$ is non-degenerate as well, then
(6.13) ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h)\wedge\mu={\bf s}(E,h)\wedge{\bf s}(E^{\prime},h^{\prime})\wedge\mu,\quad\mu\in{\mathcal{QZ}}(X).$
(iv) For any $\ell$ and $k$ the endomorphisms ${\bf s}_{k}(E,h)$ and ${\bf s}_{\ell}(E^{\prime},h^{\prime})$ commute.
Proof.
The hypothesis means that $Z=\emptyset$ and thus $s(E,h,h_{0})=s^{\prime}(E,h)$. Therefore it is independent of $h_{0}$, cf. (6.7). It follows by (2.3), (4.17) and (6.7) that if $h$ is smooth on an open set, then ${\bf s}(E,h)$ coincides with multiplication by the Segre form of $h$ there.
We now prove (ii). Fix a trivialization of $E$, and a local coordinate system $x$ on $X$ such that locally, $E\cong{\mathbb{C}}^{r}_{\alpha}\times{\mathbb{C}}^{n}_{x}$. Let $\rho_{\epsilon}$ be a family of radial mollifiers, and let
(6.14) $h_{\epsilon}=h*\rho_{\epsilon}+A{\epsilon}I,$
where $I$ is the identity matrix and $A$ is a large constant. Note that $h_{\epsilon}$ depends on both the trivialization of $E$ and the local coordinate system on $X$. It is readily verified that a convolution of positive definite matrices is positive definite, and the entries are smooth, so that each $h_{\epsilon}$ indeed defines a smooth metric on $E$.
Fix a point $x_{0}\in X$ and a small neighborhood $V\subset X$. Consider the subset of $p^{-1}V\subset{\mathbb{P}}(E)$ where $\alpha_{j}\neq 0$. With respect to the frame $\alpha/\alpha_{j}$ of $L_{E}$, the metric $e^{\psi_{\epsilon}}$ induced by $h_{\epsilon}$ is represented by the function $\psi_{\epsilon}=\log|\alpha/\alpha_{j}|_{h_{\epsilon}(x)}$. Then $\psi_{\epsilon}$ is of the form (6.9) with $\psi=\log|\alpha/\alpha_{j}|^{2}_{h(x)}$, and $v=\log(|\alpha/\alpha_{j}|^{2})$. By Proposition 6.4, if $A$ is chosen large enough, $\psi_{\epsilon}$ is decreasing when ${\epsilon}\to 0$. By compactness the same $A$ works where $\alpha_{j}\neq 0$ for any $j$. We conclude that $h_{\epsilon}$ decreases to $h$. It follows from Proposition 6.4 that
(6.15) $s(L_{E},e^{\psi_{\epsilon}})\wedge p^{*}\mu=\sum_{k}(dd^{c}\psi_{\epsilon})^{k+r-1}\wedge p^{*}\mu\to\sum_{k}(dd^{c}\psi)^{k+r-1}\wedge p^{*}\mu.$
Applying $p_{*}$ we get (6.12), keeping in mind, cf. (4.17), that ${\bf s}(E,h_{\epsilon})\wedge\mu=s(E,h_{\epsilon})\wedge\mu$ since $h_{\epsilon}$ are smooth. Thus (ii) is proved.
We now consider (iii). Let $\psi^{\prime}=\log|\alpha^{\prime}|^{2}_{h^{\prime}(x)}$. We can assume that $g$ is such that also $\psi^{\prime}+g$ is psh. Let $h^{\prime}_{{\epsilon}^{\prime}}$ be a similar regularization. It then follows that $\log(e^{\psi^{\prime}_{{\epsilon}^{\prime}}}+e^{\psi_{\epsilon}})+g$ is psh for each $\epsilon$$^{\prime}$ and $\epsilon$} and decreases to $\log(|\alpha^{\prime}|^{2}_{h^{\prime}(x)}+|\alpha|^{2}_{h(x)})+g$ when $({\epsilon}^{\prime},{\epsilon})$ decreases to $(0,0)$. With the same argument as in the proof of (ii) we conclude that
(6.16) $s(E^{\prime}\oplus E,h^{\prime}_{{\epsilon}^{\prime}}\oplus h_{\epsilon})\wedge\mu\to{\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h)\wedge\mu$
unconditionally when ${\epsilon}^{\prime},{\epsilon}\to 0$. With the product formula for smooth metrics
$s(E^{\prime}\oplus E,h^{\prime}_{{\epsilon}^{\prime}}\oplus h_{\epsilon})\wedge\mu=s(E,h_{\epsilon})\wedge s(E^{\prime},h^{\prime}_{{\epsilon}^{\prime}})\wedge\mu.$
If we first let ${\epsilon}^{\prime}\to 0$ for fixed $\epsilon$} and then let ${\epsilon}\to 0$ we get (6.13) from (6.16) and (6.12).
Part (iv) requires an additional argument; it is indeed an immediate consequence of Theorem 6.7 (vii) below. ∎
Let $E\oplus\cdots\oplus E$ be the direct sum of $k$ copies of $E$. By repeated use of (iii) we have ${\bf s}(E\oplus\cdots\oplus E,h\oplus\cdots\oplus h)={\bf s}(E,h)^{k}$ and clearly ${\bf s}(E\oplus\cdots\oplus E,h_{\epsilon}\oplus\cdots\oplus h_{\epsilon})={\bf s}(E,h_{\epsilon})^{k}$. Applying (ii) to the direct sum, noting that $(h\oplus\cdots\oplus h)_{\epsilon}=h_{\epsilon}\oplus\cdots\oplus h_{\epsilon}$, we conclude that
(6.17) ${\bf s}(E,h_{\epsilon})^{k}\to{\bf s}(E,h)^{k}.$
Since $s(E,h)={\bf s}^{\prime}(E,h)1$ it follows, cf. (6.7) that
(6.18) $s(E,h)=p_{*}\sum_{\ell=0}^{\infty}(dd^{c}\psi)^{\ell},$
which is precisely the definition (2.3) when $h$ is smooth but with the Monge-Ampère products interpreted in the Bedford-Taylor sense. Letting $\mu=1$, (ii) implies that
(6.19) $s(E,h_{\epsilon})\to s(E,h).$
Part (iv) means that $s(E\oplus E^{\prime},h\oplus h^{\prime})={\bf s}(E^{\prime},h^{\prime})\wedge s(E,h).$
Remark $6.$
If $h$ is Griffiths negative, then one can take $A=0$ in (6.14). The resulting approximation, in particular (6.19), appeared in [17]. ∎
6.2. The case with a general qas metric $h$
Here is our main theorem on the Segre endomorphism ${\bf s}(E,h,h_{0})$.
Theorem 6.7.
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$ and a smooth reference metric $h_{0}$, let $Z$ be the degeneracy locus of $h$, and let ${\bf s}(E,h,h_{0})$ be the associated Segre endomorphism on ${\mathcal{QZ}}(X)$.
(i) If $h_{\epsilon}$ is the metric defined by (1.1), then
(6.20) ${\bf s}(E,h_{\epsilon})\wedge\mu\to{\bf s}(E,h,h_{0})\wedge\mu,\quad{\epsilon}\to 0,\quad\mu\in{\mathcal{QZ}}(X).$
(ii) The endomorphisms ${\bf s}_{k}(E,h,h_{0})$ vanish for $k<0$ and ${\bf s}_{0}(E,h,h_{0})=I$, the identity mapping.
(iii) If $\tau\colon X^{\prime}\to X$ is a proper mapping, then $\tau^{*}h$ is qas and
(6.21) $\tau_{*}({\bf s}(\tau^{*}E,\tau^{*}h,\tau^{*}h_{0})\wedge\mu)={\bf s}(E,h,h_{0})\wedge\tau_{*}\mu,\quad\mu\in{\mathcal{QZ}}(X^{\prime}).$
(iv) For any $\mu\in{\mathcal{QZ}}(X)$ there is a current $w$ on $X$ such that
(6.22) $({\bf s}(E,h,h_{0})-{\bf s}(E,h_{0}))\wedge\mu=dd^{c}w.$
(v) If $E^{\prime}$ is another holomorphic vector bundle, equipped with a smooth metric $h^{\prime}$, then
(6.23) ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})\wedge\mu={\bf s}(E,h,h_{0})\wedge{\bf s}(E^{\prime},h^{\prime})\wedge\mu,\quad\mu\in{\mathcal{QZ}}(X).$
(vi) For each $k$, ${\bf s}_{k}(E,h,h_{0})$ yields a well-defined endomorphism on ${\mathcal{QB}}(X)$, independent of $h_{0}$, and if $h^{\prime}$ is a qas metric on $E$ comparable to $h$, then ${\bf s}_{k}(E,h^{\prime},h_{0})$ yields the same endomorphism as ${\bf s}_{k}(E,h,h_{0})$ on ${\mathcal{QB}}(X)$.
(vii) Assume that $(E_{1},h_{1})$ and $(E_{2},h_{2})$ are qas bundles, with degeneracy loci $Z_{1}$ and $Z_{2}$, respectively, and $h_{10}$ and $h_{20}$ be smooth reference metrics on these bundles. Let $X^{\prime}\subset\subset X$ be an open relatively compact subset. For any $k,\ell$ and $\mu\in{\mathcal{GZ}}(Y^{\prime})$, there is a current $w$ on $X^{\prime}$, which is the pushforward of a current on $Z_{1}\cap Z_{2}\cap X^{\prime}$, such that
$({\bf s}_{k}(E_{1},h_{1},h_{10})\wedge{\bf s}_{\ell}(E_{2},h_{2},h_{20})-{\bf s}_{\ell}(E_{2},h_{2},h_{20})\wedge{\bf s}_{k}(E_{1},h_{1},h_{10}))\wedge\mu=dd^{c}w.$
Remark 6.8.
(i) If $h_{0}$ and $h_{0}^{\prime}$ are two smooth reference metrics, then by the first part of (vi),
(6.24) $\big({\bf s}_{k}(E,h,h_{0})-{\bf s}_{k}(E,h,h_{0}^{\prime}))\wedge\mu,\quad\mu\in{\mathcal{QZ}}(X),$
belongs to ${\mathcal{QZ}}_{0}(X)$. Since ${\bf s}(E,h,h_{0})$ and ${\bf s}(E,h,h_{0}^{\prime})$ coincide outside of $Z$, it follows from Proposition 4.15 that if $\mu$ has bidegree $(q,q)$, then (6.24) vanishes if $k+q-1<{\rm codim\,}Z$.
(ii) If $\mu$ has bidegree $(q,q)$, then the current $w$ in (vii) has bidegree $(k+\ell+q-1,k+\ell+q-1)$. Since $w$ is the pushforward of a current on $Z$, it vanishes if $k+\ell+q-1<{\rm codim\,}Z$. ∎
Proof.
Notice that if $\chi_{\epsilon}(t)=\log(e^{t}+{\epsilon})$, then
$\psi_{\epsilon}:=\log(|\alpha|_{h}^{2}+{\epsilon}|\alpha|^{2}_{h_{0}})=\log(|\alpha|_{h}^{2}/|\alpha|^{2}_{h_{0}}+{\epsilon})+\log|\alpha|^{2}_{h_{0}}=\log(e^{\psi-\psi_{0}}+{\epsilon})+\psi_{0}=\chi_{\epsilon}\circ(\psi-\psi_{0})+\psi_{0},$
so it follows from Theorem 5.3 that
$(dd^{c}\psi_{\epsilon})^{k}\wedge p^{*}\mu\to[dd^{c}\psi]^{k}_{\omega}\wedge p^{*}\mu.$
Applying $p_{*}$ we get (6.20) and hence (i) is proved.
Now (ii) follows since it holds when the metric is smooth, and hence also when it is qas and non-degenerate by Theorem 6.5 (ii).
Notice that a holomorphic map $\tau\colon X^{\prime}\to X$ induces a holomorphic map $\tau\colon{\mathbb{P}}(\tau^{*}E)\to{\mathbb{P}}(E)$. Part (iii) now follows from Definition 6.1 and Proposition 5.2 (iii).
In (iv) we can assume that $\mu=\tau_{*}\gamma$ where $\gamma$ is a product as in (4.1) on $W$ and $\tau\colon W\to X$. In view of (iii) and by changing notation we have to prove (iv) on $X$ but with $\mu=\gamma$.
If the metric on $E$ is identically 0, i.e., $Z^{\prime}={\mathbb{P}}(E)$, then ${\bf s}(E,h,h_{0})={\bf s}(E,h_{0})$ and then the statement is trivial. Let us thus assume that $Z^{\prime}$ is not equal to ${\mathbb{P}}(E)$. It may very well happen anyhow that $Z=X$. Notice that $p^{*}\gamma$ is a product as in (4.1) on ${\mathbb{P}}(E)$. Let $\psi$ and $\psi_{0}$ be as above. Then $\psi-\psi_{0}$ is a global function on ${\mathbb{P}}(E)$ and, cf. Proposition 4.8,
(6.25) $w^{\prime}=\frac{\psi-\psi_{0}}{(1-dd^{c}\psi_{0})(1-\langle dd^{c}\psi\rangle)}\wedge p^{*}\gamma$
is a locally finite measure. In view of (4.12) and (4.13) we have
(6.26) ${\bf 1}_{{\mathbb{P}}(E)\setminus Z^{\prime}}dd^{c}w^{\prime}=\left(\frac{1}{1-\langle dd^{c}\psi\rangle}-\frac{1}{1-dd^{c}\psi_{0}}\right)\wedge p^{*}\gamma$
and
(6.27) ${\bf 1}_{Z^{\prime}}dd^{c}w^{\prime}=\frac{1}{1-dd^{c}\psi_{0}}{\bf 1}_{Z^{\prime}}dd^{c}\psi\frac{1}{1-dd^{c}\psi}\wedge p^{*}\gamma=s(L_{E},e^{\psi_{0}})\wedge\sum_{\ell=1}^{\infty}M^{\psi}_{\ell}\wedge p^{*}\gamma.$
Notice that $Z^{\prime}$ has positive codimension so that $M^{\psi}_{0}\wedge p^{*}\gamma={\bf 1}_{Z^{\prime}}p^{*}\gamma=0$. Letting $w=p_{*}w^{\prime}$ we conclude from (6.6), (6.26) and (6.27) that $dd^{c}w=({\bf s}(E,h,h_{0})-{\bf s}(E,h_{0}))\wedge\mu.$ Thus (iv) is proved.
We now consider (v). The metrics $h^{\prime}$ and $h_{\epsilon}=h+{\epsilon}h_{0}$ are both qas and non-degenerate. By Theorem 6.5 (iii) therefore
(6.28) ${\bf s}(E,h+{\epsilon}h_{0})\wedge{\bf s}(E^{\prime},(1+{\epsilon})h^{\prime})\wedge\mu={\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h+{\epsilon}(h^{\prime}\oplus h_{0}))\wedge\mu.$
Since $s(E^{\prime},(1+{\epsilon})h^{\prime})=s(E^{\prime},h^{\prime})$, (i) now implies that the left hand side of (6.28) tends to ${\bf s}(E,h)\wedge{\bf s}(E^{\prime},h^{\prime})\wedge\mu$ and the right hand side tends to ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h)\wedge\mu$.
We now consider (vi). Let $h_{0}^{\prime}$ denote another smooth reference metric. By (6.6) we see that ${\bf s}(E,h,h_{0}^{\prime})-{\bf s}(E,h,h_{0})$ is the image under $p_{*}$ of $(s(L_{E},e^{\psi_{0}^{\prime}})-s(L_{E},e^{\psi_{0}}))\wedge M^{\psi}$, which belongs to ${\mathcal{QZ}}_{0}({\mathbb{P}}(E))$. Now the first statement follows. The second one follows from Proposition 5.2 (ii) and (6.2). Thus (vi) is proved. The proof of (vii) is postponed to Section 6.3. ∎
Proof of Theorem 1.3.
Items (i) to (v) are essentially covered by Theorem 6.7 with $\mu=1$. In the decomposition (1.4) in (iii), one must have that $M^{E,h,h_{0}}={\bf 1}_{Z}s(E,h,h_{0})$. By (4.6), $M^{h,h,h_{0}}$ belongs to ${\mathcal{QZ}}(X)$, and in particular, it is closed. By Theorem 6.5 (i), $s^{\prime}(E,h)={\bf 1}_{X\setminus Z}s(E,h,h_{0})$ is independent of $h_{0}$, and it coincides with the Segre form of $h$ on $X\setminus Z$ if $h$ is smooth there.
The decomposition (1.7) holds since $M^{E,h,h_{0}}$ is in ${\mathcal{QZ}}(X)$. The statement about the codimension of the non-vanishing multiplicities of $s_{k}^{\prime}$ and $N_{k}$ follows from Theorem 4.20. From Theorem 6.7 (vi) we know that $s(E,h,h_{0})$ and $s(E,h^{\prime},h_{0}^{\prime})$ in (vii) are in the same class in ${\mathcal{QB}}(X)$. Hence $M^{E,h,h_{0}}$ and $M^{E,h^{\prime},h_{0}^{\prime}}$ are as well by (4.21). By Proposition 4.19 they therefore have the same multiplicities.
We now prove that these numbers are non-negative. In a small open neighborhood $\mathcal{U}$} of a given point on $X$, $E={\mathcal{U}}\times{\mathbb{C}}^{r}_{\alpha}$, and therefore $p^{-1}{\mathcal{U}}={\mathcal{U}}\times{\mathbb{P}}({\mathbb{C}}^{r}_{\alpha})$. If we choose a trivial metric $h_{0}$ on $E$, then $\omega_{\alpha}=s_{1}(L_{E},h_{0})$ is positive. Assume that the metric is inherited from a morphism $g\colon E\to F$ and $F$ has a trivial metric. Then $\psi=\log|g(x)\alpha|^{2}$ is psh. In that case, $s(L_{E},e^{\psi},e^{\psi_{0}})$ is a positive current, and hence, so is $s(E,h,h_{0})=p_{*}(s(L_{E},e^{\psi},e^{\psi_{0}}))$. Furthermore, each of the three terms in the decomposition $s(E,h,h_{0})=s^{\prime}(E,h)+S^{E,h}+N^{E,h,h_{0}}$ is positive as well, and in particular, all their components will have non-negative Lelong numbers. By Lemma 3.7 any qas metric $h$ on $E$ is locally equivalent to a psh metric as above. Thus (vi) is proved. ∎
Remark 6.9.
The proof of the positivity of the multiplicities refers to Lemma 3.7 which in turn requires that $E$ is qas and not only that $L_{E}$ is qas, cf. Remark 3.9. The remainder of the theorem holds assuming only that $L_{E}$ is qas. ∎
Remark 6.10.
Almost all of Theorem 1.3 and Theorem 6.7 have natural extensions to $\mathbb{Q}$}-qas or $\mathbb{R}$}-qas metrics. Indeed, if $h$ is $\mathbb{Q}$}-qas or $\mathbb{R}$}-qas, it follows readily that ${\bf s}(E,h,h_{0})$ yields an endomorphism on ${\mathcal{QZ}}_{\mathbb{Q}}(X)$ or ${\mathcal{QZ}}_{\mathbb{R}}(X)$, cf. Remark 4.11. In that case, the multiplicities referred to in the Theorem 1.3 would be non-negative rational or real numbers, cf. Remark 4.21. ∎
6.3. Composition of Segre endomorphisms
Example 6.11.
With the notation in the proof of Lemma 4.12, and $\mu=\tau_{*}\gamma$, we have
(6.29) ${\bf s}_{k}(E,h,h_{0})\wedge\mu=\Pi_{*}(([dd^{c}(\tau^{\prime})^{*}\psi]_{(\tau^{\prime})^{*}\omega}^{k+r-1}\wedge(p^{\prime})^{*}\gamma),$
where the product on the right takes place on ${\mathbb{P}}(\tau^{*}E)$ and $\Pi=p\circ\tau^{\prime}=\tau\circ p^{\prime}$. To see (6.29) just notice that
$p_{*}([dd^{c}\psi]_{\omega}^{k+r-1}\wedge\tau^{\prime}_{*}(p^{\prime})^{*}\gamma)=p_{*}\tau^{\prime}_{*}([dd^{c}(\tau^{\prime})^{*}\psi]_{(\tau^{\prime})^{*}\omega}^{k+r-1}\wedge(p^{\prime})^{*}\gamma),$
and this holds because $\Gamma=(p^{\prime})^{*}\gamma$ is a product as in (4.1), and from (5.2) therefore
$dd^{c}\psi]^{k+r-1}_{\omega}\wedge\tau^{\prime}_{*}\Gamma=\tau^{\prime}_{*}\big([dd^{c}(\tau^{\prime})^{*}\psi]^{k+r-1}_{(\tau^{\prime})^{*}\omega}\wedge\Gamma\big).$
Notice that ${\mathbb{P}}(\tau^{*}E)={\mathbb{P}}(E)\times_{X}W$. ∎
Let $(E_{1},h_{1})$ and $(E_{2},h_{2})$ be two qas bundles over $X$. Consider (4.16) with $E=E_{2}$, $W={\mathbb{P}}(E_{1})$ and $\tau=p_{1}$. We then get the fiber diagram
(6.30) $\begin{array}[c]{cccc}Y&\stackrel{{\scriptstyle\pi_{2}}}{{\longrightarrow}}&{\mathbb{P}}(E_{2})\\ \downarrow\scriptstyle{\pi_{1}}&&\downarrow\scriptstyle{p_{2}}\\ {\mathbb{P}}(E_{1})&\stackrel{{\scriptstyle p_{1}}}{{\longrightarrow}}&X\end{array}$
with
(6.31) $Y={\mathbb{P}}(E_{2})\times_{X}{\mathbb{P}}(E_{1}),\quad\Pi=p_{1}\circ\pi_{1}=p_{2}\circ\pi_{2}\colon Y\to X.$
Let $h_{10}$ and $h_{20}$ be smooth reference metrics on $E_{1}$ and $E_{2}$ of rank $r_{1}$ and $r_{2}$, respectively. In view of Example 6.11 and (6.2) we have
${\bf s}_{\ell_{2}}(E_{2},h_{2},h_{20})\wedge s_{\ell_{1}}(E_{1},h_{1},h_{10})=(p_{2})_{*}([dd^{c}\psi_{2}]^{\ell_{2}+r_{2}-1}_{\omega_{2}}\wedge p_{1}^{*}s_{\ell_{1}}(E,h_{1},h_{10}))=\\ (p_{2})_{*}([dd^{c}\psi_{2}]^{\ell_{2}+r_{2}-1}_{\omega_{2}}\wedge(\pi_{2})_{*}[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r_{1}-1}_{\pi_{1}^{*}\omega_{1}})=\Pi_{*}\big([dd^{c}\pi_{2}^{*}\psi_{2}]^{\ell_{2}+r_{2}-1}_{\pi_{2}^{*}\omega_{2}}\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r_{1}-1}_{\pi_{1}^{*}\omega_{1}}\big).$
where $e^{\psi_{j}}$ and $e^{\psi_{j0}}$ denote the metrics on $L_{E_{j}}$ induced by $h_{j}$ and $h_{j0}$, and $\omega_{j}=s_{1}(L_{E_{j}},e^{\psi_{j0}})$. In the same way
(6.32) ${\bf s}_{\ell_{2}}(E_{2},h_{2},h_{20})\wedge{\bf s}_{\ell_{1}}(E_{1},h_{1},h_{10})\wedge\mu=\Pi_{*}\big([dd^{c}\pi_{2}^{*}\psi_{2}]^{\ell_{2}+r_{2}-1}_{\pi_{2}^{*}\omega_{2}}\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r_{1}-1}_{\pi_{1}^{*}\omega_{1}}\wedge\Pi^{*}\mu\big),$
Example 6.12.
Let $(E_{1},h_{1}),\ldots,(E_{k},h_{k})$ be disjoint copies of $(E,h)$. Then, similarly,
(6.33) ${\bf s}_{\ell_{k}}(E,h,h_{0})\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h,h_{0})\wedge\mu=\Pi_{*}([dd^{c}\pi_{k}^{*}\psi_{k}]_{\pi_{k}^{*}\omega_{k}}^{\ell_{k}+r-1}\wedge\cdots\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]_{\pi_{1}^{*}\omega_{1}}^{\ell_{1}+r-1}\wedge\Pi^{*}\mu\big).$
In particular, our definition of product of Segre forms coincides with the product defined in [18, Section 6.1], cf. also Remark 6.2. ∎
Lemma 6.13.
Assume that $L_{1},L_{2}$ are qas line bundles over a manifold $Y$, and let $Z=Z_{L_{1}}\cap Z_{L_{2}}$ denote the intersection of their degeneracy loci. Let $Y^{\prime}\subset\subset Y$. For any $\mu\in{\mathcal{QZ}}(Y^{\prime})$ there exists a current $w$ on $Y^{\prime}$, that is the push-forward of a current on $Z\cap Y^{\prime}$, such that
$({\bf s}_{1}(L_{1})\wedge{\bf s}_{1}(L_{2})-{\bf s}_{1}(L_{2})\wedge{\bf s}_{1}(L_{1}))\wedge\mu=dd^{c}w.$
For simplicity we write ${\bf s}(L_{j})$ rather than ${\bf s}(L_{j},e^{\psi_{j}},e^{\psi_{j0}})$.
Proof.
By taking $\mu=\tau_{*}\gamma$, $\tau\colon W\to Y$, where $\gamma$ is of the form (4.1), we may assume by (6.21), after replacing $Y$ by $W$, that $\mu$ is of the form (4.1). Let $\psi_{j}$ denote the metric on $L_{j}$, for $j=1,2$, and let $\omega_{j}$ denote the first Segre form of the smooth reference metric on $L_{j}$. Either some of $\psi_{j}\equiv$-$\infty$, say it is the case for $\psi_{1}$. Then ${\bf s}_{1}(L_{1})$ acts by multiplication with the smooth form $\omega_{1}$, which commutes with ${\bf s}_{1}(L_{2})$, and we may take $w_{0}=0$. Hence we can assume that $\psi_{j}\not\equiv$-$\infty$ for both $j$.
In view of Proposition 5.2 (iii) we can replace $Y$ by any modification of it. As in the beginning of Section 4.1 we may therefore assume that
$dd^{c}\psi_{j}=[D_{j}]+[W_{j}]+\nu_{j},$
where $D_{j},W_{j}$ are divisors for $j=1,2$, $|D_{1}|=|D_{2}|$, and all irreducible components of $W_{1},W_{2},D_{1}$ intersect properly. Notice that ${\bf s}_{1}(L_{1})\wedge[D_{2}]\wedge\mu=\omega_{1}\wedge[D_{2}]\wedge\mu$ whereas
${\bf s}_{1}(L_{1})\wedge([W_{2}]+\nu_{2})\wedge\mu=([D_{1}]+[W_{1}]+\nu_{1})\wedge([W_{2}]+\nu_{2})\wedge\mu,$
where the right hand side can be computed formally since the intersections that occur are proper. We therefore get
(6.34) $({\bf s}_{1}(L_{1})\wedge{\bf s}_{1}(L_{2})-{\bf s}_{1}(L_{2})\wedge{\bf s}_{1}(L_{1}))\wedge\mu=([D_{1}]\wedge([W_{2}]+\nu_{2}-\omega_{2})-[D_{2}]\wedge([W_{1}]+\nu_{1}-\omega_{1}))\wedge\mu.$
If we choose a smooth metric on ${\mathcal{O}}(D_{j})$ we get a smooth representative $\alpha_{j}$ of its Bott-Chern class. Since $W_{j}]+\nu_{j}-\omega_{j}$ defines the same class there is a global function $u_{j}$ such that
$dd^{c}u_{j}=[W_{j}]+\nu_{j}-\omega_{j}-\alpha_{j}.$
Locally $u_{j}$ is a potential to $\nu_{j}$ plus a potential to $W_{j}]$ plus a smooth function, and hence it is qas. It now follows, cf. the proof of Lemma 4.7, that
$dd^{c}(u_{1}[D_{2}]\wedge\mu)=[dd^{c}u_{1}]\wedge[D_{2}]\wedge\mu=([W_{1}]+\nu_{1}-\omega_{1}-\alpha_{1})\wedge[D_{2}]\wedge\mu.$
Thus $w=(u_{1}[D_{2}]-u_{2}[D_{1}])\wedge\mu$ is a form with support on $Z=|D_{1}|=|D_{2}|$ such that, $dd^{c}w$ is equal to (6.34) minus
(6.35) $(\alpha_{1}[D_{2}]-\alpha_{2}[D_{1}])\wedge\mu.$
It remains to find a $dd^{c}$-potential to (6.35) in $Y^{\prime}$ that is the push-forward of a current on $Z\cap Y^{\prime}$. Since $Y^{\prime}\subset\subset Y$, $Z\cap Y^{\prime}$ has a finite number of irreducible components $D^{\ell}$. There are integers $a_{j}^{\ell}$ such that $D_{j}=\sum_{\ell}a_{j}^{\ell}D^{\ell},\ j=1,2.$ If $\omega^{\ell}$ are smooth representatives of the Bott-Chern class of $D^{\ell}]$, then we can choose $\alpha_{j}=\sum_{\ell}a_{j}^{\ell}\omega^{\ell}$ on $Y^{\prime}$. Thus
$(\alpha_{1}[D_{2}]-\alpha_{2}[D_{1}])\wedge\mu=\sum_{\ell\neq k}a_{1}^{\ell}a_{2}^{k}(\omega^{\ell}[D^{k}]-\omega^{k}[D^{\ell}])\wedge\mu.$
If $v_{\ell}$ are functions such that $dd^{c}v_{\ell}=[D^{\ell}]-\omega^{\ell}$, then as before, suppressing $\wedge\mu$ in the notation,
$dd^{c}(v_{\ell}[D^{k}]-v_{k}[D^{\ell}])=\omega^{\ell}[D^{k}]-\omega^{k}[D^{\ell}]-([D^{\ell}]\wedge[D^{k}]-[D^{k}]\wedge[D^{\ell}])=\omega^{\ell}[D^{k}]-\omega^{k}[D^{\ell}]$
since the $D^{\ell}]\wedge[D^{k}]$ is a proper intersection and thus commuting. Thus $w=\sum a_{1}^{\ell}a_{2}^{k}(v_{\ell}[D^{k}]-v_{k}[D^{\ell}])$ is a $dd^{c}$-potential to (6.35) as desired. ∎
Proof of Theorem 6.7 (vii).
Let $Y={\mathbb{P}}(E_{2})\times_{X}{\mathbb{P}}(E_{1})$ be as above, let $L_{j}\to Y$ be the pullback to $Y$ of the tautological line bundle $L_{E_{j}}\to{\mathbb{P}}(E_{j})$ and let $Y^{\prime}$ be the analogous set but over $X^{\prime}\subset\subset X$. Then $Y^{\prime}\subset\subset Y$. For simplicity we write $Y$ rather than $Y^{\prime}$ in what follows. By (6.32) it is enough to prove the similar statement for $L_{j}\to Y$, cf. (6.4), and $\ell,k$ replaced by $k+r_{2}-1,\ell+r_{1}-1$. Since $L_{j}$ are line bundles, ${\bf s}_{k}(L_{j})={\bf s}_{1}^{k}(L_{j})$, and so
$\big({\bf s}_{k}(L_{1})\wedge{\bf s}_{\ell}(L_{2})-{\bf s}_{\ell}(L_{2})\wedge{\bf s}_{k}(L_{1})\big)\wedge\mu=({\bf s}_{1}(L_{1})^{k}\wedge{\bf s}_{1}(L_{2})^{\ell}-{\bf s}_{1}(L_{2})^{\ell}\wedge{\bf s}_{1}(L_{1})^{k})\wedge\mu.$
We may rewrite the right-hand side as a sum of terms of the form
$\mathbf{S}^{+}\wedge\big({\bf s}_{1}(L_{1})\wedge{\bf s}_{1}(L_{2})-{\bf s}_{1}(L_{2})\wedge{\bf s}_{1}(L_{1})\big)\wedge\mathbf{S}^{-}\wedge\mu,$
where $\mathbf{S}^{+}$ and $\mathbf{S}^{-}$ are products of ${\bf s}_{1}(L_{1})$ and ${\bf s}_{1}(L_{2})$. Let $Z=Z_{L_{1}}\cap Z_{L_{2}}$. It suffices to prove that each such term is of the form $dd^{c}w$, where $w$ is a current that is the push-forward of a current on $Z$. Since $\mathbf{S}^{-}\wedge\mu$ is in $\mathcal{QZ}$}, we obtain by Lemma 6.13 a current $w_{0}$, which is the push-forward of a current on $Z$, such that
$\big({\bf s}_{1}(L_{1})\wedge{\bf s}_{1}(L_{2})-{\bf s}_{1}(L_{2})\wedge{\bf s}_{1}(L_{1})\big)\wedge\mathbf{S}^{-}\wedge\mu=dd^{c}w_{0},$
where $w_{0}$ is current that is the push-forward of a current on $Z$. Note that $w_{0}$ has support on $Z$, and both ${\bf s}_{1}(L_{1})$ and ${\bf s}_{1}(L_{2})$ act by multiplication with a closed smooth form on currents in $\mathcal{QZ}$} with support on $Z$, and hence $\mathbf{S}^{+}$ acts as multiplication with a closed smooth form $S^{+}_{0}$, so we may take $w=S^{+}_{0}\wedge w_{0}$. ∎
7. Chern forms of qas bundles
Let $(E,h)$ be a qas bundle, $h_{0}$ a smooth reference metric, and let $\hat{\bf s}(E,h,h_{0})={\bf s}(E,h,h_{0})-I$. In this section, we will sometimes drop $h$ and $h_{0}$ in the notation. In view of Theorem 6.7 (ii), $\hat{\bf s}(E):={\bf s}(E)-I={\bf s}_{1}(E)+{\bf s}_{2}(E)+\cdots$ is an endomorphism on ${\mathcal{QZ}}(X)$ of positive bidegree. Let $\hat{\bf s}(E)^{k}$ the composition of it $k$ times. We define the Chern endomorphism
(7.1) ${\bf c}(E,h,h_{0}):=\sum_{k=0}^{\infty}(-1)^{k}\hat{\bf s}(E,h,h_{0})^{k}=I-\hat{\bf s}(E,h,h_{0})+\hat{\bf s}(E,h,h_{0})^{2}+\cdots.$
Proposition 7.1.
Let $E\to X$ be an qas bundle. Then
(7.2) ${\bf c}(E,h,h_{0})\wedge{\bf s}(E,h,h_{0})={\bf s}(E,h,h_{0})\wedge{\bf c}(E,h,h_{0})=I.$
The statements in Theorem 6.5 and (ii)-(vi) of Theorem 6.7 hold if ${\bf s}(E,h,h_{0})$ is replaced by ${\bf c}(E,h,h_{0})$.
Proof.
Clearly ${\bf s}(E,h,h_{0})$ and $\hat{\bf s}(E,h,h_{0})$ commute, and hence the computation
${\bf s}(E,h,h_{0})\wedge{\bf c}(E,h,h_{0})=(I+\hat{\bf s}(E,h,h_{0}))\sum_{k=0}^{\infty}(-1)^{k}\hat{\bf s}(E,h,h_{0})^{k}=I,$
is legitimate. The other equality in (7.2) follows in the same way.
We now consider the analogue for $\bf c$} of Theorem 6.5, so assume that $h$ is qas and non-denegerate. Then clearly $c(E,h)$ is independent of $h_{0}$. If $h_{\epsilon}$ is a local regularization, then it follows from (6.17) that the analogue for ${\bf c}(E,h)$ of part (ii) holds. Part (iii) implies its analogue for $\bf c$}: Suppressing $h$ and $h^{\prime}$ in the notation, by (i) and (iii) for $\bf s$} we have $I={\bf s}(E^{\prime}\oplus E)\wedge{\bf c}(E^{\prime}\oplus E)={\bf s}(E)\wedge{\bf s}(E^{\prime})\wedge{\bf c}(E^{\prime}\oplus E)$. Multiplying by ${\bf c}(E^{\prime})$ from the right and ${\bf c}(E)$ from the left we get ${\bf c}(E)\wedge{\bf c}(E^{\prime})={\bf c}(E^{\prime}\oplus E)$.
From (2.4) we see that ${\bf c}_{k}(E,h)$ can be expressed as a sum of commuting compositions ${\bf s}_{\ell_{1}}(E,h)\cdots{\bf s}_{\ell_{m}}(E,h)$. That is,
${\bf c}_{1}(E,h)=-{\bf s}_{1}(E,h),\quad{\bf c}_{2}(E,h)={\bf s}_{2}(E,h)-{\bf s}_{1}(E,h)^{2},\quad etc.$
Now (iv) follows for $\bf c$}.
It remains to prove the analogous statements of Theorem 6.7 when ${\bf s}(E)$ is replaced by ${\bf c}(E)$. Part (ii) is obvious, and part (iii) for ${\bf c}(E)$ follows from (iii) for ${\bf s}(E)$. To see (iv), notice that ${\bf s}(E,h,h_{0})-{\bf s}(E,h_{0})=\hat{\bf s}(E,h,h_{0})-\hat{\bf s}(E,h_{0})$ and that $\hat{\bf s}(E,h,h_{0})$ and $\hat{\bf s}(E,h_{0})$ commute. By Theorem 6.7 (iv), for each $k$, there are $w_{k}$ such that
$dd^{c}w_{k}=({\bf s}(E,h,h_{0})-{\bf s}(E,h_{0}))\wedge\sum_{j=0}^{k-1}\hat{\bf s}(E,h,h_{0})^{j}\wedge\hat{\bf s}(E,h_{0})^{k-1-j}\wedge\mu.$
If $w=-w_{1}+w_{2}$-$\cdots$, then $dd^{c}w=({\bf c}(E,h,h_{0})-{\bf c}(E,h_{0}))\wedge\mu$. We get (v) for $\bf c$} from the statement for $\bf s$} by the same formal computation as above in the non-degenerate case. Finally, we get (vi) for $\bf c$} by repeated use of (vi) for $\bf s$}. ∎
Proof of Theorem 1.5.
The theorem follows from Proposition 7.1 and the fact that the Chern form $c(E,h,h_{0}):={\bf c}(E,h,h_{0})1$ is in ${\mathcal{QZ}}(X)$, cf. Section 4 in the same way that Theorem 1.3 follows from Theorem 6.7 and the fact that the Segre form $s(E,h,h_{0})$ is in ${\mathcal{QZ}}(X)$. ∎
If $h$ is qas and non-degenerate, then by the analogue of Theorem 6.5 (ii),
$c(E,h)=\lim_{{\epsilon}\to 0}c(E,h_{\epsilon})$
where $h_{\epsilon}$ are smooth metrics. In particular,
(7.3) $c_{k}(E,h)=0,\quad k>{\rm rank\,}E,$
since this holds for $c_{k}(E,h_{\epsilon})$. We do not know whether $c_{k}(E,h,h_{0})=0$ for $k>{\rm rank\,}E$ if $h$ is a general qas metric.
7.1. An alternative endomorphism
Let $(E,h)$ be a qas bundle on $X$ and a let $h_{0}$ be a smooth reference metric. We shall now consider an alternative extension to an endomorphism on ${\mathcal{QZ}}(X)$ which leads to a different Chern form, that appeared in [1]. We will use the notation $s(E):=s(E,h,h_{0})$ and $s(E_{0})=s(E,h_{0})$. Let $Z$ be the degeneracy locus of $h$.
We introduce the endomorphism
(7.4) ${\bf s}(E,Z)\wedge\mu={\bf s}(E)\wedge{\bf 1}_{X\setminus Z}\mu+{\bf s}(E_{0})\wedge{\bf 1}_{Z}\mu,\quad\mu\in{\mathcal{QZ}}(X).$
If $Z=X$, then it is just ${\bf s}(E_{0})$ so it is of interest only when $Z$ has positive codimension, which we will assume from now on. Note that if $s(E,Z):={\bf s}(E,Z)1$, then
(7.5) $s(E,Z)=s(E)$
since ${\bf 1}_{Z}1=0$ and ${\bf 1}_{X\setminus Z}1=1$ as long as $Z$ has positive codimension.
Proposition 7.2.
Let $E\to X$ be an qas bundle and $\mu\in{\mathcal{QZ}}(X)$.
(7.6) $({\bf s}(E,Z)-{\bf s}(E))\wedge\mu=({\bf s}(E)-{\bf s}(E_{0}))\wedge{\bf 1}_{Z}\mu.$
(ii) There is a current $w$, which is the pushforward of a current on $Z$, such that
(7.7) $({\bf s}(E,Z)-{\bf s}(E))\wedge\mu=dd^{c}w.$
Proof.
Equation (7.6) follows directly from the definition of ${\bf s}(E,Z)$. To prove (ii), first notice that since ${\bf 1}_{Z}\mu$ is in ${\mathcal{QZ}}(X)$ and has support on $Z$, by Proposition 4.4, there is a current $\mu_{Z}\in{\mathcal{QZ}}(Z)$ such that ${\bf 1}_{Z}\mu=i_{*}\mu_{Z}$, where $i\colon Z\to X$ denotes the inclusion. By Theorem 6.7 (iv) there is $w^{\prime}\in{\mathcal{QZ}}(Z)$ such that
$({\bf s}(i^{*}E)-{\bf s}(i^{*}E_{0}))\wedge\mu_{Z}=dd^{c}w^{\prime}.$
To be precise, we have formulated Theorem 6.7 only for a complex manifold, but the proof if parts (iii) and (iv) works just as well if $Z$ is any reduced analytic space. Hence, (7.7) follows by (i) together with Theorem 6.7 (iii) with $w=i_{*}w^{\prime}$. ∎
It follows from (7.4) that ${\bf s}_{0}(E,Z)\wedge\mu=\mu$ so that ${\bf s}_{0}(E,Z)=I$, the identity on ${\mathcal{QZ}}(X)$. We can thus define
(7.8) ${\bf c}(E,Z)\wedge\mu=\sum_{k=0}^{\infty}(-1)^{k}\hat{\bf s}(E,Z)^{k}\wedge\mu,$
where $\hat{\bf s}(E,Z)={\bf s}(E,Z)-I$.
Proposition 7.3.
Let $E\to X$ be an qas bundle and $\mu\in{\mathcal{QZ}}(X)$.
(7.9) ${\bf c}(E,Z)\wedge{\bf s}(E,Z)={\bf s}(E,Z)\wedge{\bf c}(E,Z)={I}.$
(ii) For each $\mu\in{\mathcal{QZ}}(X)$ there is a current $w$, which is the pushforward of a current on $Z$, such that
$dd^{c}w=\big({\bf c}(E,Z)-{\bf c}(E)\big)\wedge\mu.$
(iii) If $Z$ has positive codimension in $X$, then, if ${\bf c}^{\prime}(E):={\bf 1}_{X\setminus Z}{\bf c}(E)$,
(7.10) ${\bf c}(E,Z)\wedge\mu=\big({\bf c}^{\prime}(E)-{\bf c}(E_{0})\wedge M^{E}\wedge{\bf c}^{\prime}(E)\big)\wedge{\bf 1}_{X\setminus Z}\mu+{\bf c}(E_{0})\wedge{\bf 1}_{Z}\mu.$
By (ii), $c(E,Z):={\bf c}(E,Z)\wedge 1$ is cohomologous with $c(E)$ and hence it defines the expected Bott-Chern class.
By definition both $c(E)$ and $c(E,Z)$ are equal to the usual Chern form in $X\setminus Z$ if the metric on $E$ is smooth there. However, in general $c(E,Z)\neq c(E)$ across $Z$, see Example 8.1.
Remark 7.4.
In view of (iii), $c(E,Z)$ is precisely the Chern form that appeared in [1] in case the metric on $E$ is obtained from $F$ via a generically injective $g\colon E\to F$, cf. Example 6.3
Proof of Proposition 7.3.
Part (i) follows in the same way as (i) in Proposition 7.1. To see (ii) we first write
$\hat{\bf s}(E)^{k}-\hat{\bf s}(E,Z)^{k}=\sum_{\ell=0}^{k-1}{\bf s}(E,Z)^{k-\ell-1}\wedge({\bf s}(E)-{\bf s}(E,Z))\wedge{\bf s}(E)^{\ell}.$
By (7.7), for each $\ell$ there exists $w_{\ell}$, which is the push-forward of a current on $Z$, such that
$({\bf s}(E)-{\bf s}(E,Z))\wedge{\bf s}(E)^{\ell}\wedge\mu=dd^{c}w_{\ell}.$
Since $w_{\ell}$ has support on $Z$,
${\bf s}(E,Z)^{k-\ell-1}\wedge({\bf s}(E)-{\bf s}(E,Z))\wedge{\bf s}(E)^{\ell}\wedge\mu=s(E_{0})^{k-\ell-1}\wedge dd^{c}w_{\ell}=dd^{c}(s(E_{0})^{k-\ell-1}w_{\ell}),$
so we get (ii) with
$w=\sum s(E_{0})^{k-\ell-1}\wedge w_{\ell}.$
For (iii), recall that $M^{E}\wedge\mu:={\bf 1}_{Z}{\bf s}(E)\wedge\mu$ and ${\bf s}^{\prime}(E)\wedge\mu={\bf 1}_{X\setminus Z}{\bf s}(E)\wedge\mu.$ One checks that
(7.11) $\hat{\bf s}(E,Z)=\hat{\bf s}(E)\wedge{\bf 1}_{X\setminus Z}+\hat{s}(E_{0})\wedge{\bf 1}_{Z}$
We will prove by induction that
(7.12) $\hat{\bf s}(E,Z)^{k}=\hat{\bf s}^{\prime}(E)^{k}+\sum_{\ell=0}^{k-1}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell},\quad k=1,2,\ldots,$
when acting on ${\bf 1}_{X\setminus Z}\mu$. By definition it holds for $k=1$. Assume that it holds for $k$. When applying (7.11) to the right hand side of (7.12) the first mapping only acts on the first term and gives $\hat{\bf s}(E)\wedge\hat{\bf s}^{\prime}(E)^{k}=\hat{\bf s}^{\prime}(E)^{k+1}+M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k}$. The second mapping neglects the first term and just gives another factor $\hat{\bf s}(E_{0})$ on the terms in the sum. Thus (7.12) holds. Now
${\bf c}(E,Z)=I+\sum_{k=1}^{\infty}(-1)^{k}\Big(\hat{\bf s}^{\prime}(E)^{k}+\sum_{\ell=0}^{k-1}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell}\Big)=\\ {\bf c}^{\prime}(E)+\sum_{\ell=0}^{\infty}\sum_{k=\ell+1}^{\infty}(-1)^{k}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell}=\\ {\bf c}^{\prime}(E)-\sum_{\ell=0}^{\infty}(-1)^{\ell}{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf c}^{\prime}(E)={\bf c}^{\prime}(E)-{\bf c}(E_{0})\wedge M^{E}\wedge{\bf c}^{\prime}(E)$
when acting on ${\bf 1}_{X\setminus Z}\mu$. The action on ${\bf 1}_{Z}\mu$ is just $\hat{\bf s}(E_{0})^{k}\wedge{\bf 1}_{Z}\mu$. Thus (iii) follows. ∎
8. Remarks and examples
If $(E,h)$ is qas and its degeneracy locus $Z$ is empty, then, in view of (6.18) and (6.19), the Segre form $s(E,h)$ defined here seems to be the only reasonable one and thus canonical.
If $Z\neq\emptyset$, then any closed current of order 0 that coincides with $s_{k}(E,h)$ on $X\setminus Z$ is unique if $k<{\kappa}:={\rm codim\,}Z$. This follows from the dimension principle for normal currents, see, e.g., [13, Corollary III.2.11]. In particular, $s_{k}(E,h)=s_{k}^{\prime}(E,h)$, which coincides with the currents considered by Xia (for more general classes of metrics), [21], based on the non-pluripolar Monge-Ampère product. Since the class of $s_{k}(E,h,h_{0})$ in ${\mathcal{QB}}(X)$ is independent of $h_{0}$ it follows from Proposition 4.15 that $s_{\kappa}(E,h,h_{0})$ is independent of $h_{0}$. If the metric $h$ in addition is Griffiths negative, then by [17], $s_{k}(E,h)$ may be obtained as $\lim_{{\epsilon}\to 0}s_{k}(E,h_{\epsilon})$ for any sequence of smooth Griffiths negative metrics $h_{\epsilon}$ decreasing to $h$ as ${\epsilon}\to 0$ for any $k\leq{\kappa}$. On the other hand, [17, Example 8.2] shows that $s_{k}(E,h,h_{0})$ might depend on $h_{0}$ for any $k>{\kappa}$. It follows from Theorem 1.3 (i) that the limits $s(E,h_{\epsilon})$ as ${\epsilon}\to 0$ might depend on the sequence $h_{\epsilon}$, even when $h_{\epsilon}$ is a sequence of Griffiths negative smooth metrics decreasing to $h$. Therefore it does not seem to be any ’canonical’ choice of Segre form $s_{k}(E,h)$ for $k>{\kappa}$.
Similar statements as above hold also for products
$s_{k_{t}}(E,h)\wedge\dots\wedge s_{k_{1}}(E,h):={\bf s}_{k_{t}}(E,h)\wedge\dots\wedge{\bf s}_{k_{1}}(E,h)1,$
depending on how $k=k_{1}+\dots+k_{t}$ compares to $\kappa$}. Such a product is furthermore robust in the sense that it is commutative in the factors provided $k\leq{\kappa}$, which follows from Theorem 6.7 (vii), cf. Remark 6.8 (ii).
By Theorem 6.5 (iii), ${\bf s}(E,h+{\epsilon}h_{0})^{k}\mu=s(E\oplus\cdots\oplus E,h\oplus\cdots\oplus h+{\epsilon}(h_{0}\oplus\cdots\oplus h_{0})\wedge\mu$ for each $k$. Therefore $\lim_{{\epsilon}\to 0}c(E,h+{\epsilon}h_{0})\wedge\mu$ exists and is in ${\mathcal{QZ}}(X)$. Moreover it coincides with $c(E,h)\wedge\mu$ outside the degeneracy locus. However, in general it is not equal to ${\bf c}(E,h,h_{0})\wedge\mu$. This will be discussed in a forthcoming paper.
8.1. Difference between $c_{k}(E)$ and $c_{k}(E,Z)$
In general $c(E,Z)$ and $c(E)$ are different. The simplest example is when $Z=X$ and $Z^{\prime}\neq Z$. Then in general $s(E)$ will have a residue term, whereas $s(E,Z)=\omega$ and hence $c(E,Z)=c(E_{0})$.
Example 8.1.
Let $g$ be the $2\times 2$-diagonal matrix, with entries $x_{1}$ and $x_{2}$, considered as a morphism between the trivial bundles $E=X\times{\mathbb{C}}^{2}_{\alpha}=F$, and $F$ is equipped with the trivial metric. With the trivial metric on $E$ as reference metric $h_{0}$ and the singular metric $h$ induced by $g$, we have
(8.1) $s_{1}(E)=[x_{1}=0]+[x_{2}=0].$
This follows by direct computation, or, e.g., from [1, Proposition 11.1]. From [1, Example 11.2] we have that $s_{2}(E)=[x_{1}=x_{2}=0]$.
Let us now compute $c(E)$ and $c(E,Z)$ for the same qas bundle $E$. To begin with $c_{1}(E)=c_{1}(E,Z)=-s_{1}(E)=-([x_{1}=0]+[x_{2}=0])$. A calculation yields that
(8.2) ${\bf s}_{1}(E)^{2}={\bf s}_{1}(E)\wedge([x_{1}=0]+[x_{2}=0])=2[x_{1}=x_{2}=0].$
For instance, if $\tau\colon{\mathbb{C}}\to{\mathbb{C}}^{2}$, $t\mapsto(t,0)$, then ${\bf s}_{1}(\tau^{*}E)=[t=0]$. Hence, by Theorem 6.7 (iii), ${\bf s}_{1}(E)\wedge[x_{2}=0]=\tau_{*}(s_{1}(\tau^{*}E))=\tau_{*}[t=0]=[x_{1}=x_{2}=0]$. From [1, Example 11.2] we have that $s_{2}(E)=s_{2}(E,Z)=[x_{1}=x_{2}=0]$. Thus
$c_{2}(E)=s_{1}(E)^{2}-s_{2}(E)=[x_{1}=x_{2}=0].$
Since $Z=\{x_{1}x_{2}=0\}$, and $s_{1}(E)=s_{1}(E,Z)$ has support on $Z$, cf. above and (7.5),
$s_{1}(E,Z)^{2}={\bf s}_{1}(E,Z)\wedge([x_{1}=0]+[x_{2}=0])=0,$
since ${\bf s}_{1}(E_{0})=0$. We conclude that
$c_{2}(E,Z)=-[x_{1}=x_{2}=0],$
which is different from $c_{2}(E)$. ∎